### v3

Abstract
We describe the design, implementation, and evaluation of a low-cost (approximately \$15 per person) and scalable program, called Challenges, aimed at aiding women in Poland transition to technology-sector jobs. This program helps participants develop portfolios demonstrating job-relevant competencies. We conduct two independent evaluations, one of the Challenges program and the other of a traditional mentoring program – Mentoring – where experienced tech professionals work individually with mentees to support them in their job search. Exploiting the fact that both programs were oversubscribed, we randomized admissions and measured their impact on the probability of finding a job in the technology sector. We estimate that Mentoring increases the probability of finding a technology job within four months from 29% to 42% and Challenges from 20% to 29%, and the treatment effects do not attenuate over 12 months. Since both programs are capacity constrained in practice (only 28% of applicants can be accommodated), we evaluate the effectiveness of several alternative prioritization rules based on applicant characteristics. We find that a policy that selects applicants based on their predicted treatment effects increases the average treatment effect across the two programs to 22 percentage points. We further analyze how alternative prioritization rules compare to the selection that mentors used. We find that mentors selected applicants who were more likely to get a tech job even without participating in the program, and the treatment effect for applicants with similar characteristics to those selected by mentors is about half of the effect attainable when participants are prioritized optimally.
1 Introduction
An important concern of labor policymakers is identifying and implementing effective programs to support workers transitioning into occupations and industries that offer attractive jobs. One sector that has experienced rapid growth in recent decades and offers many attractive career opportunities is the technology sector. However, not all societal groups have benefited equally from opportunities in this sector; in particular, women remain underrepresented. Despite a number of programs and initiatives aimed at supporting women in the technology sector, [^1] For example, the United Nations organizes an annual event Girls in ICT Day; The European Commission carries out an annual study: Women in Digital Scoreboard. There are various NGOs dedicated to increasing women’s participation in the technology sector; For instance, Women in Technology International is a community of 160,000 women in the technology sector; Girls Who Code operates a network of 10,000 clubs providing support for women in technology. Finally, many major technology companies have initiatives aimed at supporting women in technology, including Women in Tech from Salesforce, The Women at Microsoft, or Accenture’s Women in TS&A. women held only 30% of technology jobs worldwide in 2021 (United Nations, 2021). The shortage of empirical evidence evaluating the effectiveness of specific strategies to address this issue is one of the hurdles in designing a successful policy response.
In this paper, we ask whether interventions can have an impact on this problem, and if so, which workers benefit most. We conduct randomized evaluations of two distinct programs, both implemented in Poland. The first, Mentoring, involves one-to-one mentoring and is thus not easy to scale. We designed and introduced a new program, Challenges, to be scalable. In Challenges, women work in teams to create portfolios to show to employers. Second, we ask which women benefit most from each program. Using data from the two evaluations, we identify the distinct subgroups with the largest benefits. Third, we evaluate the effectiveness of alternative prioritization rules for admission to the program in the presence of capacity constraints, illustrating that there are large gains from the optimal prioritization. Fourth, we examine the implicit selection rule used by mentors, who each selected two candidates from a pool of applicants. We find that mentors do not select the candidates with highest estimated treatment effects, but rather focus more on candidates who would be likely to succeed even without the program.
The results from our randomized evaluation suggest that the Challenges program increases the likelihood of successfully transitioning to the technology sector by approximately 45% within four months of graduation. [^2] Throughout this paper, we consider technology jobs as all jobs in firms in the technology sector other than those in finance, regulatory, legal, accounting, or human resources or jobs outside of the technology sector that involve software development and testing, IT support, and data analytics. Specifically, all jobs in UX and front-end programming are considered technology jobs. Challenges helps participants develop portfolios demonstrating capability relevant to the jobs they seek. The program was designed in partnership with technology companies, delivered online, and relies on interactions between program participants, which makes it cheap to operate (approximately \$15 per person) and suitable for scaling up. The program is also transferable across domains of the technology sector and languages; the first edition, evaluated in this paper, had specializations in user experience design and front-end programming. The program’s second edition was offered in Ukrainian instead of in Polish and offered additional specializations in software testing, cloud computing, and user interface design.
We designed the Challenges program as a scalable alternative to traditional in-person mentoring. To develop it, we partnered with Dare IT, a Polish organization offering a non-profit mentoring program – Mentoring, since 2018. Dare IT’s mentoring program serves women who, despite having invested time into acquiring IT skills, still have not successfully obtained a technology job. There is a high demand for this program, as evidenced by the fact that it has been repeatedly oversubscribed. However, scaling it up to meet the demand faces several challenges. First, it is difficult to find enough mentors. Mentors operate only on a volunteer basis and tend to be women in mid-seniority technology jobs. Since few women are in these positions, recruiting more mentors is difficult. Second, the one-to-one format of Mentoring makes it operationally demanding. Ensuring satisfaction, remedying conflicts, and tracking everyone are complicated tasks that require regular supervision by Dare IT staff.
The development of Challenges proceeded in two steps. First, together with Dare IT, we set up a series of interviews with HR experts and hiring managers of Polish technology companies to gain their perspectives on ways to support women candidates. Each of our interviewees mentioned the shortage of practical experience or clear signals of practical skills in women applicants’ profiles. Second, along with Dare IT and industry partners, we used the insights from the interviews to create Challenges. The program’s core consists of assignments sent every second week that, if completed, result in a job-relevant skill signal that can be added to participants’ resumes or mentioned during an interview. The program is designed based on three principles: (i.) it is delivered online, which lowers the costs; (ii.) it primarily involves interactions between participants, addressing the shortage of mentors; and (iii.) it is closely based on recent business cases. To operationalize the last principle, industry partners were recruited to help design the assignments and provide domain experts to review and critique selected submissions. We launched the registration to the program in December 2021 with 300 available spots; the program became oversubscribed within four hours.
We conducted two concurrent randomized experiments to evaluate the effectiveness of Challenges and Mentoring. Both experiments exploit the fact that the two programs were oversubscribed, so we were able to randomize access and evaluate the impact of participation in a program on the probability of having a technology job. To evaluate Mentoring, we asked mentors to select two mentee candidates (instead of the usual one), and within these pairs, we randomized into treatment and control. In the case of Challenges, we randomly assigned access to 200 spots in the program. Four months after the end of each program, we collected data on the outcomes. We estimate that participation in Mentoring increased the probability of having a job in technology from 29% to 42%, while the Challenges program increased this probability from 20% to 28%. [^3] The difference in the outcomes of the two control groups is likely because, in Mentoring, mentors selected applicants which were later randomized into the control group. The control group of Challenges is a random draw from the pool of applicants. We document that the outcomes of non-selected applicants are substantially lower than those of the applicants selected by mentors and later randomized into the control group. We repeated the evaluation 12 months after the end of the two programs and found that their impact did not attenuate over time.
We find that the two programs lead to different patterns of treatment effect heterogeneity. Mentoring is particularly beneficial for women from smaller towns that do not have a STEM degree, are mothers, and have been active in the labor market for over five years. Working with mentors with managerial experience and a longer tenure in technology also leads to high treatment effects. In the case of Challenges, participants who have just entered the labor market without a STEM degree exhibit high treatment effects. Different patterns of heterogeneity in treatment effects combined with insights from our discussions with Dare IT, mentors, and industry experts suggest that the two programs support participants’ employability through different mechanisms. Mentoring can positively impact the chances of obtaining employment in the technology sector by reinforcing self-efficacy, improving interview skills, or expanding professional networks. In contrast, creating portfolio items in the Challenges program that constitute signals of specific skills relevant to the job can counter a pessimistic belief (e.g., based on stereotypes) held by hiring managers, recruiters, or applicants themselves.
The core motivation for creating the Challenges program was to extend Dare IT’s impact beyond what is possible via the relatively small Mentoring program. Although Challenges expanded the number of people who receive Dare IT services, the existing capacity constraints in Mentoring and Challenges present a significant hurdle, as the number of interested applicants greatly exceeds the number of available spots in each program. Furthermore, at the time this research was conducted, Dare IT was not using paid advertising to promote its services, so the applicants to Dare IT heard about the program through word-of-mouth recommendations or through an internet search, which underscores that the potential demand for Dare IT services is even higher. In the case of Mentoring, the constraint is caused by the limited number of mentors and the operational difficulties of scaling up the program. For the Challenges program, the number of spots has been intentionally limited to 300 to maintain appropriate participant support. [^4] The first 100 applicants were automatically granted a spot in Challenges and were removed from the analysis conducted in this paper. Even though there are plans to incrementally increase the number of available spots in future Challenges cohorts, Dare IT will continue to grapple with the decisions of which applicants to admit to the Challenges program and how to select participants for the Mentoring program in the foreseeable future.
In the presence of capacity constraints and substantial heterogeneity in treatment effects, admission rules prioritizing applicants with the highest benefits can potentially enhance Dare IT’s impact. Adapting the methodology of Athey and Wager (2021) and Zhou et al. (2023), we estimate the value of counterfactual assignment policies that maximize outcomes across all applicants, given the constraints on the number of spots available in each program. These policies are then evaluated in the test set using doubly-robust estimates obtained through cross-fitting.
We estimate that the existing policy, which assigns 15% of all applicants to Challenges and 13% in Mentoring, leads to the weighted average treatment effect of 12 percentage points increase in the probability of getting a tech job. Considering the case in which the available number of spots in each program does not change, but we change the assignment policy to prioritizes applicants with the highest predicted benefits from each program (we refer to this policy as optimal policy), we show that the average treatment effect increases to 22 percentage points. Thus, we find a 10 percentage points incremental benefit of targeting. Second, we evaluate the optimal policy at various capacity levels. While expanding capacity increases the average outcomes as applicants who previously did not participate are admitted, the incremental value of targeting over random admissions might decrease, as applicants with the highest predicted benefits are allocated first, and the marginal participants have lower predicted benefits. However, we estimate that targeting has substantial value even at high capacity levels. We show that when the Challenges program is expanded to 50% of all applicants, the incremental benefit of targeting is 5 percentage points, while when there are no capacity limits, the incremental benefit of targeting is to increase the probability of finding a job by 3 percentage points. The comparison of random and targeted admissions at the full capacity level, underscores that some participants have higher benefits from one program than from the other; thus, the value of targeting is not only in targeting between programs and out of programs, but also across the two offerings.
Dare IT ensures high-quality matching by asking mentors to choose mentees from a list of pre-screened candidates. When comparing the outcomes of applicants to Mentoring, who were selected by mentors but randomized to the control group, with those not selected, we find that mentors tend to pick promising candidates. These candidates will likely find jobs in the technology sector, even without participating in the program. However, this approach to selecting participants might not be maximizing the potential impact of Mentoring. We use the off-policy approach to evaluate an alternative policy where mentors select candidates with the highest treatment effects. We find that the policy of choosing applicants based on their treatment effects leads to an average probability of getting a job in the technology sector of 23% across all applicants to Mentoring (we exclude applicants to Challenges from this analysis), compared to 20% when selecting the most promising candidates. This finding suggests that focusing on those most likely to benefit from the program rather than those who seem most promising overall could increase the program’s effectiveness.
Finally, a non-parametric assignment policy which requires solving a discrete optimization program, might not be realistic in the context of Dare IT. Therefore, to create viable alternatives which lead to similar benefits as optimal policies but are easier to implement, we propose simple rules that define priority groups based on observed characteristics (e.g., applicants from smaller cities are prioritized for Mentoring). We then evaluate these rules using predicted treatment effects. We find that prioritization can substantially increase the two programs’ effectiveness.
2 Literature review
This paper relates to several strands of literature. First, the paper contributes to the literature on the effectiveness of mentoring in improving career outcomes. Several influential papers focus on the effectiveness of mentoring in the academic context: Dennehy and Dasgupta (2017) employ a randomized experiment to show that assigning a female mentor to a female engineering student increases the sense of belonging in the discipline, self-efficacy, motivation, and retention for women majoring in engineering in college. Resnjanskij et al. (2021), also in a randomized experiment, document a mentoring program’s positive impact on disadvantaged high-school youth’s math grades and career plans. Several papers focus on the impact of mentoring on the careers of academic faculty members. Gardiner et al. (2007) provide empirical evidence that mentored individuals are more likely to stay at a university, receive more grant income, and receive promotions more frequently. In more recent studies, Ginther et al. (2020) and Blau et al. (2010) analyze the effectiveness of a mentorship workshop to support women in research careers. In a randomized access design similar to ours, they show that individuals in the treatment group are more likely to receive tenure at their institution, have more publications and citations, and receive more federal grants. We contribute to this literature by providing evidence from a randomized trial of the efficacy of mentoring outside of the context of education or academic jobs. [^5] The impact of mentoring on career outcomes has also been studied using surveys and other techniques based on observational data. For example, Ragins and Cotton (1999), in a survey of over 600 individuals, find that mentored respondents have better career outcomes than non-mentored respondents. They also find that the gender composition of the mentoring pair influences the effectiveness of mentoring. Similarly, Aryee et al. (1996) survey a group of workers in Hong Kong and found that mentored respondents reported a higher number of promotions and were more satisfied with their work than those not mentored. Kammeyer-Mueller and Judge (2008) carry out a meta-analysis of this literature and concludes that mentoring improves job and career satisfaction.
Second, there is rich evidence evaluating the efficacy of offline labor market training programs. Card et al. (2018) carry out a comprehensive meta-analysis of the estimates from over 200 recent studies. They find that on average, the analyzed training programs have close to zero short-term effects on employment and some small positive impact in 2-3 years after completion. Programs evaluated in Card et al. (2018) generally aim at unemployed individuals looking for blue-collar jobs. [^6] Other influential papers evaluating labor training programs have a similar focus: Barnow (1987); Bloom et al. (1997); Heckman et al. (1999). In this category, programs focused on the development of skills on-the-job tend to be more effective (Sianesi, 2008; Lechner and Gerfin, 2000). On-the-job training programs do not necessarily need to take the form of long apprenticeships to be effective: Biewen et al. (2014) compare a long apprenticeship with short-term on-the-job training, combined with a job search workshop, and find that the short-term program is equally effective or more effective than the much longer program. Fein and Hamadyk (2018) use a randomized experiment to evaluate YearUp, a sectoral training program, and find a 40% increase in quarterly earnings in the medium term. We contribute to this literature by bringing evidence from a randomized trial of a training program that happens online, is focused on individuals joining a highly paid technology sector, and is very cheap to operate. Our program has on-the-job training elements as participants develop content designed by domain experts to closely resemble the content they would produce on the job.
Third, a few studies have analyzed the efficacy of Massive Open Online Courses (MOOCs) in supporting learners’ job search. For instance, Zhenghao et al. (2015) shows that a significant majority of Coursera participants reported career benefits, particularly those from economically- and academically-disadvantaged backgrounds. However, through a difference-in-differences approach, Castaño-Muñoz and Rodrigues (2021) finds no effect on wages or employment among MOOC learners in Spain and a moderate impact on job retention. Hadavand et al. (2018) compare participants in a Data Science MOOC that were just above the threshold to graduate with those just below and argue that graduation from this MOOC leads to a 30% increase in the probability of finding a job. Our study contributes to this literature by using a randomized experiment approach. We also compare an online training program with an in-person Mentoring, identifying diverse patterns of heterogeneity that suggest different priority groups for online versus offline learning.
Fourth, we contribute to the literature on skill signaling (Spence, 1978). Previous research showed that credentials demonstrating capability can significantly boost earnings Tyler et al. (2000); Hadavand et al. (2018). Furthermore, in situations where companies might be reticent about hiring less experienced candidates, establishing trustworthy indicators of ability can bring advantages at the market level Pallais (2014). Our research, through a randomized trial, sheds light on the efficacy of a program aimed at developing labor market signals. Our subjects are women seeking tech roles who can utilize skill signaling to counter possible negative stereotypes. However, since the programs we evaluate incorporate both learning and signaling, we cannot separate the signaling effect.
Fifth, we document a high degree of heterogeneity in treatment effects from both programs, which we use to define treatment assignment policies. We show that targeting based on treatment effects leads to substantially higher average outcomes than random targeting, or targeting based on lowest outcomes in the absence of treatment, or highest outcomes with treatment. In doing so, we contribute to a small but growing literature on targeting treatment based on estimated treatment effects. An early contribution to this area was made by Ascarza (2018), who using randomized experiments, documents that treatment effect-based targeting is more effective in customer retention than churn-based prediction. Similar findings have been noted in marketing and development economics and medical contexts (Hitsch and Misra (2018), Devriendt et al. (2021), Haushofer et al. (2022), Inoue et al. (2023)). Lastly, a comparison of assignment policies for financial aid application nudges by Athey et al. (2023) highlights the superiority of maximizing treatment effects over targeting based on the predicted risk of non-compliance.
Finally, this paper adds to the literature on the gender gap in the tech sector. Various reasons for low women’s participation in the technology sector have been studied. For example, Murciano-Goroff (2022) identify women’s tendency to under-report programming skills, impacting their job prospects. Strategies proposed to counter this include increasing role model visibility (Correll and Mackenzie (2016)) and shifting the perception of tech roles being predominantly male (Cheryan et al. (2013), Del Carpio and Guadalupe (2022)). We contribute to this literature by demonstrating how mentoring and job-training programs can enhance women’s success rates in obtaining tech jobs. Notably, our work focuses on the primary outcome of finding a job.
3 Mentoring and Challenges Programs
Over recent decades, Poland’s technology sector has witnessed significant growth, culminating in a 2021 revenue of roughly USD 51 billion, which accounts for 8% of the national GDP and provides employment for 250,000 individuals (GUS, 2022a). The sector’s growth has contributed to an escalating demand for skilled tech professionals, subsequently driving an approximate 9% wage growth in 2021. This led to the sector’s median monthly wage reaching USD 2,600, substantially outpacing the country’s overall median monthly wage for full-time workers, which stands at approximately USD 950 (GUS, 2022b). A survey of gender demographics within this industry unveils an underrepresentation of women, who make up 30% of the total Polish tech workforce (Gawlowska-Bujok and Bujok, 2021). The European Commission (European Commission, 2021) reports that the tech industry accounts for 1.1% of Polish women active in the labor market, compared to 5.2% of men.
Dare IT is a social impact organization that aims at reducing the gender gap in the tech sector. They do so by helping qualified women transition from other careers. Dare IT offers a variety of services for women looking for jobs in tech or making their first steps in the profession, including a mentoring program, career fairs, networking events, and an online social network.
Mentoring.
Dare IT launched Mentoring in 2018 as a not-for-profit initiative to support women aspiring to tech jobs. Since then, the program has been offered twice per year, and the number of participants has increased in each cohort. The program is funded by grants and donations, which allows participation to be free of charge. In recent years, demand for the program has been far greater than supply. For example, for the spring 2021 cohort, Dare IT received over 1,000 applications for 150 spots.
The program follows several steps. Each cycle begins with Dare IT recruiting potential mentors through past participant referrals, social media engagement, and corporate partnerships. All mentors are women who are active professionals in the tech sector, and they contribute their time and expertise to the program on a pro-bono basis. Mentors are screened and selected to participate based on an application process; Dare IT considers multiple criteria, including candidates’ expertise in domains that attract the most interest from mentees, length of professional experience, and reputation.
Simultaneously, the recruitment of potential mentees occurs primarily via social media and word-of-mouth recommendations. During this study, Dare IT did not utilize paid advertisements for recruitment. Dare IT selects participants based on several criteria. Only fully-completed applications by women not currently engaged in a tech role are considered. The program primarily targets individuals with necessary skills — quantified by self-reported learning hours in a particular domain and manual review of applications by Dare IT employees — who require additional support in their tech job search. Applicants who do not meet a baseline skill level are excluded. After Dare IT finalizes the selection process, mentors receive access to a database of eligible applicants. They can then select their preferred mentee.
Once mentors and mentees are paired, they typically meet on a weekly basis for four months in order to support the mentee in their job search. Dare IT provides mentors with guidelines and best practices to increase the likelihood of a successful mentorship experience, although mentors have the flexibility to modify these guidelines to cater to their mentee’s needs. As a result, mentoring takes various forms, from coaching in job searching and interviewing or resume preparation to reviewing code and developing technical skills. Our discussions with mentors and Dare IT staff suggest four potential channels through which Mentoring increases the employability of mentees: (i.) reinforcement of self-efficacy through the provision of emotional support and encouragement, (ii.) help in the development of a professional network by relaying on connections of mentors [^7] Pallais and Sands (2016) emphasizes how referrals can augment job opportunities and delves into the underlying mechanisms., (iii.) explanation of unwritten norms of professional behavior and conduct, especially those intrinsic to the tech sector and, more specifically, during tech recruitment phases, and (iv.) direct transfer of skills through code review and discussions.
Challenges.
Due to the excess demand for Mentoring, Dare IT explored the possibility of expanding its capacity. However, scaling proved to be a complex issue. First, selecting more mentors without sacrificing the quality of mentoring is difficult due to the scarcity of women occupying suitable roles in tech. Second, implementing a large 1:1 mentoring program presents operational difficulties.
We partnered with Dare IT to develop a more easily scalable approach. Our first step was forming a set of hypotheses on barriers that impede women from securing roles in the Polish tech sector. We conducted interviews with hiring managers and HR experts from a dozen prominent tech companies in Poland. Each of our interviewees mentioned a shortage of practical experience or signals of practical skills on job application materials provided by female candidates. They emphasized that tangible proof of hands-on experience, even a short project, could significantly alter the interview dynamics, typically favoring candidates. They also expressed low trust in skill credentials that are available online. Many of our interviewees described online credentials as needing to be more informative about the capability to carry out real-world business tasks. Motivated by these findings, we developed a new program called Challenges, designed to address these specific concerns.
The Challenges program comprises six bi-weekly mini-projects. Successful completion of all mini-projects results in a portfolio item, e.g., a UX design for a mobile application, which functions as proof of practical experience in the chosen domain (see Figure 1 for an example of wireframes from a completed product). The program strongly emphasizes on-the-job training: assignments are curated by domain experts and are readily applicable in a business setting. Each weekly mini-project undergoes evaluation and feedback from domain practitioners, including existing mentors and industry professionals, facilitating ongoing skill and knowledge enhancement for participants. A more comprehensive discussion of the way the Challenges program is organized and it’s content is in Appendix A.
Figure 1: Challenges final product.
Note: Selected wireframes from a design of a mobile app for pets adoption prepared during Challenges. See: https://www.behance.net/gallery/142047393/Pet-Adoption-App-UXUI for the entire design.
We launched the program in December 2021, setting the start date for the initial cohort in mid-January 2022. Dare IT promoted the program on its social media channels to encourage applications. Enrollment was capped at 300 participants. [^8] The decision to limit program capacity to 300 spots was made to account for potential logistical challenges and to ensure that we could provide personalized assistance to all participants if necessary. As the program progressed, it became evident that only a few participants required hands-on support. Consequently, in subsequent cohorts, Dare IT increased the number of available spots. The first 100 slots were allocated to the initial 100 applicants (excluded from this analysis), with the remaining 200 spots randomly assigned among subsequent applicants. The program was oversubscribed within four hours, at which point we closed the registration form. Applications received post-closure were not included in the randomization. However, we collected data on these late applicants’ outcomes for comparison.
4 Experimental design and data
4.1 Experimental design
Mentoring.
We evaluate the mentoring program using the cohort of Autumn 2021. As in previous cohorts, Dare IT received a significantly higher number of applications than the available number of mentors. Even after the screening process to verify eligibility, the ratio of applicants to mentors remained at 4:1. The eligibility for the Mentoring program underwent manual verification by Dare IT staff. Only applicants who did not currently hold positions in the technology field and who, according to Dare IT’s reviewer assessments, have technical skills at a level typical of a junior employee in a tech company were considered eligible for participation. The surplus of eligible applicants compared to the number of available mentors provided us with an opportunity to implement a randomized access experiment. In order to maintain high-quality mentor-mentee pairings while incorporating randomized access, we asked mentors to select two potential mentees instead of just one, as they did in the previous program’s cohorts. Within these pairs, we then randomly determined which of the two applicants would be granted access to the program (the treatment group) and which would not (the control group). Thus, we created a paired experiment.
In the experiment, we initially had 156 mentor-mentee pairs and an equivalent number of mentor-control pairs. However, four pairs did not start the program because the mentee secured a tech job between applying for the program and its commencement. In these cases, the mentor-selected applicant, who was initially randomized into the control group, participated in the program. [^9] We decided against having idle mentors to ensure broader access to the program, as well as, to respect mentors’ time dedicated to selecting candidates and willingness to volunteer their time in mentoring. Consequently, we excluded these mentors and participants from our analysis. As a result, our analysis included a total of 152 pairs. Additionally, we gathered data on 468 eligible applicants whom mentors did not select for participation.
Challenges.
Regarding the Challenges program, we concluded the recruitment process once we received 500 applications that met the established eligibility criteria. To qualify for participation, applicants were required to declare that they were not currently employed in the technology sector, possess a fluency level in English (B1 level), and be proficient in Figma or React depending on the specialization. The first 100 applicants were granted program spots, while the remaining 400 were randomly assigned to either the treatment or control groups. Randomization was conducted with stratification based on variables obtained from the registration survey. [^10] These stratification variables included specialization (front-end development or UX), age, type of residence, and employment status (student, employed, neither employed nor student). Additionally, we gathered data on 160 applicants who submitted their forms after the registration process had concluded. These late applicants were not included in the primary program evaluation. However, they are considered in the off-policy evaluation, as detailed in Section 6.
Furthermore, it’s worth noting that some applicants randomized into the treatment group did not actively participate in the program (did not submit any assignments). Nevertheless, since these individuals still received all program materials and had the opportunity to engage in webinars discussing projects, we have included all such individuals in our analysis.
4.2 Data
We have four data sources: applicant registration surveys, mentor registration surveys, outcome surveys conducted following the program’s conclusion, and information collected from participants’ LinkedIn profiles.
Applicant registration survey.
Applicants in both programs were asked to complete a simple registration survey. Data from the survey included the city of residence, age, employment situation, and time spent, which measures how much time the applicant had spent developing their IT skills. The mentoring application form was richer; first, applicants were asked if they have family or friends in IT and whether they have children; second, the form also had several open-ended questions related to the applicant’s motivation and the possibility of attaching a short introductory video, but we do not use either in this analysis. Additionally, we collected information on participants’ preferred domains: the main categories were UX design, manager (project or product), or developer (front or back-end).
Mentors registration survey.
Mentors were asked a number of questions before the start of the program, which allowed us to construct the following variables: first-time mentor, career changer (which indicates whether the mentor herself had worked outside of IT before moving to the sector), managerial experience (which indicates whether the mentor has any managerial experience or not), and years of experience (years of experience in the tech sector).
Outcomes survey.
We surveyed participants of each program and those in control groups four months after the end of the program. In the survey, we were primarily interested in the salaries in the new jobs: we asked the question whether the respondent found a new job and in which range the salary was. We also inquired about the number of job offers and whether the participant negotiated the proposed offer or accepted the initial one. Finally, we carried out a shorter survey right after the end of Mentoring, the “short survey,” where we asked for salaries for the new jobs. Altogether we conducted three surveys: the main mentoring survey in which we received 162 responses (with 111 responded from treatment group subjects), the initial survey of the mentoring program (with 194 responses), and the Challenges survey, for which we received responses from 68 subjects in the control group and 61 from the treatment group. We analyze the outcomes from surveys in Appendix F.
LinkedIn data.
Our primary source of outcome data is the LinkedIn profiles of the participants. This data was collected at two different points: first, four months, and second, 12 months after the conclusion of the programs. During the initial data collection phase, outcomes for all experimental subjects were gathered. In the second phase, Dare IT was unable to observe the outcome of one Mentoring participant, 17 Challenges participants, and 7 Challenges control group members. The principal cause of the missing data is the change in the LinkedIn profile address. [^11] It’s plausible that the propensity to change LinkedIn’s profile address was impacted by participation in the program. In Appendix ate_bounds we provide robustness checks to our primary results by assuming the most pessimistic and optimistic outcomes for the missing subjects. We also estimate AT bounds following the methodology in Lee and Weidner (2021). The primary outcome metric was whether a participant secured a technology job following a program’s start. As new technology jobs, we consider jobs with a starting date of January 2022 or later for Mentoring and February 2022 or later for Challenges that fall into one of the following two categories:
  1. 1.
    All jobs in technology companies other than positions in finance, regulatory, legal, accounting, and HR, where technology companies include firms in software development, testing, and sales; data analytics; IT services; digital marketing; and online platforms (including peer-to-peer platforms, and online shops).
  2. 2.
    Jobs in non-technology companies that involve software development and testing, IT support, and data analytics. In our context, this category includes jobs in banks and management consulting agencies.
Based on these categories, we define three outcomes: new job, tech job 4 months, and tech job 12 months. New job refers to any job, whether in technology or not, that started between January and May 2022 for Mentoring and between February and June for Challenges. Tech job 4 months denotes a job that aligns with category (1) or (2) and is also classified as a new job. Finally, tech job 12 months is a job, corresponding to either category (1) or (2), that began after January 2022 for Mentoring and February for Challenges and before May 2023. We also collect the date on which subjects got their new jobs. In some cases, we do not observe the month in which the new job started, but only the year; for such observations, we assign the start date as December 2022 or May 2023, depending on whether the job started in 2022 or 2023.
We also leveraged LinkedIn to gather data on the highest education level attained by participants, which was organized into the following categories: bachelor’s degree, master’s degree, postgraduate degree, non-traditional credential, and other. The other category includes profiles lacking education information or information that could not be classified into any of the pre-existing categories. Furthermore, we compiled data on college majors, forming two binary variables: STEM and social science or business. The former includes all STEM degrees, excluding social sciences, while the latter comprises degrees in social science, business, finance, or management. Lastly, we deduced a person’s age from a combination of age reported in the registration survey and the commencement date of their latest educational degree. Finally, we also compute their professional experience length by subtracting the start date of their first job recorded on LinkedIn from 2022. Table 1 summarizes the key variables for both programs.
Table 1: Summary statistics of the main variables: treatment, control, and applicants
Variables Treated group Control group
N Mean St. Dev. Min Max N Mean St. Dev. Min Max
Mentoring: mentor characteristics
First time mentor 152 0.467 0.501 0 1 145 0.434 0.497 0 1
Managerial experience 152 0.520 0.501 0 1 145 0.510 0.502 0 1
Career changer 152 0.691 0.464 0 1 145 0.710 0.455 0. 1
Years of tech experience 152 5.856 3.777 2 35 145 6.021 3.794 2 35
Mentoring: experimental groups
New job 152 0.467 0.501 0 1 147 0.422 0.496 0 1
Tech job 4 months 152 0.421 0.495 0 1 147 0.293 0.456 0 1
Tech job 12 months 151 0.609 0.499 0 1 147 0.456 0.500 0 1
Social science 152 0.132 0.339 0 1 147 0.136 0.344 0 1
STEM 152 0.487 0.501 0 1 147 0.531 0.501 0 1
Family friends IT 152 0.697 0.461 0 1 147 0.701 0.460 0 1
Warsaw 152 0.276 0.449 0 1 147 0.259 0.439 0 1
Mother 152 0.257 0.438 0 1 147 0.259 0.439 0 1
Applied before 152 0.243 0.431 0 1 147 0.231 0.423 0 1
Years of professional experience 152 7.053 5.049 0 31 147 7.204 4.816 0 23
Challenges: experimental groups
New job 183 0.333 0.473 0 1 225 0.293 0.456 0 1
Tech job 4 months 183 0.284 0.452 0 1 225 0.196 0.398 0 1
Tech job 12 months 166 0.380 0.487 0 1 218 0.266 0.443 0 1
Social science 183 0.208 0.407 0 1 225 0.267 0.443 0 1
STEM 183 0.421 0.495 0 1 225 0.391 0.489 0 1
Years of professional experience 183 7.645 4.976 0 31 225 7.671 6.237 0 31
Applicants both programs
Tech job 4 months - - - - - 628 0.152 0.359 0 1
Tech job 12 months - - - - - 628 0.279 0.449 0 1
STEM - - - - - 628 0.196 0.367 0 1
Years of professional experience - - - - - 628 7.474 6.318 0 37
Note: Summary statistics of selected variables. The top panel characteristics of mentors. The second panel characteristics of treatment and control groups in the Mentoring experiment. The third panel characteristics of treatment and control groups in the Challenges experiment. The bottom panel, applicants to both programs that are not in experimental groups; this group includes subsets of applicants to Mentoring that were not selected by mentors and applicants to Challenges that submitted their application past registration closure. Career changer takes the value of one when the mentor reported in the registration survey having changed occupations. The variable years of tech experience is reported by mentors in the survey. STEM excludes social sciences and architecture. Social science includes business, management, finance, and accounting. The variable years of professional experience is measured based on LinkedIn profiles. Appendix B shows the covariate balance between treatment and control.
5 Effectiveness of Dare IT programs
Analytical framework
Let each Dare IT applicant be characterized by $(X_{i},Z_{i},W_{i}^{M},W_{i}^{C},S_{i},Y_{i})$, where $i$ is the applicant’s index $i=1,..,\mathcal{I}$, $X_{i}\in\mathcal{X}$ is a matrix of characteristics of applicants that are observed by mentors and recorded in data, $\zeta_{i}\in\mathcal{Z}$ are characteristic of applicants that are not observed in the data (though some might be observed by mentors prior to treatment assignment), and $S_{i}\in\mathcal{S}$ is a random variable indicating whether the applicant has been selected by a mentor. [^12] To select mentees, mentors received information from the application form; the important part of the application that mentors observed, which we did not, are short introductory videos recorded by the applicants. Finally, $W_{i}^{P}\in\left\{0,1\right\}$ is the treatment assignment into program $P$ and $Y_{i}\in\mathcal{Y}$ is the observed outcome.
Let $Y_{i}(P)$ and $Y_{i}(0)$ be potential outcomes when applicant $i$ participates in program $P$ and when she does not; $P\in\left\{M,C\right\}$, where $M$ stands for Mentoring and $C$ for Challenges. We are primarily interested in estimating the magnitude of the average treatment effect, defined as $\tau^{P}\coloneqq E\left[Y_{i}(P)-Y_{i}(0)\right]$, and testing whether average treatment effects are higher than zero. We are also interested in comparing the treatment outcomes to one another, although our estimates of such contrasts are noisy due to the limited sample size in each experiment.
Selected sample in Mentoring.
Dare IT asked mentors to choose two applicants they wanted to work with. Thus, selected applicants are not a random draw from the pool. Let $S_{i}$ be a random variable indicating whether a mentor selected the applicant $i$ or not; $S_{i}=1$ when the applicant has been selected by a mentor while $S_{i}=0$ otherwise. Only selected applicants were included in the randomization. Thus, we do not observe outcomes under treatment for the non-selected applicants. The average treatment effect of Mentoring that we estimate is defined as
$\tau^{M}\coloneqq E\left[Y_{i}(M)-Y_{i}(0)|S_{i}=1\right].$ (1)
Table 1 provides summary statistics of subjects in the experimental groups of the mentoring program and the non-selected applicants. There are notable differences between the mentoring control group (those selected by mentors ($S_{i}=1$) but not assigned to treatment ($W_{i}^{M}=0$) and the applicant group (where $S_{i}=0$ and thus, by construction, $W_{i}^{M}=0$), most notably in terms of the primary outcome of securing a tech job. In the Challenges program, participants were selected randomly from eligible applicants, where eligibility was determined based on observable characteristics.
5.1 Average treatment effects
In Figure 2, we display the proportions of participants employed in technology jobs over the span of 16 months, post-application to the program. This is illustrated across all experimental groups. [^13] We construct outcomes over time based on the new job’s start date. In some cases, we only observe the year the new job started rather than the specific month. For those observations, we assign December of that year as the starting month. Consequently, the average outcomes at the 4th month after the end of either program are lower than the average of tech job 4 months across the experimental groups. In the left figure, we consider three main groups: participants of Mentoring, participants of Challenges, and all other applicants to Dare IT (we refer to this group as Out of Dare IT). We can see that all groups start at zero, with none of the applicants having a job in technology; over time, the share of subjects that have a tech job in each group increases, with the Mentoring group increasing right from the moment the program started, and achieving over 50% probability of having a job in tech after 8 months. In contrast, the Challenges group follows a similar trajectory to the Out of Dare IT group until the 8th month, which is approximately 3 months after the end of the program, but then the gap between those two groups increases to almost 20 percentage points.
The difference in average outcomes across the three groups appears to be substantial, particularly in the case of Mentoring. However, the subjects in the Mentoring group were selected by mentors; thus, the program’s impact on the probability of getting a tech job might be compounded by the influence of the non-random selection of the sample. In the right figure, we decompose the effect of selection from the impact of Mentoring. We show the outcomes across applicants selected by mentors and randomized into Mentoring, selected, but randomized into control, and those that were not selected (the non-selected group consists of the non-selected applicants to Mentoring, and the control group, and late applicants to Challenges). We observe that in the 16th month, the probability of having a job in technology differs by over 20 percentage points between selected and not-treated and non-selected applicants, and it is substantially larger than the difference between the two groups that were selected but randomized into treatment or control. On average, mentors selected candidates that were more likely to get a job in technology. This observation highlights that, in this context, the ability to carry out a randomized trial is crucial for obtaining reliable estimates of the program’s effectiveness.
Figure 2: Mean outcome per group over time.
Note: Share of subjects with a tech job per group across time. The month zero is the month of application to Dare IT. In the left figure, we group subjects according to the experimental groups: the treatment group in Mentoring (in green), the treatment group in Challenges (in red), and Out of Dare IT (in blue), where the last group consists of the applicants in control groups, non-selected applicants to Mentoring, and late applicants to Challenges. In the right figure, we group applicants into those selected and those not selected by mentors. "Selected and treated" group (in green) shows the outcomes of the treatment group in Mentoring, "Selected but not treated" group (in blue) is the control group of Mentoring "Not selected" group (in red) includes applicants to Mentoring that were not selected by mentors, both experimental groups of Challenges, and late applicants to Challenges.
In Table 2, we present the estimates of the average treatment effects for both the Mentoring and Challenges programs. Our analysis covers both short-term and longer-term outcomes: securing a job in tech within 4 months and 12 months after the programs, respectively. [^14] Tech job 4 months refers to 4 months after the end of a program which corresponds to approximately 8 months after application; analogously for Tech job 12 months. Additionally, we assess the influence of the programs on the acquisition of any new job within 4 months following the end of a program. For the short-term outcomes, we employ the difference-in-means estimator. [^15] Given the paired design of the mentoring experiment, we estimate variance as $\hat{V}(\hat{\tau}^{M})=\frac{1}{J\times(J-1)}\sum_{g=1}^{J}(\hat{\tau}^{M}_{j}-\hat{\tau}^{M})^{2}.$ (2) Athey and Imbens (2017) show that this is a conservative estimator of variance in a paired experiment. In the case of Challenges, we use the Splawa-Neyman estimator of variance. When estimating the longer-term effect, we utilize the fact that we observe monthly outcomes to estimate a Cox proportional hazard model (Cox, 1972). In Appendix C, we provide an estimate for short-term outcomes using the Augmented Inverse Propensity Weighting estimator (Robins et al., 1994; Athey et al., 2019). Furthermore, Appendix D contains alternative specifications of the Cox proportional hazard model, and the results align with those shown in Table 2. The test results of proportionality of hazard rates can be found in Appendix G.
Table 2: Average treatment effect estimates.
Mentoring Challenges
Difference-in-means Proportional hazard Difference-in-means Proportional hazard
New job Tech job 4 months Tech job 12 months New job Tech job 4 months Tech job 12 months
ATE 0.045 0.129 0.126 0.040 0.089 0.071
(0.05) (0.05) (0.025) (0.05) (0.04) (0.02)
ATE % baseline 10.749 43.941 36.887 13.636 45.306 46.972
(14.23) (17.09) (7.34) (12.00) (14.07) (10.01)
No. of obs. treatment 152 152 2265 183 183 2490
No. of obs. control 147 147 2205 225 225 3330
Note: The estimates of the average treatment effects of Mentoring and Challenges on the probability of finding a new job, a new job in technology within 4 months, and within 12 months. The first row shows estimates of the average treatment effect using the difference-in-means estimator in the case of new job and tech job 4 months (see Appendix C for AIPW estimates). The average treatment effect on Tech job 12 months is estimated using Cox proportional hazard model (we present estimates from a model with no covariates, see Appendix D for Cox proportional models with covariates and with time-varrying treatment effects). The second row presents the estimates from the first row divided by the mean outcome in the control group. Rows three and four present the number of observations per experimental group. The first, second, and third columns present the results of Mentoring intervention. Columns four, five, and six show the results from the Challenges program. Standard errors are presented below the average treatment effect estimates. The variance of the difference-in-means estimates of the average treatment effect of Mentoring are estimated using Equation 2. The variance of the difference-in-means estimate of the Challenges intervention is estimated using the classical Neyman-Splawa estimator. Standard errors of the estimates from Cox proportional hazard model are estimated using the partial likelihood function (coxph implementation in R).
We find that participation in the Mentoring program significantly increases the probability of getting a job in the tech industry, by 13 percentage points within both 4 and 12 months after the program’s conclusion. The short-term effect’s estimate has a standard error of 5 percentage points, whereas the longer-term effect’s standard error is reduced to 2.5 percentage points, thanks to the proportional hazard model’s functional form assumption and our time-series data. The Challenges program also significantly enhances the likelihood of securing a tech job, although its impact is somewhat less pronounced than that of the Mentoring program. We estimate that participation in Challenges increases the chance of securing a tech job by 8.9 percentage points within 4 months (with a standard error of 4), and by 7.1 percentage points within 12 months (with a standard error of 2).
In the second row of Table 2, we express the average treatment effects as a percentage of the control group’s mean outcome. As previously noted, mentor selection leads to higher average outcomes in the Mentoring control group compared to the Challenges control group. As a result, when expressed in percentage terms, the impact of both programs appears quite similar. Participation in Mentoring increases the odds of securing a tech job within 12 months by 37%, whereas in Challenges by 47%. Importantly, we find no evidence of the treatment effect diminishing over time when comparing the programs’ effectiveness 4 and 12 months after their conclusion.
Finally, columns 1 and 4 of Table 2 display the estimated impact of participating in Dare IT programs on the probability of acquiring any new job within 4 months post-graduation. We do not observe any statistically significant difference in this outcome between treatment and control. Therefore, our results suggest that the treatment and control groups of both experiments managed to secure new jobs at similar rates. However, those in the treatment groups were more likely to find employment in the technology sector.
In Appendix F, we show the results based on the outcomes survey. We find that the survey respondents in both treatment groups report a higher number of job offers and higher salaries than the control group. The differences between treatment and control are greater in the case of Mentoring than Challenges.
5.2 Heterogeneous treatment effects
We explore heterogeneity in treatment effects by conducting an analysis of treatment effects across different subgroups using the Cox proportional hazard model. We categorize applicants to the Mentoring and Challenges programs into non-overlapping groups based on four different types of characteristics: STEM education, length of professional experience, type of residence, and age. For each of these categories, we estimate a hazard model that includes treatment and subgroup indicators along with their interaction. The results of this analysis are presented in Table 3. It’s important to note that applicants in each of the non-overlapping groups may have other varying characteristics; to account for this, we include a full set of controls and interactions in Appendix I in Table 10. Furthermore, here we present only selected subgroups; Appendix I presents all the interaction terms.
Table 3: Heterogeneity in treatment effects: Mentoring vs. Challenges.
Mentoring Challenges
Baseline CATE Baseline CATE Diff. Mentoring - Challenges
STEM degree
Yes 0.437 (0.02) 0.030 (0.03) 0.191 (0.02) 0.030 (0.03) -0.001 (0.04)
No 0.239 (0.02) 0.232 (0.03) 0.135 (0.01) 0.086 (0.02) 0.146 (0.03)
Difference subgroups 0.199 (0.03) -0.202 (0.05) 0.055 (0.02) -0.055 (0.03) -0.147 (0.05)
Length of professional experience
Less than 5 years 0.429 (0.03) 0.008 (0.04) 0.119 (0.01) 0.136 (0.03) -0.128 (0.05)
More than 5 years 0.344 (0.02) 0.142 (0.03) 0.178 (0.01) 0.021 (0.02) 0.121 (0.03)
Difference subgroups -0.085 (0.03) 0.134 (0.05) 0.059 (0.02) -0.115 (0.03) 0.248 (0.06)
Type of residence
Small town 0.307 (0.03) 0.285 (0.05) 0.161 (0.02) 0.160 (0.03) 0.126 (0.06)
Warsaw 0.353 (0.02) 0.075 (0.03) 0.148 (0.01) 0.051 (0.02) 0.023 (0.03)
Difference subgroups -0.046 (0.03) 0.211 (0.05) 0.013 (0.02) 0.109 (0.04) 0.102 (0.07)
Age
Over 30 years old 0.261 (0.02) 0.125 (0.03) 0.130 (0.01) 0.114 (0.02) 0.011 (0.04)
Under 30 years old 0.428 (0.02) 0.117 (0.04) 0.178 (0.01) 0.025 (0.02) 0.092 (0.04)
Difference subgroups -0.167 (0.03) 0.008 (0.05) -0.049 (0.02) 0.089 (0.03) 0.011 (0.05)
Note: Estimates of average treatment effects conditioned on subgroups in Mentoring and Challenges. The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazard models, which include treatment indicator, subgroup indicator, and interaction. The first and third columns show the subgroup’s average outcome within control subjects. The second and fourth columns show the difference between treatment and control in the subgroup. The last column shows the difference between Mentoring and Challenges of the treatment effects. Standard errors are in parentheses.
First, we observe that participants with a STEM degree have higher baseline outcomes than those without a STEM degree. Both programs, Mentoring and Challenges, are more effective for participants without STEM degrees, bridging the outcome gap across these two groups. Specifically, in the case of Mentoring, we find that the treatment entirely eliminates the STEM outcomes gap. When comparing the two programs, we find no statistically significant difference for participants with STEM degrees, while for those without STEM degrees, Mentoring leads to nearly 15 percentage points higher treatment effects than Challenges. Second, when focusing on the subgroups with more or less than 5 years of professional experience, we find that Challenges is more effective for participants starting their professional career, with nearly 13 percentage points higher treatment effects in this subgroup. Conversely, for participants with longer professional experiences, Mentoring appears more effective. Third, we find that participants from smaller towns benefit more from both programs. However, this difference is particularly pronounced in the case of Mentoring. Participants from smaller towns have lower baselines than participants from Warsaw, but after the treatment, the gap is entirely closed. In fact, Mentoring participants from smaller towns achieve, on average, higher outcomes than Mentoring participants from Warsaw. Finally, participants below 30 years old have higher baselines than those over 30 years old. However, both programs help close the gap between the two age groups by having higher treatment effects for older participants.
Additional characteristics from Mentoring.
In the case of Mentoring, we observe several additional characteristics; these come from the Mentoring registration survey, which was more extensive than the Challenges one, and we also observe from characteristics of mentors. In Table 4, we present additional results documenting heterogeneity in treatment effects along those additional dimensions.
Table 4: Heterogeneity in treatment effects: Mentoring.
Variable value is 1 Variable value is 0
Baseline CATE Baseline CATE Diff. Subgroups
Mentee characteristics
Mother 0.348 (0.02) 0.288 (0.04) 0.398 (0.02) 0.068 (0.03) 0.220 (0.05)
Applied before 0.384 (0.03) 0.141 (0.05) 0.330 (0.02) 0.121 (0.03) 0.020 (0.05)
Mentor characteristics
Managerial exp. 0.285 (0.02) 0.223 (0.03) 0.414 (0.02) 0.012 (0.03) 0.211 (0.05)
Long exp. tech 0.288 (0.02) 0.226 (0.04) 0.389 (0.02) 0.052 (0.04) 0.174 (0.05)
1st time mentor 0.354 (0.02) 0.131 (0.03) 0.342 (0.02) 0.112 (0.03) 0.019 (0.05)
Note: Estimates of average treatment effects conditioned on subgroups in Mentoring. The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazard models, which include treatment indicator, subgroup indicator, and interaction. Applied before takes the value of 1 if the mentee has applied to the program in the past, but did not get selected into the program, and reapplied, and the value of zero if this is the first time applying to the program. Long exp. in tech takes the value of 1 if the mentor worked for over 5 years in the technology sector (the median value), and zero otherwise; 1st time mentor takes the value of 1 if this is the first time the mentor provides formalized mentoring service, and zero otherwise. Columns 2 and 3 describe the values for subgroups where the subgroup indicator takes the value 1, while columns 4 and 5 when the indicator takes the value of 0. The last column is the difference in the effectiveness of Mentoring across the two subgroups. Columns 1 and 3 are average outcomes in control groups, and columns 2 and 4 are the estimates of the average treatment effect. Standard errors are in parentheses.
First, we observe that the Mentoring program is particularly effective for mothers. Mothers who were selected by mentors but randomized into the control group had a baseline probability of 35% of finding a job in technology, compared to 40% for those without children. However, the average treatment effect for mothers is significantly larger at 29 percentage points, compared to just 7 percentage points for those without children. As a result, Mentoring not only enables mothers to catch up with applicants without children but even surpass them in the probability of getting a job in technology. In contrast, we do not observe a high difference in the effectiveness of Mentoring between first-time applicants and those who applied in previous editions.
Turning our attention to the characteristics of mentors, we find that mentors with managerial experience and those with a long history in the technology sector (over 5 years) tend to select, on average, less promising candidates, as indicated by baseline outcome probabilities of 29% for both groups, compared to the baseline of 41% for mentors without managerial experience and 39% for those with short tenure in tech. However, these mentors with managerial experience and long tech sector tenure are highly effective, with their mentees achieving a probability of over 50% of securing a job in technology. Interestingly, we also find that first-time mentors are as effective as mentors who have had previous mentoring experience. These findings are consistent with the notion that effective mentors leverage their knowledge, skills, and relationships accumulated over years of working in the technology sector, rather than improving their mentoring capabilities through the practice of mentoring.
Patterns of treatment effects heterogeneity are consistent with some of the mechanisms through which the two programs might impact employability that we hypothesized in Section 3. Participants from smaller towns and those without family or friends working in tech (see Table 10) might have less extensive professional networks in the technology sector, and mentors with longer experience in tech are likely to have richer networks. Thus, the finding that participants with these characteristics benefit more from Mentoring is consistent with the mechanisms that Mentoring helps mentees by bolstering their professional network and enhancing their capacity to utilize these networks effectively. As for Challenges, we hypothesized that the program would equip participants with credible signals of their technical skills. This mechanism is consistent with the finding that participants without STEM degrees and those with shorter professional experience, thus fewer other signals on their resumes, exhibit larger benefits.
6 Counterfactual program assignment policies
The number of spots in each program is limited. The capacity of Mentoring is constrained by the availability of mentors, while in Challenges, we capped spots to 300 to accommodate potential operational complexities and logistical issues. Considering the substantial treatment effects of both programs, one might posit that relaxing these capacity constraints will create significant value. Nonetheless, capacity constraints will likely remain a long-term issue for the mentoring program and may persist, albeit to a lesser degree, for Challenges. Consequently, Dare IT will continue being faced with the question of which applicants to accept for each program. With heterogeneous treatment effects, admission policies favoring applicants with high predicted benefits could enhance Dare IT programs’ impact. This section aims to quantify such potential gains.
We carry out three analyses: (i.) we estimate gains from the introduction of a policy that, for given capacity levels in each program, assigns applicants to programs to maximize the sum of predicted treatment effects across the two programs; (ii.) we compare the overall impact of Mentoring when mentors select most promising candidates to a case when they select to maximize expected benefits; ((iii). finally, we evaluate several simple policies based on heuristics.
An assignment policy is a mapping from applicants’ characteristics $X\in\mathcal{X}$ and programs capacity levels $Q=(Q^{M},Q^{C})\in\mathcal{Q}$ to assignments: $\pi:\mathcal{X}\times\mathcal{Q}\rightarrow\mathcal{P}$, where $\mathcal{P}\equiv\left\{M,C,O\right\}$ and $M$ stands for Mentoring, $C$ for Challenges, and $O$ for Out of Dare IT. $\mathcal{Q}$ is the grid of possible capacity levels under consideration. Policy estimation and evaluation are carried out separately. The objective of policy estimation is to obtain a mapping from user characteristics and programs’ capacity level to program assignments’. This mapping $\hat{\pi}$ is estimated in a training set. Across all analyses, we obtain training and testing sets by randomly splitting data at the applicant level with equal probability. Next, to evaluate a policy, we proceed in two steps: first, we apply the estimated mapping to the testing set, where for each $i$ in the test set and each $Q\in\mathcal{Q}$, we obtain the counterfactual assignment of applicants to programs $\hat{\pi}(X_{i},Q)\in\mathcal{P}$. Second, for each $Q\in\mathcal{Q}$, we construct the AIPW estimates of the “policy value,” that is, the value of the outcome when treatments are assigned according to $\pi(X_{i},Q)\in\mathcal{P}$. We do that using cross-fitting in the testing sample, i.e., we estimate new outcome and propensity models and obtain AIPW estimates of the policy values in held-out folds (we use 100 repetitions with 10 folds and consider the median outcome across repetitions) [^16] We follow the approach suggested in Chernozhukov et al. (2017).. Throughout this section, we use Cox proportional hazard model with the full set of interactions as the outcome model and the logit model as the propensity model. Appendix J provides details of the two models and of the cross-fitting procedure.
6.1 Optimal assignment policies
Let $z_{ip}\in\left\{0,1\right\}$ be a variable indicating whether applicant $i$ is assigned to program $p\in P$, and $z_{ip}^{*}$ be the arguments maximizing the sum of treatment effects given by Equation 3:
$\max_{z_{ip}}\sum_{i=1}^{I}\sum_{p\in\mathcal{P}}z_{ip}\tau_{i}^{p}\text{ s.t. }\sum_{i=1}^{I}z_{ip}\leq Q^{p}\thickspace\forall p\text{ \& }\sum_{p\in\mathcal{P}}z_{ip}=1\thickspace\forall i,$ (3)
where $\tau_{i}^{p}=\mathbb{E}\left[Y(p)-Y(O)\right]$. The first constraint ensures that the capacity limits are not violated, and the second one is that each applicant is assigned to one program. A policy that induces allocation $z_{ip}^{*}$ is what we call optimal policy. In Algorithm 1 in Appendix K, we provide further details of the estimation of optimal policy using training data, where we let $\hat{\pi}$ denote the estimate.
We estimate the optimal policy at three capacity levels: first, with the existing capacity (allowing 13% of all applicants into Mentoring and 15% into Challenges) – Optimal 1, second, increasing the capacity of Challenges to accommodate 50% of applicants – Optimal 2, and third, a scenario where each program can admit all applicants – Optimal 3. We compare these optimal policies against several alternatives: a policy accepting all applicants to Mentoring, a policy enrolling all applicants in Challenges (both these policies have increased capacities to accept all applicants), and the existing allocation policy (Status quo). We use the average treatment effect as the metric to compare policies. The results are presented in Figure 3.
Figure 3: Evaluation of counterfactual policies.
Note: Estimates of values of policies. The policy value is the mean probability of having a tech job across all participants and is measured by the mean of AIPW scores under allocation induced by each policy across all participants. Standard errors clustered at a participant level. Whiskers show confidence intervals. Optimal 1 has the existing capacity levels. Optimal 2 increases the capacity of Challenges to 50%, and Optimal 3 has no capacity limits.
First, we find that the optimal policy, with existing capacity levels, markedly improves the average treatment effect when compared to the Status quo, with an increase from 12 to 22 percentage points. Furthermore, this policy is also substantially more effective than policies that assign all applicants to Mentoring or to Challenges. This result follows from the high degree of treatment effect heterogeneity across both programs and shows that a targeting policy can substantially augment Dare IT’s effectiveness. Second, we find that the policy expanding the number of Challenges spots to 50% of applicants (Optimal 2) increases the average treatment effect to 17 percentage points and Optimal 3 to 15 percentage points. Such an increment highlights that the there are sizeable subpopulations amongst Dare IT applicants that have higher treatment effects. Moreover, we find only a small and statistically insignificant difference in policy value when contrasting Optimal 2 to the Optimal 3 policy.
In Table 5, we present the gains from optimal policy per assignment group. A program assignment group consists of all applicants that under a policy are assigned to this program, i.e., Mentoring assignment group are all applicants that are assigned to Mentoring. We are interested in comparing the benefits from the participation in each program of each of these groups under. We find that the group allocated to Mentoring substantially gains from their allocation, yielding an ATE of 0.199 (S.E. 0.017). This group would also substantially benefit from Challenges with the average treatment effect of 0.276 (S.E. 0.019). The difference between assigning them to Mentoring over Challenges is significant -0.077 (S.E. 0.021). However, the group assigned to Challenges benefits from this assignment even more. The average treatment effect of Challenges amounts to 0.244 (S.E. 0.017) compared to 0.109 (S.E. 0.012) of Mentoring. Lastly, the non-participating group has substantially lower benefits from assignment to Mentoring or to Challenges, average treatment effects of 0.101 (S.E. 0.006) and 0.073 (S.E. 0.003) respectively.
Table 5: Gains from the optimal assignment
Mentoring group Challenges group Nothing group
Share 0.15 0.50 0.35
ATE under assignment to:
Optimal 0.199 (0.017) 0.244 (0.017)
Mentoring 0.199 (0.017) 0.109 (0.012) 0.101 (0.006)
Challenges 0.276 (0.019) 0.244 (0.017) 0.073 (0.003)
Difference in ATE between assignments:
Mentoring - Challenges -0.077 (0.021) -0.135 (0.025) 0.028 (0.006)
Note: Gains from the optimal assignment by assignment group from Optimal 1 - 13% of applicants get be admitted to Mentoring and 15% to Challenges. Groups are defined as all applicants assigned to the specific program. The columns describe the outcomes of groups under various allocations - the group membership is determined using Algorithm 1 and is constant across the comparisons here, the allocations change. The first row is the share of applicants assigned to a group. Rows two to four describe the average treatment effect of the group under various assignments. The 5th row is the average treatment effect under Mentoring less the average treatment effect under Challenges. Standard errors are next to each value in parentheses and are clustered at the applicant level.
When comparing the three policies based on optimal allocation, it’s evident that the current capacity levels limit Dare IT’s impact but that additional spots do not create more value beyond a certain capacity. To further unpack this, in Figure 4, we analyze the value of optimal policy under a sequence of capacity values and compare with random assignment to either Mentoring or Challenges. Purple dots in Figure 4, show the average probability of getting a job in technology across all applicants (including those not participating in Dare IT programs), for the total capacity of the two programs ranging from 0 to 100% (in all cases, the two programs have the same capacity, so for example the total capacity of 50% means that, up to 25% of applicants get be assigned to either program). We compare the outcome under these policies with policies that assign applicants randomly to Mentoring (in yellow) and to Challenges (in red).
Figure 4: The value of targeting at various capacity levels
Note: Estimates of mean outcomes across all applicants at various capacity levels of Dare IT programs and allocation mechanisms. The policy value is the predicted share of applicants getting a tech job. The share of applicants getting a spot refers to the total capacity in both programs. Challenges Random (in orange) and Mentoring Random (in red) assign applicants at random between the program and Out of Dare IT. Targeting policy (in purple) follow the optimal policy approach.
Our analysis indicates that policies randomly admitting applicants to either the Mentoring or Challenges programs lead to consistently lower policy values when compared to the targeted policy across all capacity levels. When capacities are low, targeted policies prioritize applicants with the highest benefits. A meaningful difference in the average probability of getting a job across all applicants quickly emerges between the targeted and random policies; consequently, already at the 25% capacity levels, there is an approximately 4 percentage points gap between targeted and random policies. Interestingly, even when capacity constraints are lifted to accommodate all applicants, the targeted strategy continues to lead to higher outcomes than the random policies based on one program. This discrepancy remains because there of the difference in benefits that applicants receive across the two programs, i.e., there are some applicants that benefit more from Challenges than they do from Mentoring.
6.2 Counterfactual selection policies for mentors
Applicants selected by mentors and randomized to be the control group had a significantly higher likelihood of securing a tech job compared to those who were not selected (see Figure 2). This suggests that mentors tend to prefer highly promising applicants who are likely to secure a tech job irrespective of their participation in Mentoring. In this section, we examine alternative selection strategies that mentors could adopt and their potential implications for the effectiveness of the mentoring program. Specifically, we consider four scenarios: (i) Highest CATE – targeting applicants predicted to have the highest treatment effects, (ii) Most Promising – targeting applicants expected to have the highest outcomes without participating in Mentoring, (iii) Least Promising – selecting applicants expected to have the lowest outcomes if they do not participate, and (iv) Highest Chances if Treated – where mentors select participants that have the highest predicted outcomes if they participate in Mentoring; additionally, we include a case in which mentors select at random (Random Choice) and the observed allocation (Observed) as benchmarks. In each policy, we assume that 36% of applicants participate in Mentoring, which corresponds to the share of applicants that were selected by mentors. [^17] We consider a higher number of participants than the current number of spots in Mentoring to amplify the impact of policies on outcomes across all applicants; the ordering of policies does not depend on the share of applicants admitted to Mentoring as long as this number is constant across policies.
We follow the same off-policy approach as previously. We consider the dataset of all applicants to Mentoring, which consists of 151 individuals who participated in the program, 147 mentors selected but randomized to the control group, and 468 non-selected applicants. We split data into train and test with equal probability. We use the training dataset to estimate the mapping from applicants’ characteristics to assignments. To do that, we train the Cox proportional hazard model in the training data and predict outcomes with and without treatment in the testing data. Then, we use these predictions to determine which applicants get selected under each policy. To evaluate policies, we compute AIPW scores analogously to the optimal policy evaluation (Appendix J.1): using only testing data, we train the propensity model (logit model) and a new outcome model (Cox hazard model) using cross-fitting. We compute 100 cross-fitting rounds and take the median outcome. Table 6 shows the results.
Table 6: Mentors’ selection policies.
Policy Mean ATE Baseline outcome treated group Mean outcome all applicants
Highest CATE 0.178 (0.012) 0.187 (0.008) 0.232 (0.005)
Most Promising 0.062 (0.012) 0.285 (0.009) 0.192 (0.005)
Least Promising 0.127 (0.007) 0.055 (0.005) 0.211 (0.005)
Highest Chances if Treated 0.130 0.012 0.237 (0.008) 0.218 (0.005)
Observed 0.096 (0.012) 0.275 (0.008) 0.205 (0.005)
Random 0.094 (0.011) 0.191 (0.008) 0.203 (0.005)
Note: Outcomes and treatment effects across different mentors’ selection rules. Mean ATE is estimated using AIPW estimator with cross-fitting. Baseline outcome treated group is the predicted outcome without treatment for the group that receives treatment in the considered policy. Mean outcome all applicants corresponds to the average probability of getting a job in tech across all applicants, where the group assigned to treatment under considered policy receives treatment, and the remaining applicants do not participate. Standard errors clustered at the applicant level.
We find that prioritizing applicants with the highest anticipated treatment benefits results in higher average outcomes for all applicants and a more significant treatment effect than other strategies. Using the Highest CATE policy, 23% of applicants secure a tech job, a noticeable improvement over the 20% success rate under the Random or Observed policies. The average treatment effect under Highest CATE is an 18 percentage points increase in the probability of having a tech job for the treated group. Interestingly, targeting applicants with the highest chances of having a job if not treated (Most Promising) leads to treatment effects that are lower than selecting applicants randomly. Moreover, a policy aimed at the seemingly less promising applicants (Least Promising) — which goes against current mentor preferences — produces an average treatment effect of 13 percentage points. This results in an average outcome for all applicants surpassing those achieved by the Most Promising and Random policies. The policy of targetting applicants that have the highest chances of getting a tech job if treated (Highest Chances if Treated) results in a 13 percentage points average treatment effect, which is similar to Least Promising.
74% of applicants assigned to Mentoring under Highest CATE are also assigned to the program under Highest Chances if Treated, which is expected considering the fact that these two policies are most promising. However, there is a low overlap between Highest CATE and Least Promising only 7%. Furthermore, there is a substantial overlap between the observed selection by mentors and the Highest CATE policy – 54%, but a very low one between Observed and Least Promising, only 4%. These observations suggest that a group of applicants has low chances of securing a tech job when not treated and are systematically overlooked by mentors, even though their benefits from the program would be high. We analyze the overlaps and the characteristics of selected candidates in Appendix M.
6.3 Simple priority rules
Optimal assignment policy increases the average probability of getting a tech job across all applicants compared to the status quo. However, a non-parametric policy requiring constraint integer optimization is difficult to implement in Dare IT’s context. In this section, we examine whether a simple priority rule based on heuristics can enhance the average effectiveness of Dare IT’s programs.
Consider the admission policy wherein Dare IT first screens participants based on their priority level and then randomly assigns them to either the Mentoring or Challenges program (with the probabilities proportional to the number of spots in either of the programs). We propose the following priority groups: Priority 1 – Applicants who are from small towns, do not have a degree in a STEM field, and do not have a master’s or postgraduate degree; Priority 2 – Applicants who are from small towns, do not have a STEM degree, but have a master’s or postgraduate degree; Priority 3 – Applicants who are not from small towns and do not have a STEM degree; Priority 4 – Applicants who have a STEM degree.
The proposed priority groups are non-overlapping and together encompass all applicants. Thus, under this proposed admission policy, Dare IT would first consider the applicants in the first group. If there is more capacity after admitting the applicants in the first group, they will move to the second group, and so on. We are interested in evaluating whether the groups of applicants with high priority have higher predicted treatment effects than groups with lower priority levels and whether the high-priority groups are large enough to fill in all the existing spots in the two programs.
Given that this policy is defined using heuristics instead of data, we do not split the dataset for policy estimation and separate evaluation. Instead, we use the entire dataset to evaluate this pre-defined policy. To estimate the value of the policy, we consider both experimental populations and the remaining applicants. For each applicant, we estimate outcomes under three conditions: Mentoring, Challenges, and Out of Dare IT. We do that using cross-fitting as described in Appendix J.1. The outcome model is the proportional hazard Cox model, and the propensity model is the multinomial logit model; we combine them to obtain AIPW estimates. Admitted applicants are assigned with probability 45% to Mentoring and 55% to Challenges (reflecting the proportions of treated subjects in the dataset). Once we have the allocation, we compute the average treatment effect per group using the AIPW predictions. Figure 5 shows the results.
Figure 5: Evaluation of simple policies.
Note: Estimates of the average treatment effect per priority group; AIPW estimates obtained using cross-fitting. The treatment effect is the percentage point increase in the probability of a tech job. The red dotted line is the average treatment effect across the entire population. Whiskers around point estimates represent confidence intervals. Standard errors are clustered at the participant level.
We find that groups 1 and 2 have predicted average treatment effects that are higher than the average across the entire population of applicants (red scatter line) and higher than groups 3 and 4. Furthermore, group 1 encompasses 14% of applicants and group 2 10%. The combined capacity of Mentoring and Challenges is 28%. Thus, the two groups can almost fill in all the spots in the two programs. We find that this simple priority rule can increase Dare IT’s effectiveness.
7 Conclusion
Changing labor markets push workers to want to transition into high-growth sectors, and facilitating these transitions is an important goal for labor policy. However, effectively designing such policies requires principled evaluation of various training programs in terms of their overall impact and their heterogeneity. In this paper, we provide evidence of the high effectiveness of two programs supporting women transitioning into the technology sector. This way we identify a class of interventions that can effectively narrow the gender gap in technology.
We show that a mentoring program and an online training program called Challenges aimed at developing labor market signals increase the probability of having a job in the technology sector by more than 40% four months after the end of the program. Furthermore, the online training program developed during this research project is cheap to operate, at approximately \$15 per person, and suitable for scaling up, suggesting that it can have an impact extending beyond the current setting.
Our findings reveal strong patterns of heterogeneity in treatment effects. The benefits derived from the two programs depend on a number of participants’ characteristics including the educational background, length of professional experience, type of residence, and age. For example, we find that older participants that come from small towns have particularly high treatment effects in both programs, especially in Mentoring. Furthermore, certain groups benefit distinctly from one program over the other. Specifically, participants with short professional experience benefit from Challenges, while those with over five years of experience from Mentoring.
For the foreseeable future, the demand for the two programs, particularly for Mentoring will exceed the number of available spots, and Dare IT will grapple with the question of whom to admit. We propose an admission policy that maximizes Dare IT’s impact, by prioritizing applicants with the highest predicted treatment effects. Consistent with the high degree of heterogeneity in the effectiveness of the two programs we find that such targeted admission policies can substantially increase Dare IT’s effectiveness. Specifically, we estimate a seven percentage points increase in the average probability of getting a job in the technology sector across all applicants. Notably, such targeted policy utilizing the current combined 28% capacity of both programs, yields outcomes comparable to policies that accommodate all applicants into either Mentoring or Challenges. This showcases that in this context targeting admission has a very high potential impact.
A noteworthy aspect of Mentoring is that mentors have the autonomy to choose their mentees. To facilitate randomized access for our experiment, we asked mentors to shortlist two candidates and we randomized within these pairs. This resulted in three categories: applicants not chosen by mentors, those shortlisted but placed in the control group through randomization, and those shortlisted and included in the treatment group. Importantly, all applicants underwent an assessment of their technical skills before being presented to mentors. By the end of our observation period, there was a 20 percentage point difference in securing tech jobs between those selected but placed in control and the non-selected applicants, with the latter securing job in technology at the probability of 22%. This underscores that mentors lean towards candidates they deem more promising. While this approach boosts the success rate of Mentoring, reflected in the proportion of participants landing tech jobs, it does not necessarily maximize impact since many of these candidates might have secured tech roles even without Mentoring. Through an off-policy evaluation, we explored potential alternative mentor selection strategies. Our findings suggest that mentors focusing on participants with the greatest treatment effects, rather than those already poised for success, can significantly amplify the efficacy of Mentoring. We estimate that when mentors select 40% of participants based on the highest treatment effect, the share of applicants eventually finding a job in technology is 31% compared to 21% when they choose participants that have the highest outcomes if treated, and 24% when they select participants at random.
This study comes with certain limitations worth noting. First, the limited participant size in both programs led to wide confidence intervals in several estimates, leaving room for uncertainty regarding the exact impacts of Mentoring and Challenges. Secondly, we could not pinpoint the exact mechanisms driving the outcomes for both programs. Although our findings suggest that Mentoring likely enhances participants’ professional networks and teaches how to utilize them, and Challenges supports participants in the development of a credible signal of their skills, a more in-depth exploration is required to understand the specific mechanisms through which the training provided by both programs impacts employability. Such an analysis could also help scale up Mentoring by identifying the crucial aspects of the program and directing Dare IT resources to them. Lastly, the participants from the first edition of Challenges may not be representative of those in later editions. Early adopters of such programs might experience greater benefits. To validate the program’s continued efficacy, subsequent evaluations are essential.
Appendix
Appendix A Challenges program
The Challenges program is designed for the same population as the Mentoring program: individuals who possess technical competencies but have not yet secured employment in the technology sector. The primary aim of the Challenges program is to assist these individuals in developing credible credentials that can effectively signal their skill set. In its first edition, the Challenges program provided two distinct tracks: User Experience Design (UX) and Front-end Development in React (front-end). The selection of these specific tracks was influenced by several determinants. Primarily, there was a significant volume of applications for the Mentoring program related to these two domains. Furthermore, Dare IT possessed both internal expertise and established partnerships in these areas. Lastly, a pronounced market demand existed for professionals in these specializations in Poland.
To achieve the program’s scalability goal and streamline the application process, unlike the Mentoring program, the review of the minimum technical skills during application was based solely on applicants’ declaration. The registration website explicitly mentioned the necessity for an advanced understanding of Figma for those interested in UX design, and React for those leaning towards front-end engineering as prerequisites for the Challenges program. The applicants had to confirm that they indeed have the minimum skill level. Additionally, English proficiency was mandated due to the inclusion of course content delivered in English.
The Challenges program spanned 16 weeks. Initially, the first two weeks focused on an introductory event and a refresher course in Figma and React. The concluding two weeks were allocated for the refinement of the final projects, HR workshops, and a closing event. Central to the program, the intermediate 12 weeks were primarily aimed at the progression of an individual project. Within both specializations, experts from partner organizations, specifically Flying Bisons and Grover – Polish tech companies, were asked to prepare six tasks. These tasks, inspired by their recent undertakings, encapsulated the comprehensive procedure of designing a mobile application for the UX domain and detailed the facets of front-end engineering for the respective specialization.
The program is based on two-week cycles culminating on Thursdays when submissions are expected and news assignments are posted. Every subsequent Friday, post-submission, an evaluative session with a subject-matter specialist takes place. During this session, a selection of submissions undergoes review, identifying prevalent errors and underscoring industry best practices. This grants participants the opportunity to amend and refine their submissions from the previous week. The domain experts, either mentors from Dare IT or other professionals with an appropriate level of seniority, carry out the reviews. Each assignment has been designed to take approximately eight hours of time to finalize. As the program approaches its conclusion, participants engage with a Human Resources specialist. This guidance is primarily aimed at instructing participants on leveraging their newly-acquired qualifications, both as skill indicators in job applications and as a resource during subsequent interviews.
In the first week of the UX path, participants conceptualize their project. They are given the flexibility to either devise their own application idea or choose from two provided project themes, followed by a detailed analysis of market competitors and benchmarks. The subsequent phases guide participants through a structured journey from ideation to delivery. They familiarize themselves with tools like Personas, User Journey Maps, and User Stories, ensuring a systematic organization of the information gathered. As the program advances, participants delve deeper into content hierarchy, Information Architecture, and user pathways, culminating in the creation of UX wireframes. The latter weeks emphasize usability testing, where participants craft interactive prototypes, refining them based on feedback. Crucially, they also submit project presentations and receive feedback on them.
In the case of the front-end path, in the first weeks, participants created reusable UI components, often termed Atoms. This initial phase emphasized the foundational importance of efficient design. In the subsequent weeks, attendees delved into topics of server communication: creating APIs, handling different requests, and DELETE queries. Following these modules, the middle segment of the program trained participants in creating models and form-based data collection. Later, the curriculum shifted towards data visualization using charts and the integration of real-time user feedback systems, e.g., notifications. These modules were designed to underscore the advantages of adept data presentation and heightened user experience. Finally, the program introduced participants to TypeScript, showcasing its emerging prevalence in modern web applications.
In addition to the assignments and review sessions, participants were granted access to the exclusive Dare IT community on the Slack platform. During the kick-off meeting, the participants were recommended to collaborate in teams to tackle the assignments. A dedicated process was created to support collaboration; participants were asked whether they wanted to be matched with a group. If they were interested, they received access to a group’s channel. [^18] The emphasis on collaborative efforts and peer support was further amplified in the ensuing editions of the Challenges program, where matching events were organized. Within each path’s Slack channel, distinct discussion forums were established, allowing participants to pose queries, circulate resources, and engage in discourse about assignment components. Activity within this Slack community was consistently high, with frequent discussions across multiple channels.
The first edition of Challenges has attracted a lot of popularity amongst the participants, applicants that we could not accommodate due to capacity limits, and the Polish tech community more broadly, thereby galvanizing its further development. Currently, the program is instituted on a trimestral cycle, extending its offerings to encompass Manual Testing, Automated Testing in Python, Cloud Computing, and User Interface Design. Furthermore, in the summer of 2022, the program was adapted to cater to Ukrainian refugees in Poland. This edition was delivered in Ukrainian, incorporating bespoke modules tailored to guide foreigners in navigating the Polish job market.
Appendix B Covariate balance
In Figure 6 we show the balance of covariates in the mentoring experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances. The only difference that is statistically significant is the graduate degree, where 66% of treated subjects have a graduate degree and 55% of control (p-value 0.05).
Figure 6: Covariate balance mentoring.
Note: Balance across treatment and control of the main variables. Points correspond to the absolute difference in means between treatment and control divided by the square root of the sum of variances in treatment and control.
In Figure 7 we show the balance of covariates in the Challenges experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances. None of the differences is statistically significant.
Figure 7: Covariate balance Challenges.
Note: Balance across treatment and control of the main variables. Points correspond to the absolute difference in means between treatment and control divided by the square root of the sum of variances in treatment and control.
Appendix C Short-term average treatment effects estimates from the AIPW estimator
In Table 7, we present the estimates of the average treatment effects using the Augmented Inverse Propensity Weighing estimator (Robins et al., 1994). AIPW is a doubly-robust method which means that it adjusts for covariates in the outcome model and the propensity score. We use the grf implementation of the AIPW estimator, as in Athey et al. (2019). We use the full set of covariates as controls. We find that the estimates from the AIPW estimator are very similar to the estimates obtained using the difference-in-means estimator.
Table 7: Average treatment effect estimates.
Mentoring Challenges
Tech job 4 months New job Tech job 4 months New job
ATE AIPW 0.129 0.037 0.088 0.0437
(0.05) (0.06) (0.04) (0.04)
ATE 0.129 0.045 0.089 0.040
(0.05) (0.05) (0.04) (0.05)
Mean treatment 0.421 0.467 0.284 0.333
(0.04) (0.04) (0.03) (0.03)
Mean control 0.293 0.422 0.196 0.293
(0.04) (0.04) (0.03) (0.03)
No. of obs. treatment 152 152 183 183
No. of obs. control 147 147 225 225
Note: The estimates of the average treatment effects of both the mentoring program and Challenges on the probability of finding a new job in technology and finding any new job. The first row shows the estimates using the augmented inverse propensity weighting (AIPW) estimator. The second row shows estimates of the average treatment effect using the difference-in-means estimator. Rows three and four present mean outcomes in treatment and control groups and rows five and six the number of observations per experimental group. The first and second columns present the results of the mentoring intervention on the probability of finding a job in technology and the probability of finding any new job. Columns three and four show the results of the impact of participation in Challenges on the probability of finding a job in technology and any new job. Standard errors are presented below the average treatment effect estimates. The variance of the difference-in-means estimates the average treatment effect of mentoring estimated using equation 2. The variance of the difference-in-means estimate of the Challenges intervention is estimated using the classical Neyman-Splawa estimator. AIPW variance estimated using grf implementation of AIPW estimator (see Athey et al. (2019) section 4 for details).
Appendix D Long term treatment effects
In this, section we present the estimates from two Cox proportional hazard models for both Mentoring and Challenges. First, we estimate a model with a full set of controls and second, we consider a model where the treatment variable has time-varying coefficients and the full set of controls. Table 8 shows the results. We estimate the coefficient of 0.26 in the case of Mentoring, which corresponds to a 29% increase in the probability of having a tech job. We find higher treatment effects in the case of Challenges, a coefficient of 0.37, corresponding to a 45% increase. We also consider the model in which we analyze the impact of participation in Dare IT programs across time - columns 2 and 4 of Table 8. We confirm the observation from Figure 2, treatment effects (especially in the case of Challenges) are realized predominantly after the end of the two programs. Furthermore, we do not observe attenuation of the treatment effect over time.
Table 8: Long-term treatment effects
Dependent variable: tech job
Mentoring Mentoring time-varying Challenges Challenges time-varying
Treated ${}^{***}$ (0.0501) ${}^{***}$ (0.065)
Treated:during program 0.196 (0.130) --0.074 (0.239)
Treated: 4 months after the program ${}^{***}$ (0.071) ${}^{***}$ (0.094)
Treated: 5 months plus after the program ${}^{***}$ (0.082) ${}^{***}$ (0.096)
Full set of controls x x x x
Observations 4,275 4,275 5,490 5,490
Note: Selected coefficient estimate from the proportional hazard models. The first column is the mentoring experiment. The second column is the mentoring experiment with time-varying coefficients. The third column is the challenges experiment. In the last column, the challenges experiment with time-varying coefficients. In Appendix G, we present the Cox proportional hazard model diagnostics.
Appendix E Comparison of salary levels from Glassdoor
We consider two types of jobs as a job in technology:
  1. 1.
    All jobs in tech companies other than finance, regulatory, legal, accounting, and HR, where tech companies include firms in software development, testing, and sales; data analytics; IT services; digital marketing; and online platforms (including peer-to-peer platforms, and online sales).
  2. 2.
    Jobs in non-tech companies that involve software development and testing, IT support, and data analytics. In our context, this category includes jobs in banks and management consulting agencies.
We analyze whether the two categories of jobs pay similar salaries. To do that we searched for salary estimates of all unique job titles that fall into the two categories using Glassdoor.com. We considered only exact matches and salary estimates for jobs in Poland. In total, we have 265 unique salary estimates. We find that the mean monthly salary for tech jobs in tech firms (category 1) is PLN 8634 and for tech jobs in non-tech firms it is PLN 8866. The difference is statistically insignificant (p-value 0.9). Figure 8 shows histograms of of salary estimates.
Figure 8: Histograms of salary estimates from Glassdoor.
Note: Histograms of salary estimates from Glassdoor. In green, we show tech jobs in tech firms (category 1), in blue tech jobs in non-tech firms (category 2).
Appendix F Outcomes measured in surveys
We use data from outcomes surveys to provide additional evidence of the effectiveness of the two programs that give insights into the mechanisms through which participation boosts outcomes. There were three surveys: two main ones carried out approximately four months after the end of each program, and a shorter survey right after the end of the mentoring program from which we obtained information on salaries in jobs found during or right after the mentoring program. As discussed in Section 4, results based on the outcomes survey suffer from a selection problem. Thus, we interpret the findings based on this data as suggestive evidence.
The number of job offers.
Figure 9 shows the number of job offers received by survey respondents broken by experimental groups. We find that subjects in the control group (in green) were more likely to report receiving zero job offers than the subjects in the treated group (in orange); for any non-zero range, there is a higher share of the treatment group rather than the control group. The largest difference between the experimental groups is amongst subjects reporting more than ten job offers.
Figure 9: Reported numbers of job offers: Mentoring.
Note: The number of job offers received by subjects in the mentoring experiment that responded to the survey. Each bar corresponds to the share of respondents from an experimental group. Orange represents the treated group, and green represents the control group.
In Figure 10, we present evidence of the differences in the number of job offers across treatment and control groups of the Challenges experiment. We find results analogous to those from the mentoring program. The lowest category, here less than two job offers, has a higher share of control group respondents than treated group respondents. In contrast, treated subjects are more likely to report a higher number of job offers. The difference between treatment and control is more pronounced in the mentoring program.
Figure 10: Reported numbers of job offers: Challenges.
Note: The number of job offers received by subjects in the Challenges experiment that responded to the survey. Each bar corresponds to the share of respondents from an experimental group. Orange represents the treated group, and green represents the control group.
Salary ranges.
Figure 11 presents the share of respondents to mentoring outcomes surveys declaring a salary within a salary range (ranges were pre-defined in the survey). Orange bars indicate the treatment group, and the green ones are the control group. The right panel is based on the survey carried out right after the end of the program, while the left panel is from the survey conducted four months after the end of the program.
Figure 11: Salary ranges: Mentoring.
Note: Reported salary ranges by subjects who found new jobs after the start of the mentoring program. Orange represents the treated group, and green represents the control group. The length of a bar represent the share of respondents declaring a salary within the range. The right panel survey was conducted right after the end of the program; the left panel survey was conducted four months after the program.
Evidence presented in Figure 11 suggests that subjects in the treated group of the mentoring experiment that found a new job are more likely to have a salary in a high salary range.
In Figure 12, we report analogous results from the Challenges outcome survey. We find that a higher share of the survey respondents in the control reported salary in the lowest range as compared to the treatment group. We also find that the opposite is true for the highest salary range.
Figure 12: Reported salary range Challenges.
Note:Reported salary ranges by subjects who found new jobs after the start of the Challenges experiment. The length of a bar represent the share of respondents declaring a salary within the range. Orange represents the treated group, and green represents the control group.
Negotiation.
Respondents that reported finding a job were asked whether they negotiated the conditions of their employment. Amongst the mentoring group, 33% of treated subjects that found a new job negotiated the terms of the new job, while only 24% of the control group did so. The difference is statistically insignificant.
Appendix G Proportional hazards assumption
The Cox proportional hazards model assumes the hazard rate of an event is a product of a baseline hazard and an exponential function of the covariates. Thus, in the analysis, we assume the effects of different predictors are constant over time. To test that, we correlate Schoenfeld residuals from the baseline Mentoring and Challenges models with months. Table 9 presents the results. We do not find any statistically significant correlations.
Table 9: Proportional hazards assumption test
Mentoring model Challenges model
chisq1 df1 p1 chisq2 df2 p2
Treatment 0.232 1 0.630 1.301 1 0.254
City 0.149 2 0.928 1.863 2 0.394
Degree 1.909 4 0.752 5.832 4 0.212
STEM 0.001 1 0.971 0.002 1 0.968
Prof experience 0.187 1 0.665 1.723 1 0.189
Social science 0.074 1 0.785 - - -
Family friends IT 0.074 1 0.785 - - -
Mother 0.164 1 0.685 - - -
Applied before 0.087 1 0.768 - - -
First time mentor 0.956 1 0.328 - - -
Managerial experience 0.762 1 0.383 - - -
Career changer 0.043 1 0.836 - - -
Years of experience Mentor 0.428 1 0.513 - - -
Age 2.716 3 0.438 - - -
GLOBAL 9.595 20 0.975 13.130 9 0.157
Note: Diagnostics of the proportional hazard model. The first three columns show results for the Mentoring model, and the last three for Challenges. City is a discrete variable with levels: Small Town, Large City, and Warsaw; degree is also a discrete variable with levels at: no education mentioned, non-traditional credential, college degree, masters, and postgraduate. Prof experience is the number of years between the start date of the first employment of program participants and 2023.
Appendix H Bounds on the average treatment effect
In this section, we perform three robustness checks on the average treatment effects. First, we consider the AT bounds method introduced by Lee and Weidner (2021). Second, due to changes in LinkedIn profile URLs we could not collect outcome data from 17 Challenges participants. To obtain bounds on the treatment effects, we consider two extreme cases - none of the 17 subjects found a job in technology, and all of them did. We estimate treatment effects in both cases. Lastly, we had four participants drop out of the Mentoring program because they secured jobs in the tech industry before the start of the program. We establish bounds based on this information, assuming two distinct scenarios: (i) all corresponding control group participants also found jobs in technology, and (ii) none of them did.
Lee and Weidner (2021) proposes bounds on the average treatment effects that are designed to be robust in situations where there are many conditioning variables, or they have many values, or when the overlap condition is violated. We use R implementation of their algorithm (Lee and Weidner, ).
In the case of Mentoring we consider the following variables: Social science, STEM, Family or Friends in IT, Mother, Applied Before, First-time mentor, Managerial Experience, Career changer, Mentor’s Years of Experience, Professional Experience. We estimate the bounds on the average treatment effect to be 0.11 to 0.18 with a confidence interval of 0.03 to 0.26. In the case of Challenges, we consider the following characteristics: social science, STEM, Specialization, Residence, Age, Professional experience, Degree. We obtain bounds of 0.10 to 0.12 with a confidence interval of 0.02 to 0.20.
Next, we address the unavailability of LinkedIn data for 17 Challenges participants. We explore two extreme scenarios: one where all 17 did not secure tech-sector jobs and another where all 17 did. We estimate the average treatment effect in these scenarios using a difference-in-means estimator, without controlling for observable features due to the lack of LinkedIn data. When we assign zero outcomes for the missing subjects, the estimated average treatment effect is 6.6 percentage points, with a confidence interval ranging from -1 to 15 percentage points. Assuming positive outcomes for all missing variables, we estimate the average treatment effect at 15 percentage points, with a confidence interval spanning 7 to 20 percentage points.
Finally, we conduct a robustness check of our evaluation of the Mentoring experiment by examining two scenarios: (i) all four control group subjects also found tech jobs, and (ii) none of them did. We estimate treatment effects using the difference-in-means estimator. In the pessimistic scenario, the estimated treatment effect is 12 percentage points, with a confidence interval of 2 to 23 percentage points. In the optimistic scenario, we estimate the treatment effect at 15 percentage points, with a confidence interval spanning 4 to 26 percentage points.
Appendix I Heterogeneous treatment effects
In Table 10 we show estimates from the Cox model including the full list of interactions.
Table 10: Heterogeneous treatment effects of the interventions.
Dependent variable: tech job
Mentoring Challenges
Treatment ${}^{***}$ (0.265) ${}^{***}$ (0.239)
Social science --0.124 (0.125)
STEM ${}^{***}$ (0.099) ${}^{***}$ (0.111)
Family friends IT ${}^{***}$ (0.092)
Small Town $*{}^{*}$ (0.102) $*{}^{*}$ (0.126)
Warsaw 0.088 (0.094) ${}^{***}$ (0.137)
Mother ${}^{***}$ (0.137)
Applied before ${}^{***}$ (0.096)
First time mentor 0.123 (0.084)
Managerial experience ${}^{***}$ (0.087)
Career changer ${}^{***}$ (0.099)
Years of experience mentor --0.003 (0.013)
Non-traditional credential ${}^{***}$ (0.188) ${}^{***}$ (0.179)
Masters ${}^{***}$ (0.171) 0.218 (0.135)
No education mentioned ${}^{**}$ (0.193) $*{}^{*}$ (0.287)
Postgraduate $*{}^{*}$ (0.188) ${}^{***}$ (0.194)
Age: 25-30 $*{}^{*}$ (0.145) ${}^{**}$ (0.155)
Age 31-35 ${}^{***}$ (0.168) 0.170 (0.171)
Age: 36+ ${}^{**}$ (0.198) --0.144 (0.184)
Prof. experience ${}^{***}$ (0.013) 0.008 (0.011)
Treatment×$\times$×Social science 0.273 (0.179)
Treatment×$\times$×STEM ${}^{***}$ (0.136) ${}^{***}$ (0.160)
Treatment×$\times$×Family friends IT $*{}^{*}$ (0.119)
Treatment×$\times$×Small ${}^{***}$ (0.131) ${}^{***}$ (0.173)
Treatment×$\times$×Warsaw ${}^{***}$ (0.129) ${}^{***}$ (0.182)
Treatment×$\times$×Mother ${}^{***}$ (0.164)
Treatment×$\times$×Applied before --0.155 (0.123)
Treatment×$\times$×First time mentor --0.079 (0.108)
Treatment×$\times$×Managerial experience ${}^{***}$ (0.112)
Treatment×$\times$×Career changer ${}^{***}$ (0.125)
Treatment:$\times$×Years of experience mentor 0.001 (0.016)
Treatment×$\times$ Non-traditional credential ${}^{***}$ (0.235) ${}^{***}$ (0.251)
Treatment×$\times$ masters ${}^{***}$ (0.209) --0.267 (0.192)
Treatment×$\times$ No education mentioned ${}^{***}$ (0.347) --0.139 (0.349)
Treatment×$\times$×postgraduate ${}^{**}$ (0.236) ${}^{***}$ (0.243)
Treatment×$\times$×Age: 25-30 ${}^{***}$ (0.203) ${}^{***}$ (0.209)
Treatment×$\times$×Age:31-35 ${}^{***}$ (0.241) ${}^{***}$ (0.234)
Treatment×$\times$×Age: 36+ 0.273 (0.277) --0.013 (0.235)
Treatment×$\times$×Prof experience ${}^{***}$ (0.016) $*{}^{*}$ (0.015)
Observations 4,440 4,965
[…]
Log Likelihood --11,845.370 --6,329.277
Wald Test ${}^{***}$ (df = 39) ${}^{***}$ (df = 23)
LR Test ${}^{***}$ (df = 39) ${}^{***}$ (df = 23)
Score (Logrank) Test ${}^{***}$ (df = 39) ${}^{***}$ (df = 23)
Note: $*{}^{*}$p$<$0.1; ${}^{**}$p$<$0.05; ${}^{***}$p$<$0.01
Note: Estimates from the Cox proportional hazard model. Outcome variable tech job.
Appendix J Models used in off-policy sections
This section presents further details about the propensity and outcome models used in Section 6 for policy estimation. To estimate both models, we first join the Mentoring, Challenges, and applicants’ datasets. We obtain a dataset of 1314 individuals who applied to Dare IT, out of whom 151 participated in Mentoring and 166 in Challenges; the remaining 997 are the non-selected Mentoring applicants, late applicants to Challenges, and the control groups. All individuals are observed over the span of 15 months. The dataset is randomly split between training and testing at the individual applicant level, 50-50.
Propensity model.
To estimate the propensity model for policy estimation, we use multinomial logit regression with the full set of covariates that overlap between all datasets. We train the model in the training sample and predict into the testing set. We find that the propensity of being assigned to Mentoring ranges from 4% to 37%, to Challenges from 5% to 32%, and to Out of Dare IT from 53% to 89%. Figure 13 presents the histogram of the three propensities.
Figure 13: Predicted program propensities.
Note: Test set propensity of assignment across groups, in blue to Challenges, in red to Mentoring, and in green to Out of Dare IT.
We also evaluate the adjustment with the propensity weights by compare the standardized absolute mean differences in covaraites values with and without the adjustment. In Figure 14, we show the difference in average values of covariates between Mentoring and Nothing groups, with and without adjustment in each cross-fitting round. The same analysis for the case of Challenges and Nothing is in Figure 15. We find that the propensity model successfully adjusts for the differences in covariates.
Figure 14: Covariate balance: Mentoring.
Note: Differences in mean covariates across Mentoring and Nothing group with adjustment (in yellow) and without (in blue). Each dot corresponds to one round of cross-fitting.
Figure 15: Covariate balance: Challenges.
Note: Differences in mean covariates across Challenges and Nothing group with adjustment (in yellow) and without (in blue). Each dot corresponds to one round of cross-fitting.
Outcome model.
We use Cox proportional hazard model to model outcomes. We estimate a model with a full set of covariates and interactions of covariates and treatments. We model outcomes at the person and month level. In Figure 16, we show an evaluation of the model’s predictive performance. First, we train a model on the training sample and predict probabilities of having a job in a specific month using the testing sample; then, we take the average per user and group users based on these averages into quintiles. Finally, we compare the predicted probability within the quintile to the observed probability. We find that the differences are statistically insignificant, and the magnitudes are small. We conclude that the model is well-calibrated.
Figure 16: Comparison of predicted and observed outcomes.
Note: Comparison of predicted and observed outcomes in the test set. Subjects are grouped into quintiles based on the predicted probability of having a tech job. In red, the average per quintile predicted probability. In blue, is the average per quintile observed outcome. Whiskers show confidence intervals.
J.1 Cross-fitting.
We estimate predicted outcomes under a policy using cross-fitting with 10 folds and 100 repetitions. Specifically, for each repetition, $m\in\left\{1,...,100\right\}$, we randomly assign observations into 10 folds $k\in\left\{1,..,10\right\}$. Randomization is at the individual subject level with equal probability. Next, for each fold $k$ we train an outcome and propensity models using data from all other folds and predict outcomes and treatments in the $k$-th fold: $\hat{\mu}_{k}(X_{i})_{P}^{m}$ and $\hat{e}_{k}(X_{i})_{P}^{m}$, where $P\in\left\{0,M,C\right\}$ refers to a program. We use these predictions to construct AIPW estimates; for example, the treatment effect from Mentoring instead of Out of Dare IT writes:
$\tau^{m}_{OM}=\frac{1}{n}\sum_{i=1}^{n}\left[\hat{\mu}_{k}(X_{i})_{M}^{m}+\frac{\mathbb{1}_{W_{i}=M}}{\hat{e}_{k}(X_{i})^{m}_{M}}\times(Y_{i}-\hat{\mu}_{k}(X_{i})^{m}_{M})-\hat{\mu}(X_{i})_{0}^{m}+\frac{\mathbb{1}_{W_{i}=0}}{\hat{e}_{k}(X_{i})^{m}_{O}}\times(Y_{i}-\hat{\mu}_{k}(X_{i})^{m}_{O})\right],$
analogously for Challenges. Finally, following Chernozhukov et al. (2017), we take the median over 100 repetitions so that $\hat{\tau}_{OM}=Median(\tau^{m}_{0M})$.
Appendix K Policy estimation algorithms
Algorithm 1 describes the estimation and evaluation of optimal policy.
Algorithm 1 Optimal policy
1. Considering the dataset of all applicants to both programs, randomly split the dataset into train and test with equal probability at the applicant level, denote the resulting sets as $I^{train}=\left\{1,..,\mathcal{I}^{train}\right\}$ and $I^{test}=\left\{1,..,\mathcal{I}^{test}\right\}$.
2. In the training dataset, estimate outcome model using cross-validation: $\mu_{train}$
3. In the test set, construct predicted treatment effects using predictions from model $\mu_{train}$. Obtain $\tau^{Op,train}_{i,t}=E[\hat{Y}(p)_{train}-\hat{Y}(O)_{train}|X=x_{i,t},i\in I^{test}]$, where $\hat{Y}(p)_{train}$ is the predicted outcome under program $p$ constructed using $\mu_{train}$ and $\hat{Y}(O)_{train}$ is the predicted outcome under Out of Dare IT and $t\in\left\{1,...,15\right\}$ denotes months. Compute mean treatment effects per user as: $\tau^{Op,train}_{i}=1/t\times\sum_{t=1}^{15}\tau^{Op,train}_{i,t}$,
4. Assign treatment to maximize treatment effects subject to capacity constraint. Let $Q^{p}$ be the capacity limit of program $p$ and $z_{ip}$ an indicator variable taking the value of one when applicant $i$ is assigned to program $p$ and zero otherwise. We solve the following constrained optimization problem:
$\max_{z_{ip}}\sum_{i=1}^{I}\sum_{p=1}^{P}z_{ip}\tau_{i}^{Op,train}\text{ s.t. }\sum_{i=1}^{I}z_{ip}\leq Q^{p}\forall_{p}\text{ \& }\sum_{i=p}^{P}z_{ip}=1\forall_{i\in I^{test}}.$
The first constraint ensures that the capacity constraints are not violated. The second one is that every applicant is assigned to one program. There is no capacity limit on being Out of Dare IT. We use LP Solve algorithm to solve the problem. We obtain optimal allocation $\mathcal{A}_{X,Q}^{*}=\left\{a_{i}^{*},...,a_{\mathcal{I}^{test}}^{*}\right\}$,
5. Using the test set, estimate new outcome and propensity models using cross-fitting and obtain predictions: $\hat{\mu}_{i,k}$ and $\hat{e}_{i,k}$ for all $i\in I^{test}$. See Appendix J.1 for details of the cross-fitting procedure. Obtain $\hat{Y}_{i,k}(a^{*})$ the AIPW estimates of the predicted outcomes using cross-fitted models trained in the test set,
6. Obtain $\hat{V}_{X,Q}^{*}=\frac{1}{|I^{test}|}\times\sum_{i=1}^{\mathcal{I}^{test}}\hat{Y}_{i,k}(a^{*})$ as the mean of predicted outcomes under the allocation $\mathcal{A}_{X,Q}^{*}$. Estimate standard errors clustered at the applicant level: $\sigma_{X,Q}^{*}$. $\triangleright$
return $(\mathcal{A}_{X,Q}^{*},\hat{V}_{X,Q}^{*},\sigma_{X,Q}^{*},\hat{Y}_{1,k}(a^{*}),...,\hat{Y}_{\mathcal{I}^{test},k}(a^{*}))$
Appendix L Validate AIPW approach
In Section 6, we used estimated AIPW scores to construct and evaluate counterfactual policies. Now, we compare the estimates of treatment effects using the AIPW approach in the testing set with cross-fitting and the experimental results. Thus, we want to test the hypotheses that $\tau^{AIPW}=\tau^{RCT}$. To estimate $\tau^{AIPW}$, we proceed in the same steps as in Section 6: first, we pool all datasets: Mentoring treatment and control, mentoring non-selected applicants, Challenges treatment and control, challenges late applicants. We split the dataset into train and test, and in the testing dataset, we randomly assign 10 folds. Then for each fold, we train a proportional hazard model for the outcomes and a multinomial logistic regression for propensities using other folds, and then we predict outcomes and propensities for the left-out fold and construct AIPW scores. In the outcome model, we include an indicator variable of whether the applicant was selected by a mentor or not. We repeat the process of cross-fitting 100 times and take the average AIPW score per observation.
To assess Mentoring, we take the difference between predicted AIPW scores under mentoring treatment and under mentoring control, i.e., the estimate of the average treatment effect for the entire sample, analogously for Challenges. We compare obtained quantities with the estimates of the treatment effects using Cox models on the experimental samples. We also compute the difference between the estimates. Figure 17 shows the results
Figure 17: Difference between AIPW and RCT estimates.
Note: Comparison of the average treatment effect estimates using the randomized experiment and the AIPW scores. Point estimates and confidence intervals.
We find that the RCT and AIPW estimates are similar, and the differences are statistically not distinguishable from zero. Recall that the randomization into treatment and control was carried out within experimental samples, that is, the sample of Mentoring applicants who mentors selected and the subset of applicants to Challenges that applied to the program before we closed the registration. Thus, it was not a priori clear that the AIPW procedure with cross-fitting would recover similar estimates of the treatment effect after we pooled all datasets together. The finding that these estimates are very close in magnitude and statistically not different from one another bolsters the credibility of results presented in Section 6.
Appendix M Mentors’ selection strategies
In this section, we dive deeper into four mentors’ strategies: Observed, Highest CATE, Highest if treated, and Least promising. We analyze the degree to which they overlap in terms of which candidates get selected and how the characteristics of potential mentees differ across these strategies.
First, we present the Venn diagram in Figure 18. We observe considerable overlap between Observed, Highest CATE, Highest if treated; approximately half of the candidates selected for Mentoring under Highest CATE are also selected in the observed allocation. Furthermore, almost two-thirds of selections under Highest if treated and Highest CATE overlap. Similarly, 100 out of 138 candidates overlap between Highest if trated and Observed selection. In contrast, there is low overlap between Least promising and the other three policies, with 124 candidates that are selected under Least promising not being selected under any other policy.
Figure 18: Venn diagram of selected policies.
Note: Each policy has 138 selected candidates. Percentages refer to the share of the 138 that fall into the region of the graph.
Next, we consider the following sets of candidates: (i) Observed selection – all candidates selected under the current allocation, (ii) Only least promising – candidates that are selected under the Least promising policy and are not selected under the remaining three policies, (iii) Highest CATE & Observed selection – the candidates that are selected both under the Highest CATE and under Observed selection, and (iv) Highest CATE but not selected - where we consider candidates that are selected under Highest CATE, but under mentors’ rule. We analyze the baseline outcomes, average treatment effects, and means of selected covariates. We present results in Table 11
Table 11: ATEs, baselines, and covariate means and standard errors for different policies
policies ATE Baseline STEM Prof. exp 30 plus Small town
Observed selection $0.095(0.012)$ $0.275(0.008)$ $0.457(0.011)$ $7.072(0.110)$ $0.514(0.168)$ $0.232(0.009)$
Only least promising $0.125(0.008)$ $0.048(0.005)$ $0.076(0.006)$ $6.277(0.160)$ $1.000(0.000)$ $0.588(0.012)$
Highest CATE & Observed selection $0.158(0.017)$ $0.244(0.011)$ $0.493(0.015)$ $8.587(0.161)$ $0.507(0.015)$ $0.400(0.015)$
Highest CATE but not selected $0.208(0.016)$ $0.112(0.011)$ $0.365(0.016)$ $10.095(0.168)$ $0.365(0.016)$ $0.873(0.011)$
Note: Average treatment effects and baselines from Section 6.2 standard errors clustered at the candidate level. STEM is the share of candidates selected under the considered policy that have a degree in STEM, Prof. exp is the average length of professional experience, 30 plus is the share of candidates that are above 30 years old, and Small town is the share of candidates that reside in a smaller town.
Comparing candidates selected only under Least promising to other categories, we find that these applicants are less likely to have a STEM degree, have shorter professional experience, and are more likely to be above 30 years old. Even though both mentors overlook this group and we do not select these candidates under the Highest CATE policy, the estimated treatment effects for this group are high. Furthermore, the treatment effect of 0.125 increase in the probability of getting a tech job is particularly high considering the baseline of only 0.048.
Finally, comparing the candidates that have high estimated treatment effects and get selected to those with high treatment effects who are not selected, we notice that first, mentors select high CATE candidates with higher baselines and lower average treatment effects. They tend not to select enough candidates who have longer professional experience and come from smaller towns. These patterns are consistent with our estimates of the heterogenous treatment effects shown in Section 5.2.