### v4

# Susan Athey    Emil Palikot

Abstract
Women remain underrepresented in technology despite many support programs, partly due to limited evidence on which interventions work and how to allocate scarce program slots. We evaluate two approaches to facilitating technology-sector transitions for women in Poland: Challenges, a scalable and cheap-to-operate (\$15/person) online program developing job-relevant portfolios (designed and implemented during this research), and one-on-one mentoring. The programs address distinct labor market barriers: expanding professional networks and teaching their effective use (Mentoring) versus providing credible skill signals (Challenges). Randomizing oversubscribed admissions, we find both programs substantially increase technology employment twelve months post-completion, Mentoring by 15 percentage points (p.p.) and Challenges by 11 p.p. Exploiting treatment effect heterogeneity, we show that algorithmic targeting based on predicted treatment effects achieves outcomes 86% larger than random assignment at current capacity levels. While mentors successfully identify high-benefit candidates (outcomes 48% above random selection), algorithmic allocation using observable characteristics yields an additional 11% improvement. The results demonstrate that (i) relatively low-cost interventions can meaningfully address gender gaps in technology, and (ii) data-driven allocation rules can substantially enhance program effectiveness even when expert judgment performs well. JEL: J24, J08, J16, C93
1 Introduction
An important concern for labor policymakers is identifying and implementing effective programs to support workers transitioning into high-growth sectors. The technology sector exemplifies this challenge: despite rapid employment growth and attractive wages, women hold only 30% of technology jobs worldwide.111unitednations2021 estimates the share of women in the technology sector at approximately 30%. Many programs exist to support women in technology, including NGO initiatives (Women in Technology International, Girls Who Code), corporate programs (Women in Tech from Salesforce, The Women at Microsoft), and government efforts (UN’s Girls in ICT Day, European Commission’s Women in Digital Scoreboard). While women face multiple barriers entering technology, this paper examines two specific labor market frictions. [^2] The gender gap in technology has been attributed to discrimination in hiring and promotion (reuben2014gender; moss-racusin2012science), work-family conflict and child penalties (goldin2014grand; kleven2019children; bertrand2010dynamics), and underrepresentation in STEM education pipelines (bayer2016diversity; carrell2010sex). We focus on network and information barriers because short-term labor market interventions can directly address these constraints, unlike structural factors requiring longer-term policy responses. First, women have weaker professional networks than men, limiting access to job opportunities (greguletz2019women; forret2004networking). Second, gender stereotypes lead employers to discount women’s technical abilities, creating information frictions that credible skill signals could reduce (cheryan2013stereotypical; del2022more; murciano2022missing). Understanding which constraints bind, and for whom, is critical for designing effective policy. If network barriers dominate, interventions should expand professional connections; if information frictions dominate, scalable credentialing programs may suffice. Alternatively, both frictions may bind simultaneously for different populations, necessitating multiple program types and optimal assignment rules. We provide experimental evidence distinguishing between these possibilities.
We evaluate two interventions implemented in Poland by Dare IT, a social impact organization, each targeting one of these frictions. Mentoring pairs participants with senior female technology professionals for four months of one-on-one sessions. While sessions are personalized, mentors consistently emphasize networking: facilitating introductions to hiring managers, expanding mentees’ professional connections, and teaching effective relationship leverage. Challenges is a scalable online program (\$15 per participant) we designed during this research to address skill signaling barriers. The program helps participants build portfolios demonstrating technical proficiency through six bi-weekly industry-designed assignments, culminating in complete projects, mobile app designs or React applications, that participants reference in job applications. [^3] Program development began with structured interviews with hiring managers at twelve Polish technology firms, who uniformly cited the absence of credible practical experience signals in female applicants’ profiles. We designed Challenges to be delivered entirely online, centered on peer interaction rather than scarce mentors, and based on realistic business cases to ensure employer credibility. The program has been replicated across multiple specializations (software testing, cloud computing, UI design) and delivered in Ukrainian for refugees. We exploit both programs’ substantial oversubscription to randomize access and measure employment outcomes via LinkedIn profiles over twelve months. [^4] Technology jobs include positions in technology firms (excluding finance, legal, HR, accounting roles) or technical positions elsewhere (software development, IT support, data analytics). UX and front-end programming roles are classified as technology jobs regardless of firm type.
Both programs generate large employment effects, but with markedly different patterns of treatment effect heterogeneity that reveal which constraints bind for different populations. Mentoring increases technology employment by 15 percentage points (SE 5.7, from a 46% control baseline); Challenges by 11 percentage points (SE 4.8, from 27%). Evidence supports distinct mechanisms. Mentoring increases LinkedIn connections by 52.7 (SE 20.8), a 23% increase, while Challenges shows no detectable network expansion.
The programs benefit different populations, with heterogeneity patterns aligning with each mechanism. Professional experience reveals the starkest contrast: Challenges generates 28 percentage points larger effects for participants with under five years of experience compared to those with longer careers, while Mentoring shows the opposite pattern with 20 percentage points larger effects for experienced workers. This indicates that early-career women benefit from portfolio signals that compensate for limited work history, while experienced women benefit from network expansion that leverages their accumulated professional standing. Geography provides additional evidence: Mentoring generates 26 percentage point effects for women in small towns compared to only 3 percentage points for Warsaw residents (difference: 23 percentage points, SE 14), while Challenges shows no significant geographic heterogeneity (difference: -2 percentage points, SE 12). This pattern is consistent with network barriers binding tightly outside major urban centers where local technology ecosystems are sparse, while online portfolio signals function equally regardless of location. Mentor characteristics also predict treatment effects: participants matched to mentors with managerial experience or over five years in technology show larger treatment effects, consistent with senior mentors possessing richer networks and better access to hiring decision-makers. These differential patterns demonstrate that both network barriers and information frictions constrain technology employment, but for distinct populations defined by career stage or geography.
The differential treatment effects create an allocation problem when program capacity is limited. Both programs remain substantially oversubscribed (5:1 applicant-to-spot ratios for Mentoring; 300 spots filled in four hours for Challenges), requiring selection among applicants. We evaluate counterfactual assignment policies using the methods of athey2021policy and zhou2023offline, estimating which applicants should receive which program to maximize aggregate outcomes. A targeted policy assigning 15% of applicants to each program based on predicted treatment effects achieves outcomes 86% larger than random assignment. This gain reflects two sources of value: selecting individuals with highest treatment effects and matching individuals to their optimal program type. The matching component generates persistent value: even when capacity expands to 50% per program (universal coverage), targeted allocation maintains outcomes 23% larger than random assignment. This demonstrates that complementarity between individual characteristics and program mechanisms creates allocation gains beyond simple capacity expansion, treatment effect heterogeneity enables welfare-improving assignment even when most applicants can be served.
We further compare algorithmic targeting to expert human judgment. Mentoring relies on experienced female technology professionals to select mentees from pre-screened applicants based on pre-recorded short videos and the assessment of the application materials (cover letter, resume, and registration survey). This expert selection substantially outperforms random allocation, achieving outcomes 48% larger. Yet algorithmic allocation using only six observable characteristics (education, location, age, professional experience, degree level, domain preference) achieves outcomes 63% larger than random assignment, an additional 11% improvement over mentor judgment. The two approaches, by experts and by the algorithm, achieve high-degree of concordance (70%): the algorithm largely validates expert choices while identifying additional high-benefit candidates. Despite mentors’ rich soft information, systematic patterns in treatment response remain that simple demographic targeting can exploit. The result extends recent work showing algorithms can complement expert decisions in domains from bail (kleinberg2018human) to medical diagnosis (mullainathan2022diagnosing), demonstrating that even sophisticated expert judgment leaves room for data-driven improvement when treatment effects vary systematically with observable characteristics.
The remainder of the paper proceeds as follows. Section 2 reviews related literature. Section 3 describes the programs, theoretical framework, and testable predictions. Section 4 details the experimental design and data. Section 5 presents treatment effect estimates, mechanisms, and heterogeneity analysis. Section 6 evaluates counterfactual allocation policies. Section 7 concludes.
2 Literature Review
This paper contributes to the literature on treatment effect heterogeneity and optimal policy assignment. Building on recent advances in policy learning using machine learning methods (athey2017econometrics; kitagawa2018should; athey2021policy), we examine whether data-driven targeting can improve program allocation decisions in active labor market policy. Theoretical work demonstrates that when treatment effects vary across individuals and program capacity is limited, policies that assign treatment based on predicted treatment effects can substantially improve welfare relative to random allocation or allocation based on baseline characteristics (kitagawa2018should; bhattacharya2012inferring). Empirical applications of these methods have demonstrated value across diverse domains: customer retention (ascarza2018retention), marketing (hitsch2018heterogeneous), development programs (haushofer2022targeting), agriculture (kasy2021adaptive), and medical treatment (inoue2023machine).
We extend this literature to active labor market policy. Optimal allocation based on predicted treatment effects achieves outcomes 86% larger than random assignment at 15% capacity and 23% larger even at universal coverage. Crucially, targeting based on predicted treatment effects (CATE) dominates targeting based on baseline outcomes or predicted final outcomes, demonstrating that treatment response heterogeneity does not align with baseline heterogeneity. The persistent gains at near-universal coverage reflect complementarity between individual characteristics and program type, creating allocation value beyond capacity expansion.
We further contribute to the growing literature examining whether algorithmic decision-making can complement or improve upon expert human judgment. Recent work demonstrates that machine learning algorithms can match or exceed expert performance in domains from bail decisions (kleinberg2018human) to medical diagnosis (mullainathan2022diagnosing), often by identifying systematic patterns that experts miss or cannot consistently apply. Similar to findings in online labor markets where revealing information about worker quality improves matching outcomes (pallais2014inefficient), we show that observable characteristics predict treatment effects systematically.
We demonstrate that data-driven targeting can improve upon expert human judgment. Mentor-based selection achieves treatment effects 48% larger than random allocation. Yet algorithmic targeting using only six observable characteristics achieves an additional 11% improvement. The 70% concordance between methods indicates complementarity: algorithms largely validate expert choices while identifying additional high-benefit candidates, demonstrating that systematic patterns in treatment response exist even after expert selection captures most available heterogeneity.
This paper contributes to the literature on mentoring effectiveness in improving career outcomes. Several influential papers focus on mentoring in academic contexts: dennehy2017female; resnjanskij2021can; gardiner2007show; ginther2020can; blau2010can show that mentorship increases sense of belonging among female engineering students, youth’ learning outcomes, and tenure rates, publications, and federal grants for women in research careers. [^5] The impact of mentoring on career outcomes has also been studied using observational data. ragins1999mentor survey over 600 individuals and find that mentored respondents have better career outcomes than non-mentored respondents, with effects varying by the gender composition of the mentoring pair. kammeyer2008quantitative conduct a meta-analysis concluding that mentoring improves job and career satisfaction. Mentoring programs may operate partly through expanding professional networks, which play a critical role in labor market outcomes (granovetter1973strength; ioannides2004job; pallais2016referential). We provide randomized trial evidence on mentoring effectiveness in labor market transitions outside academic settings, demonstrating large effects on technology employment for women career-changers. Crucially, we provide experimental evidence that mentoring causally expands professional networks: Mentoring increases participants’ LinkedIn connections by 52.7 (SE 20.8), a 23% increase from baseline, while the non-network-based Challenges program shows no statistically significant effect. We further examine the evaluative selection mechanism inherent in mentor-based programs, where mentors exercise discretion in choosing mentees, and demonstrate that this selection generates substantial treatment effect heterogeneity with observable characteristics of both mentors and mentees predicting treatment effects systematically.
We contribute to the extensive literature evaluating active labor market training programs. card2018works conduct a comprehensive meta-analysis of over 200 studies, finding that traditional training programs aimed at unemployed workers exhibit near-zero short-term employment effects, with modest positive impacts emerging only 2-3 years post-completion. [^6] See also heckman1999economics, card2010active, and crepon2016active for reviews of ALMP effectiveness. Programs incorporating on-the-job skill development tend to outperform classroom-based alternatives (sianesi2008differential; lechner2000microeconometric), though even intensive apprenticeships show mixed results (biewen2014effectiveness). fein2018bridging show that YearUp, a sectoral training program targeting disadvantaged youth, generates 40% earnings gains. However, recent evidence suggests that programs targeting skill signaling and information frictions generate substantially larger employment gains than traditional classroom training (abebe2021anonymity; bassi2022screening; carranza2020job; athey2024moocs). Building on the theoretical work of spence1978job, this literature demonstrates that credentials demonstrating capability can significantly boost earnings (tyler2000estimating; hadavand2018can) and generate market-level benefits when employers face uncertainty about worker quality (pallais2014inefficient). Recent evidence further suggests that training effectiveness may depend on complementary interventions: batista2022closing find that business training for female microentrepreneurs generates significant profit increases only when combined with mobile money access, highlighting that releasing multiple constraints simultaneously may be necessary for training programs to translate into economic gains.
We provide experimental evidence on training program effectiveness in technology sector transitions, a context involving higher-skilled, career-changing women rather than the unemployed workers typically studied. We document substantial short-term effects (visible within twelve months) for both programs, aligning with the signaling literature rather than the null findings in card2018works’ meta-analysis. We directly compare two program modalities: personalized one-on-one mentoring versus scalable online project-based training. This head-to-head comparison reveals complementarity rather than substitutability: the programs serve different populations and generate heterogeneous treatment effects that enable targeted assignment. Our subjects are women seeking technology roles who can utilize skill signaling to counter gender-based statistical discrimination.
Finally, this paper adds to the literature on the gender gap in the technology sector. The persistent gender wage gap reflects both occupational segregation and within-occupation differences (juhn2017specialization). Various reasons for low women’s participation in the technology sector have been studied, including women’s tendency to under-report programming skills (murciano2022missing) and the impact of occupational stereotypes (cheryan2013stereotypical; del2022more). Strategies proposed to counter these barriers include increasing role model visibility (correll2016succeed) and shifting perceptions of technology roles. We contribute to this literature by demonstrating how mentoring and credentialing programs can enhance women’s success rates in obtaining technology jobs.
3 Experimental Interventions and Theoretical Framework
Labor market transitions face two fundamental frictions: imperfect information about worker abilities and search costs in matching workers to opportunities. Workers address these through skill signaling (spence1978job) and professional networks (granovetter1973strength; ioannides2004job; pallais2016referential). For women entering male-dominated technology sectors, both mechanisms face gender-specific barriers. First, occupational stereotypes bias employer perceptions of female candidates’ technical abilities, creating information frictions even for qualified women (cheryan2013stereotypical; del2022more; murciano2022missing). Second, network structures systematically exclude women from informal professional connections that facilitate hiring, with women reporting smaller networks and facing both structural barriers and personal hesitation in networking behaviors that benefit men (greguletz2019women; forret2004networking). Third, women in male-dominated fields face psychological barriers that may reduce job search effectiveness. Women systematically under-report their technical abilities on resumes compared to similarly skilled men (murciano2022missing), use less self-promoting language in professional contexts (lerchenmueller2019gender), and may experience reduced confidence when confronting occupational stereotypes (spencer1999stereotype). We study the relative importance of these barriers by evaluating two interventions with distinct mechanisms: a mentoring program that addresses network gaps and builds confidence through connections to successful female tech professionals, and a scalable portfolio-building program designed to counter information frictions and stereotypes through credible skill signals.
Context.
Poland’s technology sector employed 250,000 workers in 2021, with median monthly wages of USD 2,600, nearly triple the national median (gus2022). Despite strong labor demand driving 9% annual wage growth, women held only 30% of technology positions and comprised just 1.1% of the female labor force, compared to 5.2% of men (EC2021; fluff2021). This persistent underrepresentation mirrors patterns across developed economies where women hold approximately 30% of tech jobs and persists despite numerous initiatives supporting women in technology (unitednations2021).
3.1 Program Designs and Mechanisms
We partnered with Dare IT, a Polish social impact organization addressing gender barriers in technology employment, to evaluate two interventions targeting different constraints. Both programs serve women with baseline technical skills who have not yet secured technology jobs. The programs’ contrasting designs allow us to contrast network effects from signaling effects.
Mentoring: Network Expansion and Tacit Knowledge.
Dare IT launched its Mentoring program in 2018 to support women transitioning into technology careers. The program pairs participants with experienced female technology professionals for weekly one-on-one sessions over four months. Mentors, all women actively working in tech, volunteer their time and are selected based on technical expertise and professional standing. From its inception, demand has substantially exceeded capacity. The spring 2021 cohort received over 1,000 applications for 150 available spots, a ratio typical of recent cohorts. This persistent oversubscription reflects a fundamental scalability constraint: the program’s reliance on volunteer mentors limits expansion, as the pool of qualified female tech professionals willing and able to dedicate four months of weekly mentoring remains finite.
Mentoring addresses gender-specific network barriers documented in prior research. Women in technology report smaller professional networks and face structural exclusions from informal networking opportunities as compared to their male colleagues (greguletz2019women). Female professionals may also hesitate to engage in networking behaviors that appear effective for men (forret2004networking). By connecting mentees to established female professionals, the program potentially overcomes both barriers and provides same-gender role models who can navigate gender-specific challenges in recruitment.
Our qualitative interviews with mentors reveal four operational channels of the Mentoring program: (i.) direct network expansion through mentor connections and introductions to hiring managers, (ii.) transmission of tacit knowledge about navigating tech sector recruitment as a woman, including addressing potential bias in interviews, (iii.) self-efficacy reinforcement to build confidence, and (iv.) technical skill development through code review. Mentors consistently emphasize networking as the primary mechanism, actively facilitating introductions and teaching mentees how to leverage professional connections effectively.
Challenges: Skill Signaling.
To overcome Mentoring’s capacity constraints, we designed Challenges as a scalable alternative targeting information frictions exacerbated by gender stereotypes. Program development began with structured interviews of hiring managers at twelve prominent Polish technology firms. These conversations revealed a systematic pattern: hiring managers perceived female applicants as lacking practical technical experience, even when formal credentials suggested otherwise. Managers emphasized that portfolio projects demonstrating hands-on capability could substantially alter hiring decisions, but expressed skepticism about generic online certifications. These perceptions align with documented gender bias in technical hiring: employers underestimate women’s programming abilities (murciano2022missing) and stereotypes about women’s technical competence create additional screening burdens for female candidates (cheryan2013stereotypical). A credible skill signal may be particularly valuable for women if it counters these negative stereotypes by providing concrete, verifiable evidence of technical proficiency.
Challenges provides six bi-weekly assignments in User Experience Design (UX) or front-end programing in React. All assignments are designed by industry practitioners, that culminate in a complete portfolio project (UX design or React application; see Figure 1 for an example from the UX path). Each assignment replicates realistic business tasks and receives expert evaluation and feedback. Participants can add completed portfolios to resumes and reference them in interviews, creating a signal that hiring managers indicated they would find credible. The program operates online, accommodates cohorts of 300+ participants, and costs approximately \$15 per person.
Figure 1: An Example of Challenges Final Product.
Note: Selected wireframes from a design of a mobile app for pets adoption prepared during Challenges. See: https://www.behance.net/gallery/142047393/Pet-Adoption-App-UXUI for the entire design.
A more comprehensive discussion of the way the Challenges program is organized and it’s content is in Appendix A.
3.2 Theoretical Predictions and Empirical Strategy.
The programs’ differential mechanisms generate testable predictions about the treatment effect heterogeneity. The patterns of heterogeneity reveal which populations benefit most from network expansion versus skill signaling, providing evidence on the economic importance of each mechanism.
If network barriers constrain technology employment for some women, Mentoring should be especially effective for participants with weak existing professional connections: women in smaller cities and towns where local technology networks are sparse, those without family or friends working in technology, and those with longer gaps in employment who may have lost previous professional contacts. The network mechanism also predicts effects on observable network measures: in particular, we should observe increases in LinkedIn connection counts among Mentoring participants. Additionally, if experienced mentors’ networks generate value, treatment effects should vary with mentor characteristics. Mentors with longer technology sector tenure and those in managerial positions likely possess richer professional networks and better access to hiring decision-makers, potentially generating larger effects for their mentees.
If imperfect information about technical abilities constrains hiring, particularly when exacerbated by gender stereotypes, Challenges should disproportionately benefit women whose traditional credentials provide weak or ambiguous signals: for example, those early in their careers with limited work history. For these candidates, the portfolio complements weak existing signals. In contrast, women with strong observable credentials may gain less marginal value from an additional portfolio signal.
We test these predictions by estimating treatment effects across participant subgroups defined by network strength, credential quality, and demographic characteristics. Differential patterns of heterogeneity provide evidence on whether these frictions meaningfully constrain employment and which populations each intervention serves most effectively. We further examine effects on LinkedIn connection counts to directly test the network expansion mechanism.
4 Experimental Design and Data
Mentoring.
We evaluate the Autumn 2021 cohort, for which recruitment opened in October 2021 and the program launched in January 2022. As in previous cohorts, applications substantially exceeded the number of available mentors. Even after Dare IT’s eligibility screening, which excluded applicants currently employed in technology and those lacking technical skills at a junior professional level, the applicant-to-mentor ratio remained 4:1.
This oversubscription allowed us to implement a randomized evaluation while preserving Dare IT’s mentor-mentee matching process. We modified the standard selection procedure: instead of each mentor selecting one mentee (the historical practice), we asked mentors to select their two preferred candidates from the pool of eligible applicants. We then randomized within each pair, assigning one applicant to treatment (participation in Mentoring) and one to control (no participation). This paired randomization design controls for mentor selection preferences while generating exogenous variation in program access.
Of 160 mentors, 151 agreed to this modified selection process; 9 declined, because they did not want to spend effort selecting mentee who will not end up working with them. This yielded an initial sample of 302 applicants (151 treatment, 151 control) across 151 mentor pairs. However, 4 treatment-assigned participants secured technology jobs between randomization (November 2021) and program launch (January 2022). We excluded these individuals from analysis and allowed the initially-assigned control applicant to participate instead, maintaining program access for mentors who invested time in candidate selection. Our final experimental sample comprises 151 treatment participants and 147 control participants.
Appendix G provides robustness checks: (i.) excluding the 4 mentor pairs without matched controls, (ii.) including the 4 pre-program employment cases with outcomes set to 1, and (iii.) including the 9 non-participating mentors under the assumption their hypothetical mentees did not find technology jobs. Results are substantively unchanged across specifications.
In addition to experimental subjects, we collected outcome data on 468 eligible applicants whom mentors did not select. These applicants provide a benchmark for assessing mentor selection effects independent of program participation. As shown in Figure 2, selected-but-control participants have substantially higher employment success than non-selected applicants, highlighting the importance of randomization for isolating treatment effects from selection effects.
Challenges. We launched recruitment in December 2021, with the program scheduled to begin in February 2022. To generate initial interest, we guaranteed admission to the first 100 applicants. Registration was capped at 500 applicants, but high demand led to 525 applications before we closed the portal.
Applicants self-certified that they (i.) were not currently employed in the technology sector, (ii.) possessed English language proficiency at B1 level or higher, and (iii.) had working knowledge of either Figma (User Experience Design track) or React (Front-end Development track). Unlike Mentoring, we did not manually verify technical skills prior to admission due to scalability considerations.
We exclude the first 100 guaranteed-admission applicants from our analysis, as they were not subject to randomization. Among the remaining 425 applicants, we randomly assigned 200 to treatment (program access) and 225 to control (no access). Randomization employed stratification on four registration characteristics: track specialization (UX Design vs. Front-end Development), age category, residence type (Warsaw, other large city, or small town), and employment status (student, employed outside tech, or neither).
Between randomization (December 2021) and program launch (February 2022), 17 treatment-assigned applicants secured technology employment. Consistent with our approach in Mentoring, we exclude these individuals from our primary analysis. Our final analytic sample includes 183 treatment participants and 225 control participants. Appendix G demonstrates that results are robust to including these 17 early-employment cases with outcomes coded as 1.
Among the 183 treatment participants, 41 (22%) did not submit any program assignments, suggesting passive non-compliance. However, these individuals retained access to all program materials, weekly webinars with industry experts, and the Dare IT Slack community. Following intention-to-treat principles, we include them in our analysis regardless of active participation level.
We also collected outcome data on 160 applicants who submitted forms after enrollment closed. These late applicants were not randomized and do not contribute to our primary treatment effect estimates. However, they provide an additional comparison group for the off-policy evaluation in Section 6.
Appendix B reports covariate balance tests across treatment and control groups. Of 9 baseline characteristics, only social science degree shows a statistically significant imbalance at the 5% level, consistent with chance given multiple comparisons. Our main specifications control for all observed variables to improve precision.
4.1 Data
We have four data sources: applicant registration surveys, mentor registration surveys, outcome surveys conducted four months post-program, and LinkedIn profile data.
Applicant Registration Survey.
Both programs collected basic information during registration: city of residence, age, employment situation, and time spent developing IT skills. The Mentoring application additionally asked whether applicants have family or friends in IT and whether they have children. We also collected information on preferred domains: UX design, manager (project or product), or developer (front or back-end).
Mentors Registration Survey.
Mentors provided information enabling us to construct: first-time mentor (binary), career changer (whether the mentor worked outside tech before transitioning to the sector), managerial experience (binary), and years of experience in the tech sector.
LinkedIn Data. Our primary outcome measure, pre-registered prior to data collection, is whether participants secured technology-sector employment following program completion. We collected employment data from participants’ LinkedIn profiles at two time points: four months post-program (May 2022 for Mentoring, June 2022 for Challenges) and twelve months post-program (May 2023 and June 2023, respectively).
LinkedIn provides standardized job title metadata for each employment entry. The research team and research assistants, blinded to treatment assignment, coded each profile according to our pre-registered definition. A technology job is:
  1. 1.
    Any position in a technology company—firms in software development, testing, and sales; data analytics; IT services; digital marketing; or online platforms—excluding finance, regulatory, legal, accounting, and HR roles.
  2. 2.
    Technical positions in non-technology companies involving software development and testing, IT support, or data analytics (e.g., software engineers at banks, data analysts at consulting firms).
Self-employment, freelancing, and volunteer positions were not coded as new employment. Appendix D validates this classification by comparing average salaries for categories (1) and (2) using Glassdoor data; we find no statistically significant difference (p=0.9), supporting their equivalence.
Data collection and attrition.
We obtained complete LinkedIn data for all Mentoring participants at both measurement points. For Challenges, we collected data on all participants at four months and 384 of 408 participants (94%) at twelve months. Missing data resulted from participants changing LinkedIn profile URLs or privacy settings between recruitment and follow-up waves. Appendix G provides bounds and estimates under extreme attrition assumptions; main results remain statistically significant under all scenarios.
Outcome variables.
We define three binary measures: new job (any employment starting January 2022 or later for Mentoring, February 2022 or later for Challenges), tech job (technology-sector position per above criteria), and non-tech job (new employment outside technology). In a few cases, subjects reported only the start year of new employment on their LinkedIn profiles, we conservatively assign December 2022 or May 2023 as the start date.
We also collected from LinkedIn: highest education level (bachelor’s, master’s, postgraduate, non-traditional credential, or other), college major (constructing binary variables for STEM and social science/business), age (combining registration survey data with degree completion dates), and professional experience length (from first job start date to 2022). Table 1 summarizes key variables.
Table 1: Summary Statistics of the Main Variables: Treatment, Control, and Applicants
Variables Treated group Control group
N Mean St. Dev. Min Max N Mean St. Dev. Min Max
Mentoring: mentor characteristics
First time mentor 151 0.47 0.50 0 1 145 0.43 0.50 0 1
Managerial experience 151 0.52 0.50 0 1 145 0.51 0.50 0 1
Career changer 151 5.86 3.79 2 35 145 6.02 3.79 2 35
Years of tech experience 151 0.70 0.46 0 1 145 0.71 0.46 0 1
Mentoring: experimental groups
New job 151 0.65 0.48 0 1 147 0.57 0.50 0 1
Non-tech job 151 0.04 0.20 0 1 147 0.12 0.32 0 1
Tech job 151 0.61 0.49 0 1 147 0.46 0.50 0 1
LinkedIn connections 151 277.45 173.47 0 500 147 223.20 185.60 0 500
STEM 151 0.49 0.50 0 1 147 0.53 0.50 0 1
Student 151 0.14 0.35 0 1 147 0.16 0.37 0 1
Over 30 151 0.47 0.50 0 1 147 0.50 0.50 0 1
Small town 151 0.25 0.44 0 1 147 0.24 0.43 0 1
Applied before 151 0.24 0.43 0 1 147 0.23 0.42 0 1
Mother 151 0.26 0.44 0 1 147 0.26 0.44 0 1
High proficiency 151 0.54 0.50 0 1 147 0.57 0.50 0 1
Family or friends in IT 151 0.70 0.46 0 1 147 0.70 0.46 0 1
UX 151 0.42 0.50 0 1 147 0.44 0.50 0 1
Masters 151 0.48 0.50 0 1 147 0.41 0.49 0 1
Years of professional experience 151 6.99 5.01 0 31 147 7.20 4.82 0 23
Challenges: experimental groups
New job 166 0.57 0.50 0 1 218 0.50 0.50 0 1
Non tech job 166 0.19 0.40 0 1 218 0.23 0.42 0 1
Tech job 166 0.38 0.49 0 1 218 0.27 0.44 0 1
LinkedIn connections 151 277.45 173.47 0 500 147 223.20 185.60 0 500
STEM 183 0.42 0.50 0 1 225 0.39 0.49 0 1
Social science 183 0.21 0.41 0 1 225 0.27 0.44 0 1
Warsaw 183 0.20 0.40 0 1 225 0.20 0.40 0 1
Years of professional experience 183 7.64 4.98 0 31 225 7.67 6.24 0 49
UX 183 0.49 0.50 0 1 225 0.54 0.50 0 1
Over 30 183 0.47 0.50 0 1 225 0.44 0.50 0 1
Small town 183 0.61 0.49 0 1 225 0.63 0.48 0 1
Applicants both programs
Tech job 12 months - - - - - 628 0.28 0.45 0 1
STEM - - - - - 628 0.20 0.40 0 1
Warsaw - - - - - 628 0.31 0.46 0 1
Small town - - - - - 628 0.39 0.49 0 1
Years of professional experience - - - - - 606 7.47 6.32 0 37
Over 30 - - - - - 628 0.14 0.35 0 1
Masters degree - - - - - 628 0.31 0.46 0 1
Mentoring applicant - - - - - 628 0.75 0.44 0 1
Note: Summary statistics of selected variables. The top panel characteristics of mentors. The second panel characteristics of treatment and control groups in the Mentoring experiment. The third panel characteristics of treatment and control groups in the Challenges experiment. The bottom panel, applicants to both programs that are not in experimental groups; this group includes subsets of applicants to Mentoring that were not selected by mentors and applicants to Challenges that submitted their application past registration closure. Career changer takes the value of one when the mentor reported in the registration survey having changed occupations. The variable years of tech experience is reported by mentors in the survey. STEM excludes social sciences and architecture. Social science includes business, management, finance, and accounting. The variable years of professional experience is measured based on LinkedIn profiles. Appendix B shows the covariate balance between treatment and control.
Outcomes Survey.
We surveyed participants and control group members four months post-program, asking about new employment, salary ranges, number of job offers, and negotiation behavior. We conducted three surveys: the main Mentoring survey (162 responses, 111 from treatment), an initial Mentoring survey (194 responses), and the Challenges survey (68 control, 61 treatment). Appendix E analyzes survey outcomes.
5 Effectiveness of Dare IT Programs
Analytical Framework
Let each Dare IT applicant be characterized by $(X_{i},Z_{i},W_{i}^{M},W_{i}^{C},S_{i},Y_{i})$, where $i$ is the applicant’s index $i=1,..,\mathcal{I}$, $X_{i}\in\mathcal{X}$ is a matrix of characteristics of applicants that are observed by mentors and recorded in data, $\zeta_{i}\in\mathcal{Z}$ are characteristic of applicants that are not observed in the data (though some might be observed by mentors prior to treatment assignment), and $S_{i}\in\mathcal{S}$ is a random variable indicating whether the applicant has been selected by a mentor. [^7] To select mentees, mentors received information from the application form; the important part of the application that mentors observed, which we did not, are short introductory videos recorded by the applicants. Finally, $W_{i}^{P}\in\left\{0,1\right\}$ is the treatment assignment into program $P$ and $Y_{i}\in\mathcal{Y}$ is the observed outcome.
Let $Y_{i}(P)$ and $Y_{i}(0)$ be potential outcomes when applicant $i$ participates in program $P$ and when she does not; $P\in\left\{M,C\right\}$, where $M$ stands for Mentoring and $C$ for Challenges. We are primarily interested in estimating the magnitude of the average treatment effect, defined as $\tau^{P}\coloneqq E\left[Y_{i}(P)-Y_{i}(0)\right]$, and testing whether average treatment effects are higher than zero. We are also interested in comparing the treatment outcomes to one another, although our estimates of such contrasts are noisy due to the limited sample size in each experiment.
Selected Sample in Mentoring.
Dare IT asked mentors to choose two applicants they wanted to work with. Thus, selected applicants are not a random draw from the pool. Let $S_{i}$ be a random variable indicating whether a mentor selected the applicant $i$ or not; $S_{i}=1$ when the applicant has been selected by a mentor while $S_{i}=0$ otherwise. Only selected applicants were included in the randomization. Thus, we do not observe outcomes under treatment for the non-selected applicants. The average treatment effect of Mentoring that we estimate is defined as
$\tau^{M}\coloneqq E\left[Y_{i}(M)-Y_{i}(0)|S_{i}=1\right].$ (1)
Table 1 provides summary statistics of subjects in the experimental groups of the mentoring program and the non-selected applicants. There are notable differences between the mentoring control group (those selected by mentors ($S_{i}=1$) but not assigned to treatment ($W_{i}^{M}=0$) and the applicant group (where $S_{i}=0$ and thus, by construction, $W_{i}^{M}=0$), most notably in terms of the primary outcome of securing a tech job. In the Challenges program, participants were selected randomly from eligible applicants, where eligibility was determined based on observable characteristics.
5.1 Average Treatment Effects
In Figure 2, we display the proportions of participants employed in technology jobs over the span of thirteen months post-application to the program. This is illustrated across all experimental groups. [^8] We construct outcomes over time based on the new job’s start date. In some cases, we only observe the year the new job started rather than the specific month. For those observations, we assign December of that year as the starting month. across the experimental groups. In the left figure, we consider three main groups: participants of Mentoring, participants of Challenges, and all other applicants to Dare IT (we refer to this group as Out of Dare IT). We can see that all groups start at zero, with none of the applicants having a job in technology; over time, the share of subjects that have a tech job in each group increases, with the Mentoring group increasing right from the moment the program started, and achieving over 50% probability of having a job in tech after 8 months. In contrast, the Challenges group follows a similar trajectory to the Out of Dare IT group until the 8th month, which is approximately 3 months after the end of the program, but then the gap between those two groups increases to almost 20 percentage points. This differential timing, with Mentoring effects emerging during the program while Challenges effects manifest several months post-completion, suggests distinct mechanisms. These trajectories are consistent with the distinction between immediate network-based job matching versus the creation of portfolio signals and gradual employer response to them.
The right panel decomposes selection from treatment effects in Mentoring. Applicants selected by mentors but randomized to control achieve substantially higher technology employment (43% at month thirteen) than non-selected applicants (19%), confirming that mentors identify promising candidates. Furthermore, the treatment-control gap among selected applicants is smaller than the selection effect itself, underscoring the importance of randomization for isolating program impacts from mentor selection.
Figure 2: Share of Applicants with a Tech Job per Group Over Time.
Note: Share of subjects with a tech job per group across time. The month zero is the month of application to Dare IT. In the left figure, we group subjects according to the experimental groups: the treatment group in Mentoring (in green), the treatment group in Challenges (in red), and Out of Dare IT (in blue), where the last group consists of the applicants in control groups, non-selected applicants to Mentoring, and late applicants to Challenges. In the right figure, we group applicants into those selected and those not selected by mentors. "Selected and treated" group (in green) shows the outcomes of the treatment group in Mentoring, "Selected but not treated" group (in blue) is the control group of Mentoring "Not selected" group (in red) includes applicants to Mentoring that were not selected by mentors, both experimental groups of Challenges, and late applicants to Challenges.
Table 2 presents estimated average treatment effects. We estimate the averate treatment effects using the difference-in-means estimator (columns 1-3, 5-7) and using Cox proportional hazard models that exploit monthly outcome variation (columns 4, 8). Appendix F verifies the proportional hazards assumption; Appendix C presents Kaplan-Meier survival curves confirming the timing patterns in Figure 2.
Both programs generate substantial, statistically significant increases in technology employment. Mentoring increases the probability of technology employment by 15.3 percentage points (SE 5.7), representing a 34% increase relative to the control group baseline of 45.6%. Challenges produces an 11.3 percentage point effect (SE 4.8), a 42% increase relative to the 26.6% control baseline. Both estimates use difference-in-means; results are nearly identical using Cox proportional hazard models that exploit monthly variation in job start dates (14.5pp for Mentoring, 11.2pp for Challenges).
Table 2: Average Treatment Effects of Mentoring and Challenges.
Difference-in-means Prop. hazard Difference-in-means Prop. hazard
Mentoring Challenges
New job Tech job Non-tech job Tech job Cox New job Tech job Non-tech job Tech job Cox
ATE 0.078 0.153*** -0.076** 0.145*** 0.072 0.113** -0.041 0.112**
SE (0.057) (0.057) (0.031) (0.055) (0.051) (0.048) (0.042) (0.051)
ATE % baseline 13.576 33.676** -65.641** 31.718*** 14.458 42.646** -17.600 42.232**
SE (9.893) (14.783) (26.720) (11.992) (10.268) (18.131) (17.978) (19.301)
No. obs. treated 151 151 151 2265 166 166 166 2656
No. obs. control 147 147 147 2205 218 218 218 3488
Note: The estimates of the average treatment effects of Mentoring and Challenges on the probability of finding a new job, a new job in technology, and a new job not in technology. Columns 1-3 and 5-7 present estimates using the difference-in-means estimator and columns 4 and 8 using a proportional hazard model with a full set of covariates. The first row shows estimates of the average treatment effect in the difference in the probability of reporting employment at month 15. The third row presents the estimates from the first row divided by the mean outcome in the control group. Standard errors are presented below the average treatment effect estimates. The variance of the difference-in-means estimate of the Challenges intervention is estimated using the classical Neyman-Splawa estimator. *** p<0.01, ** p<0.05, * p<0.1.
Job Search Selectivity.
Estimates in Column 3 of Table 2 suggest that Mentoring operates partly through increased job search selectivity. While Mentoring increases overall employment by only 7.8 percentage points (SE 5.7), statistically insignificant, it increases technology employment by 15.3 percentage points and decreases non-technology employment by 7.6 percentage points (SE 3.1). This implies that approximately half of Mentoring’s impact reflects participants becoming more selective, holding out for technology positions rather than accepting readily available employment in other sectors. In contrast, Challenges shows no significant effect on non-technology employment (-4.1pp, SE 4.2), consistent with this program operating through skill signaling rather than changing search behavior or persistence. The differential patterns support distinct mechanisms: Mentoring increases both participants’ capability to access technology jobs (through networks and tacit knowledge) and their willingness to persist in pursuing them (for example, through confidence-building), while Challenges primarily improves application success conditional on applying. This finding provides evidence consistent with the self-efficacy channel hypothesized in Section 3: mentors help women maintain focus on technology careers.
In Appendix E, we show the results based on the outcomes survey. We find that the survey respondents in both treatment groups report a higher number of job offers and higher salaries than the control group. The differences between treatment and control are greater in the case of Mentoring than Challenges.
5.2 Impact of Program Participation on the Number of LinkedIn Connections
We directly test the network mechanism by examining whether program participation increases professional network size, measured by LinkedIn connection counts. This analysis distinguishes network expansion from other channels: if Mentoring works solely through skill development or signaling, we should observe no differential effect on connections relative to Challenges. Conversely, if networking is a primary mechanism, Mentoring should increase connections while Challenges, which operates entirely online without mentor introductions or employer contact, should not.
The data support the network mechanism. Mentoring increases LinkedIn connections by 52.7 (SE 20.8) from a control group baseline of 225, a 23% increase. In contrast, Challenges produces a statistically insignificant reduction of 11.4 connections (SE 17.8) from a baseline of 232. The differential treatment effect of 64 connections between programs (SE 28) provides direct evidence that Mentoring expands professional networks while Challenges does not.
To understand where these effects appear in the connection distribution, Table 3 presents quantile regression estimates. Mentoring generates large, statistically significant increases at the 25th percentile (62 connections, SE 27) and median (92 connections, SE 35), but no detectable effect at the 75th percentile. This pattern indicates the program particularly benefits participants with initially sparse networks. In contrast, Challenges shows no significant effects at any quantile, confirming this program does not operate through networking channels.
Table 3: Treatment Effects on LinkedIn Professional Networks.
Quantile Treatment Effect Mentoring Treatment Effect Challenges Difference: Mentoring and Challenges
0.25 62** 5 57*
(27.159) (13.391) (30.281)
0.5 92*** 5 87*
(35.426) (29.470) (46.082)
0.75 5 -8 13
(47.004) (50.706) (69.141)
Note: The table provides the quantile regression results for the Challenges and Mentoring programs. The treatment effects of the experimental groups on LinkedIn connections are shown at the 25th, 50th, and 75th quantiles. The difference column captures the difference in the treatment effects between the Mentoring and Challenges programs. Estimates are obtained by regressing the treatment indicator on the number of connections. Standard errors are in parentheses. *** p<0.01, ** p<0.05, * p<0.1.
The magnitude of network expansion is economically meaningful. A 92-connection increase at the median represents substantial growth from the control baseline of 225, comparable to the network size differences between mid-career and senior professionals in our sample. Moreover, the increase occurs despite LinkedIn’s 500-connection display cap, which likely attenuates our estimates: 25% of Mentoring treatment participants hit this ceiling compared to 23% of controls, suggesting the true effect exceeds our measured 52.7-connection average.
This network expansion cannot be solely attributed to new employment. Both Mentoring and Challenges participants secure technology jobs at high rates, yet only Mentoring participants expand their networks. Had job acquisition been the primary driver of connection growth, through adding new colleagues and industry contacts post-employment, we should observe comparable increases in both programs. The absence of connection growth among Challenges participants, despite their 38% technology employment rate, indicates that Mentoring’s network effects operate independently of (and plausibly prior to) job acquisition. This timing is consistent with mentors actively facilitating introductions during the program, as reported in our qualitative interviews, rather than networks expanding passively after job placement.
5.3 Heterogeneous Treatment Effects
The hypothesized contrasting mechanisms underlying Mentoring and Challenges generate testable predictions about which participants benefit most from each program. We exploit treatment effect heterogeneity to provide evidence consistent with these mechanisms: if Mentoring operates primarily through network expansion (as suggested by LinkedIn connections analysis), effects should concentrate among participants with weak existing networks; if Challenges works through skill signaling, effects should be largest for those with weaker pre-treatment credentials. We estimate heterogeneous treatment effects using Cox proportional hazard models that estimate treatment effects in subgroups defined by the pre-treatment characteristics of interest. We control for the full covariate set.
Table 4 presents heterogeneous treatment effects across three participant characteristics that map to credential strength and network access. These patterns provide evidence on which populations each program serves most effectively and tests the distinct mechanisms we hypothesize.
The clearest differential pattern emerges from professional experience. Challenges generates 28 percentage points larger treatment effects for participants with under five years of experience compared to those with longer careers (SE 10), whereas Mentoring shows the opposite pattern with 20 percentage points larger effects for experienced participants (SE 11). The cross-program differential is 49 percentage points (SE 11). This contrast directly supports our theoretical predictions: early-career women lack established work history to credibly signal their technical abilities, making the Challenges portfolio particularly valuable as a complement to sparse employment records.
Educational background reveals similar patterns. Recall that all participants were pre-screened for technical competency, so heterogeneity by educational background likely reflects differences in credential signals and professional networks, and to a lesser degree underlying ability. Among Mentoring participants, those without STEM degrees show 19 percentage points larger effects than those with STEM backgrounds (SE 11). This pattern is consistent with the network access: non-STEM participants followed educational paths outside technology-adjacent fields, limiting opportunities to build connections with peers entering the sector.
In contrast, Challenges shows no STEM heterogeneity (difference: 0.3pp, SE 11). This null result demonstrates that portfolio signals are equally valuable regardless of educational credentials—employers evaluate demonstrated capability rather than inferring skill from degrees.
Geography provides additional evidence on network mechanisms. Mentoring generates 26 percentage point treatment effects for participants from small towns (SE 10) compared to only 3 percentage points for Warsaw residents (difference: 23pp, SE 14, p=0.11). This pattern aligns with network barriers binding more tightly outside major urban centers, where local technology ecosystems are sparse and professional connections to the sector are limited. Mentor networks based in Warsaw and other major cities provide greater marginal value when participants lack local access to technology professionals. Challenges shows no significant geographic heterogeneity (difference: -2pp, SE 1.2), consistent with online portfolio signals functioning equally regardless of location, employers evaluate portfolios based on technical content rather than applicant geography.
Table 4: Heterogeneity in Treatment Effects: Mentoring vs. Challenges.
Mentoring Challenges
Baseline CATE Baseline CATE Diff. Mentoring – Challenges
STEM degree
Yes 0.538 (0.057) 0.038 (0.079) 0.283 (0.059) 0.12 (0.099) -0.081 (0.082)
No 0.362 (0.058) 0.228 (0.07) 0.259 (0.035) 0.116 (0.057) 0.111 (0.068)
Difference subgroups 0.176 (0.081) -0.189 (0.106) 0.024 (0.068) 0.003 (0.114) -0.193 (0.106)
Length of professional experience
More than 5 years 0.412 (0.049) 0.202 (0.066) 0.31 (0.039) -0.003 (0.065) 0.205 (0.063)
Less than 5 years 0.556 (0.075) 0 (0.087) 0.19 (0.042) 0.281 (0.075) -0.281 (0.086)
Difference subgroups -0.144 (0.089) 0.203 (0.109) 0.120 (0.057) -0.284 (0.099) 0.486 (0.106)
Small Town
Small town 0.429 (0.085) 0.26 (0.101) 0.256 (0.067) 0.149 (0.13) 0.111 (0.108)
Warsaw 0.447 (0.082) 0.033 (0.103) 0.182 (0.044) 0.166 (0.089) -0.133 (0.093)
Difference subgroups -0.019 (0.118) 0.227 (0.144) 0.074 (0.157) -0.019 (0.118) 0.246 (0.197)
Note: Estimates of the average treatment effects conditioned on subgroups in Mentoring and Challenges. The outcome variable is Tech job, and estimates are obtained using Cox proportional hazards models per subgroup with a full set of covariates. The first and third columns show the average outcome per subgroup within the control groups. The second and fourth columns show the difference between the treatment and control groups in each subgroup. The last column shows the difference in the treatment effects between Mentoring and Challenges. Standard errors are in parentheses.
Additional Characteristics from Mentoring.
The Mentoring registration survey was more extensive than the Challenges one, allowing us to observe more characteristics. We also obtain the characteristics of mentors from their registration survey. In Table 5, we present additional results documenting heterogeneity in treatment effects along these dimensions.
Table 5: Heterogeneity in Treatment Effects: Mentoring.
Characteristic variable value is 0 Characteristic variable value is 1
Baseline CATE Baseline CATE Diff. Subgroups
Mentee characteristics
Mothers 0.237 (0.07) 0.348 (0.097) 0.532 (0.048) 0.071 (0.062) 0.277 (0.115)
Applied before 0.5 (0.087) 0.139 (0.109) 0.442 (0.047) 0.143 (0.06) -0.004 (0.124)
Mentor characteristics
Managerial exp. 0.378 (0.056) 0.259 (0.073) 0.549 (0.059) 0.002 (0.074) 0.257 (0.104)
Long exp. in tech 0.39 (0.063) 0.277 (0.085) 0.512 (0.054) 0.021 (0.066) 0.257 (0.107)
1st time mentor 0.444 (0.062) 0.155 (0.08) 0.476 (0.055) 0.1 (0.073) 0.056 (0.108)
Note: Estimates of the average treatment effects conditioned on subgroups in the Mentoring program. The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazards models, which include the treatment indicator and the set of controls. We estimate the model per subgroup. "Long exp. in tech" takes the value of 1 if the mentor has worked for over 5 years in the technology sector (the median value). Columns 2 and 3 describe the values for subgroups when the subgroup indicator takes the value 0, while columns 4 and 5 describe the values when the indicator takes the value 1. The last column shows the difference in the effectiveness of Mentoring between the two subgroups. Columns 1 and 3 are the average outcomes in the control groups, and columns 2 and 4 are the estimates of the average treatment effect. Standard errors are in parentheses.
Mentors with managerial experience generate 26 percentage point larger treatment effects than those without (SE 10). Similarly, mentors with over five years of technology sector experience produce 26 percentage point larger effects (SE 11). These patterns are consistent with the network channel: managers and senior professionals possess richer networks, better access to hiring decision-makers, and greater ability to facilitate job placements.
Interestingly, mentor selection patterns reveal that experienced, managerial mentors choose less promising candidates. Control group participants matched to managerial mentors have 38% baseline technology employment compared to 55% for those matched to non-managerial mentors. Yet treated participants of managerial mentors achieve higher outcomes (64% vs. 64%), because the treatment effect is substantially larger. This suggests that experienced mentors either (i.) deliberately select candidates who need more help, recognizing their own capacity to assist difficult cases, or (ii.) value different candidate attributes than baseline employment probability.
Notably, first-time mentors are as effective as repeat mentors (difference: 6pp, SE 11). This finding indicates that mentoring effectiveness stems from mentors’ accumulated professional networks, technical knowledge, and industry standing—not from mentoring skill developed through practice. A senior engineer with extensive connections generates value on her first mentoring experience; repeated mentoring does not improve outcomes.
Mothers and Work-Family Barriers.
The largest heterogeneity we observe is for mothers. Among non-mothers, Mentoring generates 7 percentage point treatment effects (SE 6); among mothers, the effect is 35 percentage points (SE 10), a difference of 28 percentage points (SE 12). Mothers in the control group face substantial disadvantage, with 24% technology employment compared to 50% for non-mothers. Mentoring completely eliminates this gap: treated mothers achieve 59% employment, surpassing non-mothers.
This large effect for mothers is consistent with Mentoring addressing multiple barriers simultaneously. First, mothers may have weaker networks if career interruptions weakened professional connections. Second, work-family conflict concerns may reduce mothers’ confidence in pursuing demanding technology careers; mentor encouragement and role-modeling by successful mentors in tech may counter this.
This finding contributes to understanding gender gaps in technology beyond representation. Even among women who have acquired technical skills and are actively seeking technology employment, motherhood creates additional barriers. Traditional explanations emphasize mothers’ labor supply decisions or employer discrimination; our results suggest that confidence and network deficits also play important roles, and relatively light-touch mentoring interventions can substantially mitigate these barriers.
The contrasting heterogeneity across programs aligns with our theoretical framework. Network barriers might bind most tightly for geographically isolated women and mothers, the groups showing the largest Mentoring effects. Information frictions bind most for early-career women with limited work history, those showing the largest Challenges effects. The programs’ differential impacts across populations indicate that successful policies must target multiple, distinct barriers rather than assuming a single mechanism operates uniformly. Moreover, the patterns demonstrate that even within the selected population of women with baseline technical skills, substantial heterogeneity exists in which constraints bind most tightly, with implications for optimal program assignment that we explore in Section 6.
6 Counterfactual Program Assignment Policies
The number of spots in each program is limited. The availability of mentors restricts the capacity of Mentoring. In Challenges, we limited the available spots to 300 to accommodate potential operational complexities and logistical problems. Considering the substantial treatment effects of both programs, relaxing these capacity constraints will increase Dare IT’s impact. However, capacity constraints are likely to remain a long-term issue for the Mentoring program and may persist, albeit to a lesser extent, for Challenges. Consequently, Dare IT will continue to face the question of which applicants to accept for each program. Given the heterogeneous treatment effects, admission policies that favor applicants with high predicted benefits could enhance the impact of Dare IT programs. This section aims to quantify these potential gains.
We perform three analyses: (i.) we estimate gains from the introduction of a program admission policy that, given capacity levels in each program, assigns applicants to programs, maximizing the sum of predicted treatment effects of programs’ participants; (ii.) we evaluate the gains from relaxing capacity constraints with the targeting policy in place; (iii.) we compare the overall impact of Mentoring when mentors select the most promising candidates to a case where they select to following alternative strategies.
An assignment policy is a mapping from the characteristics of the applicants $X\in\mathcal{X}$ and the capacity levels of the programs $Q=(Q^{M},Q^{C})\in\mathcal{Q}$ to assignments: $\pi:\mathcal{X}\times\mathcal{Q}\rightarrow\mathcal{P}$, where $\mathcal{P}\equiv\left\{M,C,O\right\}$ and $M$ stands for Mentoring, $C$ for Challenges, and $O$ for Out of Dare IT. $\mathcal{Q}$ is the grid of possible capacity levels under consideration.
Policy estimation and evaluation are performed separately in train and test sets. Across all analyses, we obtain train and test sets by randomly splitting data at the applicant level with equal probability ($\mathcal{I}_{train}$ and $\mathcal{I}_{test}$ are resulting sets of applicants in the train and test data sets, respectively). The objective of policy estimation is to obtain a mapping from applicant characteristics and programs’ capacity levels to program assignments. We use the train set to obtain the estimate of the policy $\pi$, which we denote $\hat{\pi}$.
Next, to evaluate a policy, we proceed in two steps: first, we apply the estimated mapping $\hat{\pi}$ to the test set, where for each $i$ in the test set and $Q\in\mathcal{Q}$, we obtain the counterfactual assignment of the applicants to programs $\hat{\pi}(X_{i},Q)\in\mathcal{P}$. Second, we construct the AIPW estimate of the “policy value,” that is, the outcome value when treatments are assigned according to $\hat{\pi}(X_{i},Q)\in\mathcal{P}$. We do that using cross-fitting at the applicant level in the test sample, i.e., we estimate new outcome and propensity models and obtain AIPW estimates of the policy values in held-out folds; specifically, we use 100 repetitions with 5 folds and consider the median outcome across repetitions in (chernozhukov2017double).
Throughout this section, we combine the two experimental data sets with data on non-selected applicants to Mentoring and late applicants to Challenges. Thus, we have 997 applicants who did not participate in any program within our sample, resulting in a total sample of 1314 applicants and 19,000 applicant-month observations. For both policy estimation and evaluation, we use survival forests to estimate outcome models and regression forests for the propensity models. The propensity model adjusts for differences between participants and non-participants determined by selection to Mentoring and differences in the populations of applicants to Mentoring and to Challenges. Appendix H provides details of the models used in policy estimation. Appendix I, presents details of the models used in the policy evaluation, the cross-fitting procedure, and validates the approach by comparing the off-policy estimates with the randomized controlled trial estimates. Our procedure recovers the experimental estimates.
Outcome measure.
In all the preceding analyses, we measured treatment effects as changes in the probability of securing technology employment. For policy evaluation, we adopt an alternative metric that captures additional dimensions of program impact: the restricted mean survival time (RMST), which measures the average time until technology job acquisition over our observation period. The RMST metric offers two advantages for comparing policies. First, it incorporates both whether participants find jobs and how quickly they do so, a policy that helps participants find jobs faster, even if it doesn’t change the ultimate employment probability, generates economic value. Second, RMST estimates aggregate more naturally across heterogeneous populations and policy scenarios, facilitating comparisons of the assignment rules. We report RMST gains in months: negative values indicate participants find technology employment sooner relative to non-participants. An RMST gain of -3 months means treated participants find technology jobs an average of three months earlier. We use survival forests to estimate average treatment effects (athey2019generalized; cui2023estimating).
6.1 Targeted Assignment Policies
Let $z_{ip}\in\left\{0,1\right\}$ be a variable that indicates whether applicant $i$ is assigned to program $p\in P$, and $z_{ip}^{*}$ be the arguments that maximize the sum of treatment effects given by Equation 2:
$\max_{z_{ip}}\sum_{i=1}^{I}\sum_{p\in\mathcal{P}}z_{ip}\tau_{i}^{p}\text{ s.t. }\sum_{i=1}^{I}z_{ip}\leq Q^{p}\thickspace\forall p\text{ \& }\sum_{p\in\mathcal{P}}z_{ip}=1\thickspace\forall i,$ (2)
where $\tau_{i}^{p}=\mathbb{E}\left[Y(p)-Y(O)\right]$. The first constraint ensures that capacity limits are not violated, and the second is that each applicant is assigned to one allocation. A policy that induces allocation $z_{ip}^{*}$ is what we call the targeted policy. $\tau_{i}^{p}$’s are estimated using the AIPW estimator trained on the train set. In Algorithm 1 in Appendix H, we provide further details of the estimation of the targeted policy.
The outcome of interest, policy value, is the participants’ average treatment effect given program allocation induced by the policy $\pi$. Specifically, we use the AIPW estimator to estimate the average difference in outcomes between being selected to participate based on the assignment policy and not participating in any Dare IT program. This emphasis on impact, rather than average outcomes of participants, aligns with Dare IT’s mission to create the greatest benefit for their participants.
6.2 Value of Targeted Policies
Given limited program capacity, which applicants should be assigned to which program to maximize overall impact? We estimate the targeted policy at three capacity levels: the 15/15/70 policy approximates existing capacity (assigning 15% of all applicants to Mentoring and 15% to Challenges); the 33/33/34 policy increases capacity to 33% for each program; and the 50/50/0 policy allows each program to admit 50% of applicants, ensuring a program slot for every applicant. We compare these policies against random allocation as a benchmark (random policy allocates all applicants at random to either Mentoring or Challenges). In Section 6.3, we examine how mentor-based selection compares to the algorithm-driven policy in greater detail. The policies vary in both capacity levels and whether participants are systematically targeted.
To understand which applicant types are assigned to each program under data-driven targeting, Figure 3 presents standardized differences in baseline characteristics between optimally assigned individuals and those not assigned to treatment. We implement the 15/15/70 targeted allocation policy and compare covariate distributions of each assigned group to the not-assigned group. Circles represent comparisons for Mentoring; triangles represent Challenges. Dashed vertical lines mark the $\pm$0.1 threshold for imbalance.
Figure 3: Characteristics of Optimally Assigned Groups
Notes: Standardized differences in baseline characteristics between optimally assigned individuals and those not assigned under the 15/15/70 allocation policy. Circles show Mentoring vs Not Assigned; triangles show Challenges vs Not Assigned. The solid vertical line at zero indicates perfect balance; dashed lines mark the $\pm$0.1 threshold. Standardized differences are calculated as the difference in means divided by the pooled standard deviation. Positive values indicate the optimally assigned group has higher values on that characteristic.
The targeted allocation exhibits clear differentiation in targeting patterns across programs. Applicants selected for Mentoring are more likely to hold STEM degrees, live outside Warsaw, have Master’s degrees, and be over 30 years old. Notably, although the algorithm does not use information about prior mentor selection, previously mentor-selected applicants are disproportionately targeted by the algorithm—suggesting the model captures similar signals to those used by mentors. For Challenges, the algorithm shows more balanced selection across characteristics. Assigned applicants are moderately more likely to have been previously selected by mentors but are less likely to hold STEM degrees, live in small towns, or have extensive professional experience compared to those not selected.
Main result.
Table 6 shows gains from data-driven targeting, with effect sizes varying by program and capacity constraint. Under the most selective policy (15/15/70), the overall treatment effect reaches 4.34-4.34 months, which corresponds to an 86% improvement over random allocation’s 2.33-2.33 months. These gains exhibit a strong capacity-selectivity tradeoff: as capacity expands from 15% to 50%, per-person treatment effects decline (from 4.34-4.34 to 2.86-2.86 months overall), though even universal coverage (50/50/0) achieves effects 23% larger than random allocation. Thus, we find that there are gains from targeting across programs.
The targeting gains differ across programs. Mentoring exhibits the largest effects under selective policies (5.55-5.55 months at 15% capacity versus 2.31-2.31 under random), while Challenges shows more modest differentiation (3.13-3.13 versus 2.35-2.35). Notably, even at 50% capacity where the algorithm must accept marginal applicants with lower predicted treatment effects, Mentoring maintains treatment effects 54% larger than random allocation, demonstrating that observable characteristics contain substantial predictive power for treatment response beyond what random assignment captures.
Table 6: Targeted Allocation Policies: Treatment Effects by Program
Capacity Allocation Mentoring Challenges Overall
(Ment. / Chal. / None) Effect Effect Effect
Panel A: Benchmark
Random Allocation -2.31 (0.022) -2.35 (0.021) -2.33 (0.022)
Panel B: Targeted Allocation Policies
15% / 15% / 70% -5.55 (0.086) -3.13 (0.069) -4.34 (0.055)
33% / 33% / 34% -4.33 (0.052) -2.68 (0.041) -3.50 (0.033)
50% / 50% / 0% -3.55 (0.039) -2.17 (0.027) -2.86 (0.024)
Notes: Treatment effects measured as change in months to employment (negative values = faster job finding). Standard errors in parentheses. Each row shows a capacity-constrained allocation policy specifying the share of applicants assigned to Mentoring, Challenges, or neither program. Random Allocation assigns each program to a random 50% of applicants.
Table 7 examines treatment effects within each assignment group under the 15/15/70 targeted policy. An assignment group consists of all applicants allocated to a specific program (Mentoring, Challenges, or neither) under this policy. For each group, we estimate counterfactual treatment effects under alternative program assignments to assess whether the targeted allocation correctly matches individuals to their highest-benefit program.
The results reveal effective targeting. The Mentoring assignment group receives large benefits from their targeted allocation (5.55-5.55 months, SE = 0.086), substantially exceeding what they would gain from Challenges (3.12-3.12 months, SE = 0.062). The difference of 2.43 months (SE = 0.106) is statistically significant. The Challenges assignment group also benefits most from their assigned program (3.13-3.13 months, SE = 0.069) compared to Mentoring (1.87-1.87 months, SE = 0.050); a difference of 1.26 months (SE = 0.085). This pattern demonstrates benefits from targeting across both programs rather than only selecting to a program.
The Out of Dare IT group (70% of applicants receiving no program assignment) exhibits substantially smaller treatment effects under either counterfactual assignment: 1.72-1.72 months (SE = 0.024) for Mentoring and 2.02-2.02 months (SE = 0.022) for Challenges. While this group would experience gains from program participation, their predicted effects are smaller than those optimally assigned to treatment.
Table 7: Gains From Assignment Policy per Assignment Group
Mentoring group Challenges group Out of Dare IT group
Share 0.15 0.15 0.70
ATE under assignment to:
Targeted -5.552 (0.086) -3.132 (0.069)
Mentoring -5.552 (0.086) -1.872 (0.050) -1.720 (0.024)
Challenges -3.122 (0.062) -3.132 (0.069) -2.017 (0.022)
Difference in ATE between assignments:
Mentoring - Challenges -2.431 (0.106) 1.259 (0.085) 0.297 (0.032)
Note: ATEs are cross-fitted AIPW estimates from causal survival forests (RMST changes up to the evaluation horizon). Negative values indicate earlier job finding (months sooner) relative to out of Dare IT programs. Standard errors in parentheses are GRF mean-of-variance plug-in SEs aggregated within arm.
Policy value across capacity levels.
When comparing the average treatment effects under targeted assignments and the random assignment, we find that there are subsets of applicants who benefit substantially from targeting. However, when the capacity of the two programs increases, the additional participants who get in have lower treatment effects, as suggested by the comparison of targeted policies in Table 6. Thus, there are diminishing returns to targeting additional slots. To identify the level of capacity that realizes most of the available gains from targeting, we estimate the outcomes from targeted assignment under a sequence of capacity values and compare the resulting average probability of getting a job in technology with this outcome under random assignments to Mentoring and to Challenges.
Figure 4: Policy Value Across Capacity Levels: Targeted vs Random Allocation
Notes: Average treatment effects under targeted and random allocation policies as capacity varies from 5% to 50% per program. The solid curve shows targeted allocation that assigns individuals with highest predicted treatment effects to each program; the dashed line shows random assignment of 50% to each program. Treatment effects measured as change in months to employment (negative = faster job finding). Shaded region indicates 95% confidence interval for targeted policy. Policy value evaluated using cross-fitted AIPW estimates on test set. At each capacity level, targeted allocation maximizes expected policy value by solving the assignment problem in Equation 2.
Figure 4 presents the capacity-selectivity tradeoff by plotting average treatment effects as program capacity varies from 5% to 50% per program. The solid curve shows targeted allocation; the dashed horizontal line represents random assignment. Several patterns emerge. First, targeted allocation consistently outperforms random allocation across all capacity levels, with gains largest when capacity is most constrained. At 15% capacity per program, the targeted policy achieves treatment effects of 4.3-4.3 months compared to random allocation’s 2.3-2.3 months, a 2-month improvement. Second, the targeted policy exhibits sharply diminishing returns as capacity expands: per-person effects decline from approximately 4.5-4.5 months at 5% capacity to 2.9-2.9 months at 50% capacity. This reflects the algorithm’s sequential assignment of highest-benefit individuals first, with marginal participants exhibiting progressively smaller treatment effects. Third, even at universal coverage (50% per program, accommodating all applicants), targeted allocation maintains a 0.6-month advantage over random assignment—demonstrating that complementarity between individual characteristics and program type generates value beyond pure capacity expansion.
6.3 Counterfactual Mentors’ Selection Policies
Applicants selected by mentors and randomly assigned to the control group had a significantly higher likelihood of securing a tech job compared to those who were not selected (see Figure 2). This suggests that mentors identify promising applicants who are likely to secure a tech job regardless of their participation in Mentoring. Table 8 compares characteristics of selected and non-selected applicants (panel A), and selected by different groups of mentors (panels B and C). Baseline is the share of control group applicants that found a job.
Table 8: Differences in Applicant Characteristics: Selection and Mentor Experience
Baseline Small town Warsaw Over 30 Prof. exp. Social science STEM Applied before Mother High skill Family/friends in IT
Panel A: Selected vs. Not Selected
Selected 0.456 0.245 0.268 0.487 7.097 0.134 0.510
(0.041) (0.025) (0.026) (0.029) (0.284) (0.020) (0.029) (–) (–) (–) (–)
Not 0.286 0.476 0.282 0.147 7.489 0.207 0.182
(0.021) (0.023) (0.021) (0.016) (0.288) (0.019) (0.018) (–) (–) (–) (–)
Panel B: Mentor Managerial Experience
High 0.378 0.229 0.281 0.588 7.595 0.137 0.497 0.261 0.288 0.614 0.699
(0.057) (0.034) (0.036) (0.040) (0.417) (0.028) (0.041) (0.036) (0.037) (0.039) (0.037)
Low 0.549 0.266 0.245 0.378 6.503 0.133 0.531 0.210 0.231 0.497 0.699
(0.059) (0.037) (0.036) (0.041) (0.384) (0.028) (0.042) (0.034) (0.035) (0.042) (0.038)
Panel C: Mentor Total Experience (above vs. below median)
Above 0.390 0.239 0.299 0.590 8.197 0.171 0.564 0.282 0.308 0.573 0.718
(0.064) (0.040) (0.043) (0.046) (0.442) (0.035) (0.046) (0.042) (0.043) (0.046) (0.042)
Below 0.512 0.251 0.240 0.419 6.330 0.112 0.480 0.207 0.229 0.547 0.687
(0.054) (0.033) (0.032) (0.037) (0.364) (0.024) (0.037) (0.030) (0.031) (0.037) (0.035)
Notes: Means shown with standard errors in parentheses. Baseline refers to the average outcome when not participating in Mentoring. Panel A compares applicants who were selected vs. not selected. Panel B compares applicants selected by mentors with vs. without managerial experience. Panel C compares applicants selected by mentors above vs. below the median years of experience.
Panel A of Table 8 reveals systematic differences between applicants selected by mentors and those not selected. Selected applicants exhibit substantially stronger baseline employment prospects (0.456 vs 0.286), despite having similar years of professional experience (7.1 vs 7.5 years). Selected applicants are more likely to be over 30 years old (0.487 vs 0.147) and to hold STEM degrees (0.510 vs 0.182), while being considerably less likely to come from small towns (0.245 vs 0.476). They are also less likely to have social science backgrounds (0.134 vs 0.207).
Panel B compares mentors with versus without managerial experience. Mentors with managerial experience select applicants with weaker baseline employment prospects (0.378 vs 0.549) but who are substantially older (0.588 vs 0.378) and have more professional experience (7.6 vs 6.5 years). Panel C shows that mentors with above-median total career experience exhibit similar but more pronounced patterns. They select applicants with weaker baselines (0.390 vs 0.512) who are older (0.590 vs 0.419), have substantially more professional experience (8.2 vs 6.3 years), and are more likely to hold a STEM degree (0.564 vs 0.480). The consistency across Panels B and C suggests a common selection philosophy among more experienced mentors: they look beyond current employability to identify "high-potential, high-barrier". Less experienced mentors appear to select more conventionally, favoring applicants already demonstrating stronger baseline employability.
We use our off-policy tools to examine alternative selection strategies that mentors could adopt and their potential implications for the mentoring program’s effectiveness. Specifically, we consider four scenarios: (i) Highest CATE—targeting applicants predicted to have the highest treatment effects, (ii) Most Promising—targeting applicants expected to have the highest outcomes without participating in Mentoring, (iii) Least Promising—selecting applicants expected to have the lowest outcomes if they do not participate, and (iv) Highest Chances if Treated—selecting participants with the highest predicted outcomes if they participate in Mentoring. Additionally, we include two benchmarks: mentors selecting at random (Random Choice) and the observed allocation (Observed). In each of the four policies, we assume that 15% of applicants participate in Mentoring, which is an approximate share of participants to applicants observed by DareIT in recent years.
Table 9 examines the extent to which different allocation policies select the same individuals Each cell shows the percentage of individuals in the row policy who also appear in the column policy. First, we consider the relationship between mentor-based selection (Current) and the Highest CATE algorithmic policy. Among current mentor-selected participants, 70.4% would also be assigned to Mentoring under the Highest CATE algorithm, demonstrating substantial concordance between expert judgment and data-driven targeting (note that we do not use the information on whether a candidate was selected or not in the policy estimation).
Current mentor selections align most closely with Highest CATE (70.4%), followed by Most Promising (29.6%), with little overlap with Least Promising (18.3%) or Highest Chances (19.7%). This pattern suggests mentors implicitly optimize for treatment effects rather than baseline characteristics or final outcomes, though with some cream-skimming tendency. The 29.6% overlap with Most Promising indicates approximately one-third of mentor selections favor strong baseline candidates, while the predominant 70% overlap with Highest CATE confirms that mentors primarily target based on perceived program responsiveness. The low overlap with Least Promising (18.3%) suggests mentors, whether by design or implicit bias, rarely select individuals with the weakest baseline labor market attachment, even when such individuals might benefit from treatment.
Table 9: Overlap Between Allocation Policies
Policy H. CATE Most Prom. Least Prom. H. Chances Current
Highest CATE 100.0
Most Promising 28.5 100.0
Least Promising 18.2 0.0 100.0
Highest Chances 27.9 26.7 31.5 100.0
Current 70.4 29.6 18.3 19.7 100.0
Notes: Each cell shows the percentage of individuals in the row policy that also appear in the column policy. Only lower triangle shown to avoid redundancy; upper triangle omitted. Sample sizes in parentheses after policy names. Diagonal entries are 100% by construction (bold). H. CATE = Highest CATE; H. Chances = Highest Chances; Most/Least Prom. = Most/Least Promising.
Table 10 compares treatment effects across six allocation policies for the Mentoring program. The results demonstrate that targeting based on predicted treatment effects substantially outperforms alternative strategies. The Highest CATE policy achieves treatment effects of 3.59-3.59 months (SE = 0.033), the largest magnitude across all policies. Random allocation serves as the baseline benchmark, generating treatment effects of 2.20-2.20 months (SE = 0.040). Current mentor-based selection performs remarkably well, achieving 3.25-3.25 months (SE = 0.100), a 48% improvement over random allocation. This validates that mentors successfully identify high-benefit participants through their evaluative selection process.
Table 10: Allocation Policy Comparison: Mentoring Program
Policy Baseline Time Treatment Effect
(months) (months)
Current 11.09 (0.131) -3.25 (0.100)
Random 11.36 (0.059) -2.20 (0.040)
Data-driven policies:
Highest CATE 11.08 (0.086) -3.59 (0.033)
Most Promising 9.71 (0.044) -2.27 (0.075)
Least Promising 13.44 (0.085) -1.94 (0.069)
Highest Chances 11.77 (0.152) -2.14 (0.079)
Notes: Comparison of allocation policies for the Mentoring program. Baseline Time is the predicted months to employment without treatment (lower is better); Treatment Effect is the predicted change in months from treatment (more negative = faster employment). Standard errors in parentheses. Current reflects mentor-selected participants. Random is random assignment benchmark. Highest CATE maximizes treatment effects; Most/Least Promising target best/worst baseline outcomes; Highest Chances minimizes post-treatment time to employment.
Notably, the Highest CATE algorithmic policy achieves an additional 10.4% improvement over mentor selection, increasing treatment effects from 3.25-3.25 to 3.59-3.59 months. We obtain the 0.34-month gain using only six readily-observable characteristics (education, location, age, experience) while mentors access rich soft information from interviews and holistic assessment. The overlap analysis (Table 9) reveals that 70.4% of mentor-selected participants would also be assigned under Highest CATE, demonstrating substantial concordance between expert judgment and data-driven targeting. The finding that simple observable characteristics can meaningfully improve upon expert selection—even when experts demonstrate strong performance—underscores the value of systematic data-driven approaches as complements to human judgment.
Policies optimizing objectives other than treatment effects yield substantially weaker results. Most Promising, which targets applicants with the strongest baseline prospects (9.71 months to employment), achieves treatment effects of only 2.27-2.27 months, barely exceeding random allocation despite successfully cream-skimming the highest-potential candidates. This demonstrates that strong baseline prospects do not predict strong treatment responsiveness. Conversely, Least Promising targets individuals with weakest baselines (13.44 months) but generates the smallest treatment effects (1.94-1.94 months), revealing an unfortunate equity-efficiency tradeoff: those who need help most benefit least from the program. Highest Chances, which optimizes final employment outcomes, achieves 2.14-2.14 months, marginally worse than random assignment. These comparisons underscore that maximizing program impact requires explicitly targeting predicted treatment effects rather than baseline characteristics or final outcomes, as treatment response heterogeneity does not align with baseline heterogeneity.
7 Conclusion
Changing labor markets push workers to want to transition into high-growth sectors, and facilitating these transitions is an important goal for labor policy. However, effectively designing such policies requires principled evaluation of various training programs in terms of their overall impact and their heterogeneity. In this paper, we provide evidence of the high effectiveness of two programs supporting women transitioning into the technology sector. This way we identify a class of interventions that can effectively narrow the gender gap in technology.
We show that a mentoring program and an online training program called Challenges aimed at developing labor market signals increase the probability of having a job in the technology sector by over 30% twelve months after the end of the program. We estimate that Mentoring increases technology employment from 46% to 61%, while Challenges increases it from 27% to 38%, with treatment effects persisting over the full observation period. Furthermore, the online training program developed during this research project is cheap to operate, at approximately \$15 per person, and suitable for scaling up, suggesting that it can have an impact extending beyond the current setting.
Our findings reveal strong patterns of heterogeneity in treatment effects. The benefits derived from the two programs depend on a number of participants’ characteristics including the educational background, length of professional experience, type of residence, and age. For example, we find that older participants that come from small towns have particularly high treatment effects in both programs, especially in Mentoring. Furthermore, certain groups benefit distinctly from one program over the other. Specifically, participants with short professional experience benefit from Challenges, while those with over five years of experience from Mentoring. These differential patterns provide evidence on the underlying mechanisms: Mentoring expands professional networks (increasing LinkedIn connections by 52.7, a 23% increase), while Challenges provides credible skill signals. Notably, network expansion occurs only in Mentoring, and heterogeneity patterns align with each program’s mechanism: those with weaker networks benefit most from Mentoring, while those with weaker credentials benefit most from Challenges.
For the foreseeable future, the demand for the two programs, particularly for Mentoring will exceed the number of available spots, and Dare IT will grapple with the question of whom to admit. We propose an admission policy that maximizes Dare IT’s impact, by prioritizing applicants with the highest predicted treatment effects. Consistent with the high degree of heterogeneity in the effectiveness of the two programs we find that such targeted admission policies can substantially increase Dare IT’s effectiveness. Specifically, we estimate that optimal targeting achieves outcomes 86% larger than random assignment at current capacity levels (30% of applicants). Notably, the value of targeting persists even as capacity expands: at 50% capacity per program, targeted allocation maintains outcomes 23% larger than random assignment. This demonstrates that targeting generates value both through selecting who participates and through matching individuals to their highest-benefit program. The complementarity between individual characteristics and program type creates allocation gains beyond simple capacity expansion.
A noteworthy aspect of Mentoring is that mentors have the autonomy to choose their mentees. To facilitate randomized access for our experiment, we asked mentors to shortlist two candidates and we randomized within these pairs. Comparing selected-but-control applicants to non-selected applicants reveals that mentors successfully identify promising candidates: mentor-selected participants achieve outcomes 48% larger than random allocation. However, mentors appear to prioritize candidates with strong baseline prospects, those likely to secure technology employment even without the program, rather than those who would benefit most from participation. We evaluate an alternative algorithmic allocation policy that selects participants based on predicted treatment effects using only six readily-observable characteristics (education, location, age, experience). Despite mentors’ rich soft information from interviews and holistic assessment, algorithmic targeting achieves outcomes 63% larger than random allocation, representing an additional 11% improvement over mentor judgment. This finding is particularly striking given that algorithmic and mentor-based selections exhibit 70% concordance: the algorithm largely validates expert choices while identifying additional high-benefit candidates. The results demonstrate that systematic patterns in treatment response exist even after expert selection captures most available heterogeneity, and that simple demographic targeting can complement human judgment to enhance program effectiveness.
This study comes with certain limitations worth noting. First, the limited participant size in both programs led to wide confidence intervals in several estimates, leaving room for uncertainty regarding the exact impacts of Mentoring and Challenges. Secondly, we could not pinpoint the exact mechanisms driving the outcomes for both programs. Although our findings suggest that Mentoring likely enhances participants’ professional networks and teaches how to utilize them, and Challenges supports participants in the development of a credible signal of their skills, a more in-depth exploration is required to understand the specific mechanisms through which the training provided by both programs impacts employability. Such an analysis could also help scale up Mentoring by identifying the crucial aspects of the program and directing Dare IT resources to them. Lastly, the participants from the first edition of Challenges may not be representative of those in later editions. Early adopters of such programs might experience greater benefits. To validate the program’s continued efficacy, subsequent evaluations are essential.
Appendix
Appendix A Challenges program
The Challenges program is designed for the same population as the Mentoring program: individuals who possess technical competencies but have not yet secured employment in the technology sector. The primary aim of the Challenges program is to assist these individuals in developing credible credentials that can effectively signal their skill set. In its first edition, the Challenges program provided two distinct tracks: User Experience Design (UX) and Front-end Development in React (front-end). The selection of these specific tracks was influenced by several determinants. Primarily, there was a significant volume of applications for the Mentoring program related to these two domains. Furthermore, Dare IT possessed both internal expertise and established partnerships in these areas. Lastly, a pronounced market demand existed for professionals in these specializations in Poland.
To achieve the program’s scalability goal and streamline the application process, unlike the Mentoring program, the review of the minimum technical skills during application was based solely on applicants’ declaration. The registration website explicitly mentioned the necessity for an advanced understanding of Figma for those interested in UX design, and React for those leaning towards front-end engineering as prerequisites for the Challenges program. The applicants had to confirm that they indeed have the minimum skill level. Additionally, English proficiency was mandated due to the inclusion of course content delivered in English.
The Challenges program spanned 16 weeks. Initially, the first two weeks focused on an introductory event and a refresher course in Figma and React. The concluding two weeks were allocated for the refinement of the final projects, HR workshops, and a closing event. Central to the program, the intermediate 12 weeks were primarily aimed at the progression of an individual project. Within both specializations, experts from partner organizations, specifically Flying Bisons and Grover – Polish tech companies, were asked to prepare six tasks. These tasks, inspired by their recent undertakings, encapsulated the comprehensive procedure of designing a mobile application for the UX domain and detailed the facets of front-end engineering for the respective specialization.
The program is based on two-week cycles culminating on Thursdays when submissions are expected and news assignments are posted. Every subsequent Friday, post-submission, an evaluative session with a subject-matter specialist takes place. During this session, a selection of submissions undergoes review, identifying prevalent errors and underscoring industry best practices. This grants participants the opportunity to amend and refine their submissions from the previous week. The domain experts, either mentors from Dare IT or other professionals with an appropriate level of seniority, carry out the reviews. Each assignment has been designed to take approximately eight hours of time to finalize. As the program approaches its conclusion, participants engage with a Human Resources specialist. This guidance is primarily aimed at instructing participants on leveraging their newly-acquired qualifications, both as skill indicators in job applications and as a resource during subsequent interviews.
In the first week of the UX path, participants conceptualize their project. They are given the flexibility to either devise their own application idea or choose from two provided project themes, followed by a detailed analysis of market competitors and benchmarks. The subsequent phases guide participants through a structured journey from ideation to delivery. They familiarize themselves with tools like Personas, User Journey Maps, and User Stories, ensuring a systematic organization of the information gathered. As the program advances, participants delve deeper into content hierarchy, Information Architecture, and user pathways, culminating in the creation of UX wireframes. The latter weeks emphasize usability testing, where participants craft interactive prototypes, refining them based on feedback. Crucially, they also submit project presentations and receive feedback on them.
In the case of the front-end path, in the first weeks, participants created reusable UI components, often termed Atoms. This initial phase emphasized the foundational importance of efficient design. In the subsequent weeks, attendees delved into topics of server communication: creating APIs, handling different requests, and DELETE queries. Following these modules, the middle segment of the program trained participants in creating models and form-based data collection. Later, the curriculum shifted towards data visualization using charts and the integration of real-time user feedback systems, e.g., notifications. These modules were designed to underscore the advantages of adept data presentation and heightened user experience. Finally, the program introduced participants to TypeScript, showcasing its emerging prevalence in modern web applications.
In addition to the assignments and review sessions, participants were granted access to the exclusive Dare IT community on the Slack platform. During the kick-off meeting, the participants were recommended to collaborate in teams to tackle the assignments. A dedicated process was created to support collaboration; participants were asked whether they wanted to be matched with a group. If they were interested, they received access to a group’s channel. [^9] The emphasis on collaborative efforts and peer support was further amplified in the ensuing editions of the Challenges program, where matching events were organized. Within each path’s Slack channel, distinct discussion forums were established, allowing participants to pose queries, circulate resources, and engage in discourse about assignment components. Activity within this Slack community was consistently high, with frequent discussions across multiple channels.
The first edition of Challenges has attracted a lot of popularity amongst the participants, applicants that we could not accommodate due to capacity limits, and the Polish tech community more broadly, thereby galvanizing its further development. Currently, the program is instituted on a trimestral cycle, extending its offerings to encompass Manual Testing, Automated Testing in Python, Cloud Computing, and User Interface Design. Furthermore, in the summer of 2022, the program was adapted to cater to Ukrainian refugees in Poland. This edition was delivered in Ukrainian, incorporating bespoke modules tailored to guide foreigners in navigating the Polish job market.
Appendix B Covariate balance
In Figure 5 we show the balance of covariates in the mentoring experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances.
Figure 5: Covariate Balance Mentoring.
Note: Balance across treatment and control of the main variables. Points correspond to the absolute difference in means between treatment and control divided by the square root of the sum of variances in treatment and control.
In Figure 6 we show the balance of covariates in the Challenges experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances. The only variable that has a statistically significant differnece is social science.
Figure 6: Covariate Balance Challenges.
Note: Balance across treatment and control of the main variables. Points correspond to the absolute difference in means between treatment and control divided by the square root of the sum of variances in treatment and control.
Appendix C Kaplan Meier Survival Curves
We present Kaplan Meier estimates of the survival function $S(t)$, defined as the probability of not yet obtaining a technology job by month $t$.
Figure 7 shows the survival functions for the Mentoring program. Relative to both the control group and applicants, participants in the Mentoring treatment group find technology jobs more rapidly: their survival curve falls more steeply, indicating earlier entry into technology jobs. The difference between treated and control groups emerges within the first few months and grows steadily over the 15-month horizon. We can also notice that there is a substantial gap between the applicants and both experimental groups which emerges already during the first four months,
Figure 7: Survival Curves Mentoring Program.
Note: Kaplan Meier survival curves for the Mentoring program, including treated participants, control group, and applicants. A steeper decline indicates earlier job placement.
Figure 8 presents analogous estimates for the Challenges program. Treated participants also transition into jobs more quickly than the control group or applicants, though the gap is smaller in magnitude than in the Mentoring program.
Figure 8: Survival Curves Challenges Program.
Note: Kaplan Meier survival curves for the Challenges program, including treated participants, control group, and applicants. Treated participants transition to jobs earlier, though the effect is smaller than in Mentoring.
Appendix D Comparison of salary levels from Glassdoor
We consider two types of jobs as a job in technology:
  1. 1.
    All jobs in tech companies other than finance, regulatory, legal, accounting, and HR, where tech companies include firms in software development, testing, and sales; data analytics; IT services; digital marketing; and online platforms (including peer-to-peer platforms, and online sales).
  2. 2.
    Jobs in non-tech companies that involve software development and testing, IT support, and data analytics. In our context, this category includes jobs in banks and management consulting agencies.
We analyze whether the two categories of jobs pay similar salaries. To do that we searched for salary estimates of all unique job titles that fall into the two categories using Glassdoor.com. We considered only exact matches and salary estimates for jobs in Poland. In total, we have 265 unique salary estimates. We find that the mean monthly salary for tech jobs in tech firms (category 1) is PLN 8634 and for tech jobs in non-tech firms it is PLN 8866. The difference is statistically insignificant (p-value 0.9). Figure 9 shows histograms of of salary estimates.
Figure 9: Histograms of salary estimates from Glassdoor.
Note: Histograms of salary estimates from Glassdoor. In green, we show tech jobs in tech firms (category 1), in blue tech jobs in non-tech firms (category 2).
Appendix E Outcomes measured in surveys
We use data from outcomes surveys to provide additional evidence of the effectiveness of the two programs that give insights into the mechanisms through which participation boosts outcomes. There were three surveys: two main ones carried out approximately four months after the end of each program, and a shorter survey right after the end of the mentoring program from which we obtained information on salaries in jobs found during or right after the mentoring program. As discussed in Section 4, results based on the outcomes survey suffer from a selection problem. Thus, we interpret the findings based on this data as suggestive evidence.
The number of job offers.
Figure 10 shows the number of job offers received by survey respondents broken by experimental groups. We find that subjects in the control group (in green) were more likely to report receiving zero job offers than the subjects in the treated group (in orange); for any non-zero range, there is a higher share of the treatment group rather than the control group. The largest difference between the experimental groups is amongst subjects reporting more than ten job offers.
Figure 10: Reported numbers of job offers: Mentoring.
Note: The number of job offers received by subjects in the mentoring experiment that responded to the survey. Each bar corresponds to the share of respondents from an experimental group. Orange represents the treated group, and green represents the control group.
In Figure 11, we present evidence of the differences in the number of job offers across treatment and control groups of the Challenges experiment. We find results analogous to those from the mentoring program. The lowest category, here less than two job offers, has a higher share of control group respondents than treated group respondents. In contrast, treated subjects are more likely to report a higher number of job offers. The difference between treatment and control is more pronounced in the mentoring program.
Figure 11: Reported numbers of job offers: Challenges.
Note: The number of job offers received by subjects in the Challenges experiment that responded to the survey. Each bar corresponds to the share of respondents from an experimental group. Orange represents the treated group, and green represents the control group.
Salary ranges.
Figure 12 presents the share of respondents to mentoring outcomes surveys declaring a salary within a salary range (ranges were pre-defined in the survey). Orange bars indicate the treatment group, and the green ones are the control group. The right panel is based on the survey carried out right after the end of the program, while the left panel is from the survey conducted four months after the end of the program.
Figure 12: Salary ranges: Mentoring.
Note: Reported salary ranges by subjects who found new jobs after the start of the mentoring program. Orange represents the treated group, and green represents the control group. The length of a bar represent the share of respondents declaring a salary within the range. The right panel survey was conducted right after the end of the program; the left panel survey was conducted four months after the program.
Evidence presented in Figure 12 suggests that subjects in the treated group of the mentoring experiment that found a new job are more likely to have a salary in a high salary range.
In Figure 13, we report analogous results from the Challenges outcome survey. We find that a higher share of the survey respondents in the control reported salary in the lowest range as compared to the treatment group. We also find that the opposite is true for the highest salary range.
Figure 13: Reported salary range Challenges.
Note:Reported salary ranges by subjects who found new jobs after the start of the Challenges experiment. The length of a bar represent the share of respondents declaring a salary within the range. Orange represents the treated group, and green represents the control group.
Negotiation.
Respondents that reported finding a job were asked whether they negotiated the conditions of their employment. Amongst the mentoring group, 33% of treated subjects that found a new job negotiated the terms of the new job, while only 24% of the control group did so. The difference is statistically insignificant.
Appendix F Proportional hazards assumption
The Cox proportional hazards model assumes the hazard rate of an event is a product of a baseline hazard and an exponential function of the covariates. Thus, in the analysis, we assume the effects of different predictors are constant over time. To test that, we correlate Schoenfeld residuals from the baseline Mentoring and Challenges models with months. Table 11 presents the results. We do not find any statistically significant correlations.
Table 11: Proportional hazards assumption test
Mentoring model Challenges model
chisq1 df1 p1 chisq2 df2 p2
Treatment 0.232 1 0.630 1.301 1 0.254
City 0.149 2 0.928 1.863 2 0.394
Degree 1.909 4 0.752 5.832 4 0.212
STEM 0.001 1 0.971 0.002 1 0.968
Prof experience 0.187 1 0.665 1.723 1 0.189
Social science 0.074 1 0.785 - - -
Family friends IT 0.074 1 0.785 - - -
Mother 0.164 1 0.685 - - -
Applied before 0.087 1 0.768 - - -
First time mentor 0.956 1 0.328 - - -
Managerial experience 0.762 1 0.383 - - -
Career changer 0.043 1 0.836 - - -
Years of experience Mentor 0.428 1 0.513 - - -
Age 2.716 3 0.438 - - -
GLOBAL 9.595 20 0.975 13.130 9 0.157
Note: Diagnostics of the proportional hazard model. The first three columns show results for the Mentoring model, and the last three for Challenges. City is a discrete variable with levels: Small Town, Large City, and Warsaw; degree is also a discrete variable with levels at: no education mentioned, non-traditional credential, college degree, masters, and postgraduate. Prof experience is the number of years between the start date of the first employment of program participants and 2023.
Appendix G Bounds on the Average Treatment Effect Estimates
In this section, we perform robustness checks on the average treatment effect estimates. First, we consider several cases in which we change which subjects are included in the analysis. Second, we estimate AT bounds following lee2021bounding. Third, we estimate the average treatment effects at 4 months after the programs.
Robustness to Sample Construction and Attrition.
Our main analysis excludes participants who secured technology employment between randomization and program launch, as well as mentors who declined to participate in the experimental design. While these exclusions yield conservative estimates, we verify that our conclusions are robust to alternative sample definitions. Table 12 presents treatment effect estimates under seven alternative specifications. For each, we report sample sizes, mean outcomes by treatment status, and the estimated average treatment effect using difference-in-means without covariate adjustment. All estimates use the twelve-month tech employment outcome.
We consider four specifications for Mentoring. First, our baseline estimate (15.3 percentage points, SE 5.7) uses 151 treatment and 147 control participants, reflecting the 4 mentor pairs where both candidates were assigned to treatment due to control participants securing early employment. Second, restricting to the 126 perfectly matched pairs yields a slightly smaller but still substantial effect (11.9 percentage points, SE 6.3). Third, including the 4 participants who found jobs before program launch—coding their outcomes as 1—increases the estimated effect to 16.4 percentage points. Fourth, additionally including the 9 non-participating mentors under the assumption their hypothetical mentees would not have found jobs reduces the estimate to 13.0 percentage points. All four specifications yield treatment effects between 11.9 and 16.4 percentage points, each statistically significant at the 5% level.
For Challenges, we examine three specifications. Our baseline estimate (11.3 percentage points, SE 4.8) excludes 17 treatment participants who secured technology jobs before program launch. Including these individuals and coding their outcomes as 1 increases the effect to 17.1 percentage points—as expected, since we add successful job-finders to the treatment group. Finally, we address the 17 treatment and 7 control participants with missing twelve-month LinkedIn data. Under the conservative assumption that all missing treatment observations had outcomes of 0 and all missing control observations had outcomes of 1—the scenario most unfavorable to finding a treatment effect—we still estimate a positive effect of 10.5 percentage points (SE 4.6, p=0.02).
Table 12: Robustness to Sample Construction and Attrition
Specification Sample Size Mean Outcome ATE
Treated Control Treated Control (SE)
Panel A: Mentoring
   Main specification 151 147 0.609 0.456 0.153***
   Balanced pairs only 126 126 0.603 0.484 0.119*
   Include pre-program employment 155 147 0.619 0.456 0.164***
   Include non-participants 164 147 0.585 0.456 0.130**
Panel B: Challenges
   Main specification 166 218 0.380 0.266 0.113**
   Include pre-program employment 183 218 0.437 0.266 0.171***
   Pessimistic attrition bounds 200 227 0.400 0.295 0.105**
Note:Estimates of average treatment effects on twelve-month technology employment under alternative sample definitions. Main specifications correspond to Table 2. Balanced pairs only (Mentoring) restricts to the 126 mentor pairs with both treatment and control observations. Include pre-program employment adds participants who secured technology jobs between randomization and program launch. Include non-participants (Mentoring) adds 9 mentors who declined experimental participation, assigning their hypothetical mentees outcomes of 0. Pessimistic attrition bounds assign outcomes of 0 to 17 missing treatment observations and 1 to 7 missing control observations. All estimates use difference-in-means without covariate adjustment. Standard errors in parentheses. *** p<0.01, ** p<0.05, * p<0.1.
These robustness checks establish three key points. First, our main estimates are insensitive to reasonable alternative sample construction decisions, with all specifications yielding economically large and statistically significant treatment effects. Second, the range of point estimates across specifications (11.9 to 16.4 for Mentoring, 10.5 to 17.1 for Challenges) provides additional bounds on the true treatment effects that are consistent with our main findings. Third, even under extremely conservative assumptions about missing data—assumptions, the Challenges program shows significant positive impacts. This addresses potential concerns that selective attrition might bias our twelve-month estimates.
Lee Bounds for Selective Attrition.
Our twelve-month outcome measurement experienced modest attrition in the Challenges program: 6% had missing LinkedIn data due to profile URL changes or privacy setting modifications. We assess whether selective attrition could bias our estimated treatment effect using the bounding procedure of lee2009training. We consider the most conservative case, in which subjects that found jobs before the start of the program are not considered; thus, we focus only on the attrition during the program, due to missing LinkedIn data.
The Lee bounding procedure addresses selective attrition by considering worst-case scenarios about which individuals are missing. Since treatment experienced higher attrition, the conservative assumption for estimating positive effects is that the best-performing treatment participants were most likely to attrite (e.g., successful job-finders who deactivated LinkedIn). To construct the lower bound, we trim 6.8% of observations from the upper tail of the observed treatment outcome distribution, the proportion representing excess attrition relative to control, and compare the mean of this trimmed distribution to the full control group mean. The upper bound uses all observed treatment outcomes compared to all observed control outcomes. We estimate Lee bounds of [0.065, 0.114]. The upper bound of 0.114 is nearly identical to our main point estimate of 0.113, while the positive lower bound of 0.065 indicates that even under worst-case assumptions about selective attrition, the Challenges program effect remains economically meaningful. The 95% confidence interval on the lower bound is [-0.029, 0.159]; while this marginally includes zero due to sampling variability in the observed outcomes, the point estimate is bounded away from zero.
Average Treatment Effect Estimates at 4 Months.
Our main analysis uses employment outcomes measured twelve months post-program. However, we were unable to collect LinkedIn data for 17 Challenges treatment participants and 7 control participants at the twelve-month measurement due to profile URL changes or privacy setting modifications. A natural concern is whether selective attrition biases our estimated treatment effects. To address this concern, we re-estimate average treatment effects using employment outcomes measured four months after program completion. At this earlier measurement point, we successfully collected complete LinkedIn data for all randomized participants in both programs. Any attrition-induced bias is therefore eliminated at this time horizon.
Table 13 presents the results. We find substantively similar treatment effects at the four-month horizon compared to our main twelve-month estimates (reported in Table 2). For Mentoring, the four-month treatment effect is 15.3 percentage points (SE 5.7), nearly identical to the twelve-month effect of 14.5 percentage points. For Challenges, the four-month effect is 8.9 percentage points (SE 4.3), compared to 11.2 percentage points at twelve months. The larger point estimate at twelve months suggests that the Challenges program’s impact may strengthen over time as participants leverage their portfolios in job applications.
Table 13: Average Treatment Effects of Mentoring and Challenges at 4 Months.
Mentoring Challenges
   ATE 0.153 0.089
   SE (0.057) (0.043)
   ATE % baseline 33.7 45.3
   SE % baseline (12.6) (21.8)
   N Treatment 151 183
   N Control 147 225
Note: Estimates of the average treatment effect of Mentoring and Challenges at 4 months after the end of each program. Estimates obtained using the difference-in-means estimator.
Importantly, these four-month estimates use the full experimental sample without any missing observations, confirming that our main findings are not an artifact of selective attrition. The consistency between four-month and twelve-month estimates also provides evidence that treatment effects do not attenuate over time—a key finding given that some labor market training programs show fadeout of impacts card2018works. If anything, the point estimates suggest persistent or growing effects, particularly for Challenges, though the conclusions are limited to 12 months from the start of each program.
Appendix H Policy estimation algorithms
Algorithm 1 describes the estimation and evaluation of optimal policy.
Algorithm 1 Optimal policy
1. Considering the dataset of all applicants to both programs, randomly split the dataset into train and test with equal probability at the applicant level, denote the resulting sets as $I^{train}=\left\{1,..,\mathcal{I}^{train}\right\}$ and $I^{test}=\left\{1,..,\mathcal{I}^{test}\right\}$.
2. In the training dataset, estimate outcome model using cross-validation: $\mu_{train}$
3. In the test set, construct predicted treatment effects using predictions from model $\mu_{train}$. Obtain $\tau^{Op,train}_{i,t}=E[\hat{Y}(p)_{train}-\hat{Y}(O)_{train}|X=x_{i,t},i\in I^{test}]$, where $\hat{Y}(p)_{train}$ is the predicted outcome under program $p$ constructed using $\mu_{train}$ and $\hat{Y}(O)_{train}$ is the predicted outcome under Out of Dare IT and $t\in\left\{1,...,15\right\}$ denotes months. Compute mean treatment effects per user as: $\tau^{Op,train}_{i}=1/t\times\sum_{t=1}^{15}\tau^{Op,train}_{i,t}$,
4. Assign treatment to maximize treatment effects subject to capacity constraint. Let $Q^{p}$ be the capacity limit of program $p$ and $z_{ip}$ an indicator variable taking the value of one when applicant $i$ is assigned to program $p$ and zero otherwise. We solve the following constrained optimization problem:
$\max_{z_{ip}}\sum_{i=1}^{I}\sum_{p=1}^{P}z_{ip}\tau_{i}^{Op,train}\text{ s.t. }\sum_{i=1}^{I}z_{ip}\leq Q^{p}\forall_{p}\text{ \& }\sum_{i=p}^{P}z_{ip}=1\forall_{i\in I^{test}}.$
The first constraint ensures that the capacity constraints are not violated. The second one is that every applicant is assigned to one program. There is no capacity limit on being Out of Dare IT. We use LP Solve algorithm to solve the problem. We obtain optimal allocation $\mathcal{A}_{X,Q}^{*}=\left\{a_{i}^{*},...,a_{\mathcal{I}^{test}}^{*}\right\}$,
5. Using the test set, estimate new outcome and propensity models using cross-fitting and obtain predictions: $\hat{\mu}_{i,k}$ and $\hat{e}_{i,k}$ for all $i\in I^{test}$. See Appendix LABEL:cross_fit_appendix for details of the cross-fitting procedure. Obtain $\hat{Y}_{i,k}(a^{*})$ the AIPW estimates of the predicted outcomes using cross-fitted models trained in the test set,
6. Obtain $\hat{V}_{X,Q}^{*}=\frac{1}{|I^{test}|}\times\sum_{i=1}^{\mathcal{I}^{test}}\hat{Y}_{i,k}(a^{*})$ as the mean of predicted outcomes under the allocation $\mathcal{A}_{X,Q}^{*}$. Estimate standard errors clustered at the applicant level: $\sigma_{X,Q}^{*}$. $\triangleright$
return $(\mathcal{A}_{X,Q}^{*},\hat{V}_{X,Q}^{*},\sigma_{X,Q}^{*},\hat{Y}_{1,k}(a^{*}),...,\hat{Y}_{\mathcal{I}^{test},k}(a^{*}))$
H.1 Assignment Models
We pool data from Mentoring experimental participants (N=298), Challenges experimental participants (N=408), non-selected Mentoring applicants (N=468), and late Challenges applicants (N=160), yielding 1,334 total applicants. All individuals are observed over fifteen months post-application. We randomly split this dataset 50-50 at the applicant level into training (N=667) and testing (N=667) sets.
Using the training set, we estimate predicted treatment effects for each program relative to non-participation. We fit two causal survival forests (athey2019generalized; cui2023estimating): one comparing Mentoring participants to applicants who did not participate in any program (including the control group of the Mentoring experiment), and another comparing Challenges participants to non-participants.
Each causal survival forest is estimated with the following parameters: 2,000 trees, honest splitting, sample fraction of 0.5 per tree, minimum node size of 5, and horizon of 15 months. Covariates include: STEM degree, residence type (Warsaw, small town), master’s degree, age over 30, professional experience (continuous), and UX specialization. These settings follow standard recommendations in the generalized random forests (grf) package (tibshirani2024grf).
The causal survival forest estimates conditional average treatment effects on the restricted mean survival time (RMST) scale—the average time until technology employment over our observation horizon. More negative predicted effects indicate faster job finding under treatment. We apply the trained models to the test set to obtain predicted treatment effects $\hat{\tau}_{i}^{M}$ (Mentoring vs. nothing) and $\hat{\tau}_{i}^{C}$ (Challenges vs. nothing) for each test-set applicant.
Appendix I Evaluation Models
While assignment models determine which applicants receive which program under various policies, evaluation models estimate the causal effects of these assignments. The credibility of our policy comparisons rests on these evaluation models accurately estimating treatment effects for the allocated applicants. If evaluation models correctly estimate treatment effects, our policy value estimates remain valid even if assignment models imperfectly rank applicants.
We use the held-out test set (N=667) for all policy evaluations. This sample includes experimental participants and non-participants from both programs, ensuring we can estimate counterfactual outcomes under different assignment policies. All individuals are observed over the same fifteen-month post-application horizon.
Model specification.
For each policy evaluation, we estimate individual-level treatment effect predictions using causal survival forests (athey2019generalized; cui2023estimating). Within each cross-fitting fold (detailed below), we fit two causal survival forests: one comparing Mentoring participants to applicants who did not participate in any program, and another comparing Challenges participants to non-participants.
Each causal survival forest uses identical parameters to the assignment models: 2,000 trees, honest splitting, sample fraction of 0.5 per tree, minimum node size of 5, and horizon of 15 months. The covariate set is identical to that used in assignment models: STEM degree, residence type (Warsaw, small town), master’s degree, age over 30, professional experience (continuous), and UX specialization. Model fitting follows the grf package implementation (tibshirani2024grf).
The causal survival forest produces augmented inverse propensity weighted (AIPW) scores $\psi_{i}^{M}$ and $\psi_{i}^{C}$ for each test-set individual, representing individual-level treatment effect estimates from Mentoring and Challenges relative to non-participation, respectively. The forest automatically constructs these doubly-robust scores by combining: (1) propensity score estimates $\hat{e}(X_{i})$ from a survival forest predicting treatment assignment, and (2) conditional mean predictions $\hat{\mu}(X_{i},W)$ from honest regression forests predicting restricted mean survival time conditional on treatment status and covariates. These components are combined via the AIPW formula, yielding estimates that remain consistent if either the propensity or outcome model is correctly specified.
Cross-fitting procedure.
To avoid overfitting and ensure valid inference, we employ repeated $K$-fold cross-fitting on the test set. Specifically, for each repetition $m\in\{1,\ldots,100\}$, we randomly assign observations into $K=5$ folds. Randomization is stratified by experimental group (Mentoring treated, Mentoring control, Challenges treated, Challenges control, non-participants) to ensure each fold contains observations from all groups. For each fold $k\in\{1,\ldots,5\}$, we:
  1. 1.
    Train causal survival forests using data from all folds except $k$ (the training sample for fold $k$).
  2. 2.
    Predict AIPW scores for observations in fold $k$: $\psi_{i,k}^{M,m}$ and $\psi_{i,k}^{C,m}$ for each individual $i$ in fold $k$, along with corresponding variance estimates $\hat{V}_{i,k}^{M,m}$ and $\hat{V}_{i,k}^{C,m}$.
After completing all folds in repetition $m$, each individual has predicted treatment effects $\psi_{i}^{M,m}$ and $\psi_{i}^{C,m}$ estimated on data that did not include that individual. We then aggregate across the 100 repetitions: following chernozhukov2017double, we take the median prediction across repetitions for each individual: $\psi_{i}^{M}=\text{Median}(\psi_{i}^{M,1},\ldots,\psi_{i}^{M,100})$ and $\psi_{i}^{C}=\text{Median}(\psi_{i}^{C,1},\ldots,\psi_{i}^{C,100})$. Variance estimates are similarly aggregated by taking the median across repetitions.
Policy value estimation.
Given an allocation policy $\pi$ that assigns each test-set individual to Mentoring, Challenges, or neither, we estimate the policy value as:
$\hat{V}(\pi)=\frac{1}{n_{M}}\sum_{i:\pi(i)=M}\psi_{i}^{M}+\frac{1}{n_{C}}\sum_{i:\pi(i)=C}\psi_{i}^{C}$
where $n_{M}$ and $n_{C}$ are the number of individuals assigned to Mentoring and Challenges, respectively. Standard errors are computed using the aggregated variance estimates from the causal survival forests, accounting for the within-arm averaging. For comparisons between policies or programs, we compute appropriate differences in AIPW scores and propagate variances assuming independence across programs.
This approach separates the assignment decision (which applicants receive which program) from the evaluation of that decision (estimating treatment effects for the assigned groups). The validity of our policy comparisons depends on the evaluation models accurately estimating treatment effects, not on the assignment models perfectly ranking applicants.
I.1 Model Validation: Recovering Experimental Effects
To validate our evaluation model specification, we assess whether cross-fitted AIPW scores correctly recover experimental average treatment effects. This diagnostic tests whether our models produce unbiased treatment effect estimates in the setting where we have the benchmark from the randomized experiment.
Validation procedure.
We train causal survival forests following the same process as before, using 5-fold cross-fitting and generate out-of-sample AIPW predictions for each participant; we use the experimental samples. We compare these individual-level predictions to the experimental ATE by examining the distribution of deviations: $\delta_{i}=\hat{\tau}_{i}^{AIPW}-\hat{\tau}^{RCT}$. Under correct model specification, these deviations should center at zero with dispersion reflecting genuine treatment effect heterogeneity plus estimation noise.
Figure 14 presents the results. The red solid vertical line indicates the RCT benchmark (zero deviation), with shaded regions marking 95% confidence intervals around both the RCT estimate (red) and the mean AIPW deviation (gray). The black dashed line shows the mean deviation from the RCT estimate.
Figure 14: Model Validation: Distribution of Deviations from RCT Estimates
Refer to caption Refer to caption
(a) Mentoring Program (b) Challenges Program
Notes: Distribution of deviations between cross-fitted AIPW predictions and experimental ATE benchmarks. Each panel shows $\delta_{i}=\hat{\tau}_{i}^{AIPW}-\hat{\tau}^{RCT}$ for experimental participants. The red vertical line marks zero deviation (perfect recovery), with red shading indicating the 95% confidence interval around the RCT estimate. The black dashed line shows the mean AIPW deviation, with gray shading for its 95% confidence interval. Causal survival forests trained exclusively on experimental samples using 5-fold cross-fitting. Mentoring: mean deviation = $0.08$ months (SE = 0.03); Challenges: mean deviation = $0.06$ months (SE = 0.04). Deviations statistically indistinguishable from zero in both programs.
For the Mentoring program (Panel a), AIPW predictions center tightly around the RCT benchmark (SE, with a mean deviation of $0.08$ months (SE = 0.03). The distribution is approximately symmetric and concentrated near zero, indicating our cross-fitted estimates successfully recover the experimental ATE. The modest spread in the distribution reflects both the treatment effect heterogeneity among participants and sampling variation in the predictions.
For the Challenges program (Panel b), AIPW predictions show similarly strong performance, with a mean deviation of $0.06$ months (SE = 0.04). While the distribution exhibits slightly more dispersion than Mentoring—consistent with Challenges’ larger experimental sample and potentially greater heterogeneity—the mean deviation remains statistically indistinguishable from zero. The slight positive mean suggests modest underestimation, but the bias is economically negligible (less than 2% of the experimental ATE).
The cross-fitted causal survival forests accurately recover experimental treatment effects, demonstrating that our evaluation approach is well-calibrated and unbiased within the experimental sample. This gives us confidence that the models correctly estimate treatment effects where treatment is randomized.